### v2

# Dependability of UAV-Based Networks and Computing Systems: A Survey

Abstract.
Uncrewed Aerial Vehicle (UAV) computing and networking are becoming a fundamental computation infrastructure for diverse cyber-physical application systems. UAVs can be empowered by AI on edge devices and can communicate with other UAVs and ground stations via wireless communication networks. Dynamic computation demands and heterogeneous computing resources are distributed in the system and need to be controlled to maintain the quality of services and to accomplish critical missions. With the evolution of UAV-based systems, dependability assurance of such systems emerges as a crucial challenge. UAV-based systems confront diverse sources of uncertainty that may threaten their dependability, such as software bugs, component failures, network disconnections, battery shortages, and disturbances from the real world. In this paper, we conduct systematic literature reviews on the dependability of UAV-based networks and computing systems. The survey report reveals emerging research trends in this field and summarizes the literature into comprehensive categories by threat types and adopted technologies. Based on our literature reviews, we identify eight research fields that require further exploration in the future to achieve dependable UAV-based systems.
1. Introduction
Uncrewed Aerial Vehicles (UAVs), also known as drones, have rapidly evolved from niche tools to critical components in a wide array of cyber-physical systems (CPS), enabling applications such as disaster response (Wei et al., 2022), precision agriculture (Vélez et al., 2024), surveillance (Zhang et al., 2021c), and logistics (Xu et al., 2024a). By integrating advanced computing paradigms and capabilities at the edge and leveraging wireless communication networks within 5/6G connectivity (De Paula Soares et al., 2023), UAVs can perform complex tasks autonomously or collaboratively, often in dynamic and unpredictable environments. Equipped with artificial intelligence (AI) and interconnected with other UAVs and ground stations (Ning et al., 2024), these systems form a distributed computing and networking infrastructure that supports real-time decision-making and mission execution. However, the increasing complexity, autonomy, and interconnectivity of UAV-based CPS present substantial challenges to their dependability. This concept encompasses reliability, availability, maintainability, and safety, all of which are critical to the secure and effective operation of these systems (Avizienis et al., 2004).
Dependability has long been a cornerstone of cyber-physical systems (Miclea and Sanislav, 2011). Traditional CPS frameworks emphasize fault tolerance, redundancy, and rigorous verification to mitigate failures. UAV-based systems, as a specialized class of CPS, inherit these principles but face unique challenges due to their aerial mobility, limited energy resources, and exposure to dynamic environmental conditions (e.g., wind gusts, electromagnetic interference) (Pacheco et al., 2021; Xue et al., 2023b; Zhan and Xu, 2024). For instance, while ground-based CPS may rely on stable power sources and fixed communication infrastructure, UAVs must balance limited battery life with computational demands, all while maintaining connectivity in ad-hoc airborne networks. These fundamental differences require a re-examination of existing dependability paradigms to effectively address the specific vulnerabilities and operational demands of UAV-based systems.
Dependability in UAV-based systems encompasses a range of attributes, including reliability, availability, safety, and resilience, all of which are essential to maintaining Quality of Service (QoS) and accomplishing mission objectives (Gupta and Gupta, 2022; Farrukh and West, 2023; Avizienis et al., 2004). These systems operate in environments where they face diverse sources of uncertainty, such as hardware and software failures, network disruptions, energy constraints, environmental disturbances, and communication breakdowns. For instance, a UAV performing a search-and-rescue mission may encounter battery depletion or signal loss, potentially leading to mission failure or safety hazards (Kaymaz et al., 2024). Recent studies have identified energy constraints as a primary threat, as battery depletion often causes mission aborts in long-duration operations (Lin et al., 2021). Morereover, network disruptions in UAV swarms can hinder coordination and data sharing due to intermittent connectivity (Alioua et al., 2018).
Additionally, environmental factors like extreme weather exacerbate hardware and software failures, further complicating dependable operation. As UAV-based systems become more integral to critical applications, ensuring their dependability has emerged as a pressing research challenge, necessitating a systematic understanding of the threats they face and the methodologies to mitigate them.
Efforts to improve dependability of UAV-based systems have led to a variety of proposed approaches, such as heuristic optimization techniques for task offloading to edge servers (Baktayan et al., 2022), energy harvesting mechanisms to extend flight duration (Kim and Kim, 2021), and fault-tolerant protocols for 5G-enabled UAV swarms to ensure reliable communication under disruptions (Damigos et al., 2023). Simulation techniques, including Markov models, have also been employed to evaluate these solutions in dynamic environments (Bai et al., 2023a; Zhao et al., 2024). Despite these advancements, the literature remains fragmented, with studies often focusing on isolated aspects of dependability, including energy management or network reliability, without integrating both networking and computing dimensions.
This survey report aims to provide a comprehensive overview of the dependability of UAV-based networks and computing systems, synthesizing the state-of-the-art research and identifying key trends, challenges, and opportunities. To achieve this, we conducted a Systematic Literature Review (SLR) following a rigorous methodology to ensure a thorough and unbiased selection of relevant studies. We first identified a set of keywords related to dependability of UAV-based systems and system type, such as ”UAV reliability,” ”dependability”, and ”UAV computing system”. Using these keywords, we searched five major literature databases, including IEEE Xplore 111https://ieeexplore.ieee.org/Xplore/home.jsp, ACM Digital Library 222https://dl.acm.org/, SpringerLink 333https://link.springer.com/, ScienceDirect 444https://www.sciencedirect.com/, and ISI Web of Science 555http://www.isiknowledge.com, collecting publications from the past decade (January 2015 to July 2024). This initial search yielded a total of 1,848 articles. We then performed a meticulous screening process by reviewing the titles and abstracts of these articles, applying inclusion and exclusion criteria to focus on studies directly addressing the dependability of UAV-based systems in networking and computing contexts. After this filtering, we selected 458 articles for in-depth analysis, forming the foundation of this survey.
To guide our analysis, we formulated the following research questions:
How have research trends on the dependability of UAV-based systems evolved in recent years?
What types of dependability threats are considered in the literature on UAV-based networks and computing systems?
What techniques are adopted to improve the dependability of UAV-based networks and computing systems?
What are the future research directions for dependable UAV-based networks and computing systems?
These questions provide a structured framework to explore the multifaceted aspects of dependability of UAV-based systems, from identifying vulnerabilities to evaluating solutions and envisioning future advancements.
The rest of the survey is organized as follows. We begin by reviewing related work in Section 2 to contextualize the evolution of UAV-based systems and their dependability requirements. Section 3 introduces the methodology used to conduct the research. We then explore research trends in Section 4, including systematic taxonomies, categorization of dependability aspects, and comparative analyses of metrics, system types, threats, and methodological approaches. In Section 5, a detailed examination of dependability threats highlights the multifaceted nature of the problem. To address these challenges, we survey methodological approaches in Section 6. Then, we discuss future research directions in Section 7 and threats to validity in Section 8, offering insights into the evolving landscape of dependability of UAV-based systems. Finally, we summarize our work in Section 9.
2. Related work
A growing body of surveys has explored various aspects of UAV-based computing and networking.
UAV-Enabled Edge and Distributed Computing.
Recent surveys have focused on UAV-enabled edge and distributed computing systems. For example, Xia et al. (Xia et al., 2023) presented a granular classification of resource management paradigms, such as computation offloading, energy-aware scheduling, and hybrid edge-cloud collaboration. Huda et al. (Huda and Moh, 2022) further compared model-based optimization and learning-driven methods in offloading decision-making frameworks, highlighting algorithmic trade-offs in responsiveness, scalability, and adaptability. Adil et al. (Adil et al., 2024) examined UAV-assisted IoT applications, outlining QoS requirements and communication challenges that directly impact system reliability. These works highlight the importance of efficient resource management and robust communication for dependable UAV operations, yet they often treat these aspects in isolation.
AI Techniques in UAV Systems.
The intersection of AI and UAVs has also garnered significant attention. Ning et al. (Ning et al., 2023) mapped ML techniques (e.g., federated learning, reinforcement learning) to UAV applications like autonomous navigation and swarm coordination. Heidari et al. (Heidari et al., 2023) explored the role of machine learning in enhancing UAV autonomy and decision-making, while also addressing energy efficiency through offloading techniques. These surveys underscore the potential of AI to mitigate dependability challenges, but they rarely connect these advancements to broader dependability frameworks that encompass both networking and computing.
Energy and Sustainability Considerations.
Sustainable operation remains a pressing challenge for UAV-based systems. Ahmed et al. (Ahmed et al., 2025) emphasized energy sustainability, addressing optimization in power consumption and flight planning. Cyrille et al. (Cyrille et al., 2024) categorized energy management strategies across hardware optimization and renewable energy harvesting. They highlight the need for integrated energy solutions that balance computational and networking demands, an area that intersects with dependability but is often studied independently.
Security and Privacy in UAV Networks.
Additionally, surveys on cyber-physical security and privacy, such as those discussing secure communication and data protection in UAV networks, highlight the intersection of dependability and security, a critical concern in safety-critical applications. Zhi et al. (Zhi et al., 2020) provided an overview of UAV security threats across sensors, communication links, and multi-UAV coordination, while also addressing privacy concerns such as unauthorized surveillance and data leakage. Mekdad et al. (Mekdad et al., 2023) analyzed vulnerabilities in both civilian and military UAV applications, highlighting challenges in secure transmission and adversarial defense. These studies are complementary to our study, as security and privacy are also essential properties when systems are threatened by malicious attackers. In this paper, we limit our scope to dependability (i.e., reliability, availability, maintainability, and safety) and do not cover security and privacy-related issues.
Survey Gaps and Motivation.
While these surveys provide valuable insights into specific topics of UAV computing and networking, they focus on isolated aspects, such as communication, computation, or security, without a holistic examination of dependability across both networking and computing domains. This fragmentation highlights the need for a comprehensive survey, such as the one proposed in this work, that systematically integrates these perspectives to address the multifaceted nature of dependability of UAV-based systems.
3. Methodology
Our systematic mapping was guided by the methodology proposed by Petersen et al.(Petersen et al., 2008), which we employed to identify works related to UAV dependability under different computing systems and approaches.
3.1. Planning the Review
To ensure systematic coverage, we formulated a set of focused research questions (RQs) as follows.
How have research trends on the dependability of UAV-based systems evolved in recent years?
What types of dependability threats are considered in the literature on UAV-based networks and computing systems?
What techniques are adopted to improve the dependability of UAV-based networks and computing systems?
What are the future research directions for dependable UAV-based networks and computing systems?
Based on these RQs, we developed a comprehensive search string encompassing three key conceptual components: UAV terminology, computing and networking domains, and dependability attributes. The final Boolean search string was designed as follows:
("UAV" OR "UAS" OR "drone" OR "quadcopter" OR "aircraft") AND ("computing system" OR "network system" OR "edge computing" OR "fog computing" OR "cloud computing" OR "distributed system") AND ("dependability" OR "reliability" OR "availability" OR "safety" OR "fault-tolerance" OR "resiliency" OR "resilience" OR "survivability" OR "robustness" OR "performability")
We selected five major digital libraries as data sources: IEEE Xplore, ACM Digital Library, SpringerLink, ScienceDirect, and ISI Web of Science, and selected papers published between 2015 and 2024.
3.2. Data Collection and Extraction
Based on the selected databases, we first conducted an initial screening of articles using our structured search queries. This preliminary step aimed to retrieve a broad but relevant set of publications related to our survey scope. Subsequently, we applied the predefined inclusion and exclusion criteria to refine the pool of studies. The inclusion criteria focused on primary research that introduced novel methods, frameworks, or experimental insights addressing the dependability of UAV-based systems. We excluded studies primarily focused on cybersecurity, hardware-level reliability, or other secondary literature, such as review and survey papers, to ensure alignment with our scope.
The filtering process was conducted in two stages by multiple researchers: first, by reviewing titles and abstracts, and then by performing full-text screening. Any disagreements regarding article inclusion were resolved through consensus discussions to ensure consistency. For the final set of selected studies, we extracted detailed information relevant to our research questions, including addressed dependability threats, adopted metrics and methodologies, and application scenarios.
3.3. Classification, Synthesis, and Reporting
After data extraction, we systematically categorized the selected articles according to the views defined in our research questions. These aspects include system types (e.g., edge-connected UAVs), threat types (e.g., network disconnection), methodological approaches (e.g., modeling, optimization), and applications (e.g., monitoring, rescue). The general research trends on these keywords are presented in Section 4.
Then, we synthesized the responses to research questions in Sections 5 and Section 6, identifying dependability threats covered in the literature and dependability techniques to counteract these threats. Finally, based on the findings from comprehensive literature reviews, we discuss possible future research directions to answer RQ4 in Section 7.
4. Research trends
To provide a comprehensive understanding of the dependability of UAV-based systems, we begin with an overview of the primary studies conducted in this field. Table 1 shows the number of studies identified by the search strings across selected databases, as well as the number of studies retained after applying the defined inclusion and exclusion criteria. Initially, a total of 1,848 articles were retrieved from five databases. After removing duplicates and applying the selection criteria, 458 articles were retained for full-text analysis and used to address the research questions.
| Source | Selected | Accepted |
| ACM Digital Library | 19 | 7 |
| IEEE Digital Library | 167 | 111 |
| ISI Web of Science | 376 | 217 |
| Science@Direct | 582 | 72 |
| Springer Link | 704 | 51 |
In particular, we survey the key papers published from 2015 to July 2024, which amount to 458 papers after filtering. These include journal articles, conference proceedings, and some articles in open-source archives. The analysis of publication trends over the years reveals a consistent and accelerating growth in research on dependability of UAV-based systems, as shown in Figure 1. The field began modestly, with only three publications in both 2015 and 2016. A noticeable increase occurred from 2018 onward, with 25 papers published that year. This growth continued steadily: 29 papers in 2019, 41 in 2020, and a sharp rise to 73 papers in 2021. These figures reflect the growing interest in UAV reliability, driven by technological advances and expanding application domains. In recent years, the trend has intensified. In 2023 alone, 101 papers were published, followed by 109 publications in the first seven months of 2024. This numbers shows the increasing relevance of UAV dependability, reflecting their broader adoption in sectors such as logistics, surveillance, and communications. This steady growth indicates a maturing and dynamic research area, marked by strong interdisciplinary contributions.
4.1. Research trends
To systematically analyze the research trends in the dependability of UAV-based systems, we closely examined the keyword frequencies and their temporal variations over the investigation period.
4.1.1. Keyward frequency analysis
First, we examined the total number of occurrences over the ten years for selected keywords in various subjects. The comparison of keyword frequencies highlights the hot areas discussed in the existing literature, as shown in Figure 2. In terms of dependability metrics shown in Figure 2(a), the most prominent emphasis is placed on reliability (380 mentions) and performance (364 mentions), exhibiting their foundational role in ensuring consistent UAV operations. Energy (258 mentions) and availability (195 mentions) also receive considerable attention, reflecting concerns around power constraints and operational readiness. Less frequently discussed metrics, such as robustness, fault tolerance, and especially safety and maintainability, highlight potential opportunities for future research in enhancing system resilience and post-deployment usability.
Regarding system types, as presented in Figure 2(b), edge-connected UAVs (344 mentions) and multi-drone systems (292 mentions) dominate the landscape, illustrating the growing importance of distributed, collaborative UAV networks. Single-drone systems, although less prevalent, remain relevant for specific applications, while cloud-connected architectures appear to be less emphasized, possibly due to latency or connectivity challenges. When analyzing dependability threats shown in Figure 2(c), the primary concern is network disconnection (330 mentions), reflecting the critical role of stable communication links in UAV-based systems. Hardware faults and software faults are also recognized as significant challenges, while the substantial frequency of the ”other” category suggests a wide array of additional threats, such as environmental or operational factors.
In the realm of methodological approaches, as presented in Figure 2(d), optimization algorithms (300 mentions) and modeling & simulation (299 mentions) are the dominant techniques employed to enhance UAV system performance and predictability. Other commonly used methods include resource allocation, task assignment, and offloading strategies. The notable application of machine learning, game theory, and detection and diagnosis techniques demonstrates an increasing trend toward intelligent, adaptive UAV systems. Finally, the application domain analysis in Figure 2(e) reveals that monitoring (246 mentions) and surveillance (187 mentions) are the primary use cases for UAVs, emphasizing their critical role in security, environmental observation, and emergency response scenarios. Emerging applications such as rescue missions, maintenance, and delivery services appear less frequently, indicating potential areas for future expansion and innovation.
4.1.2. Major research trends
To capture the evolving focus in different aspects, we conducted a temporal analysis of keyword frequencies over the years (see Figure 3). By generating a year-wise pivot graphics, we systematically tracked how often each metric was mentioned in the literature from 2015 to 2024. This approach enables us to observe shifts in research priorities and emerging areas of interest over time.
The yearly distribution of dependability metrics reveals several notable trends, as shown in Figure 3(a). Reliability and performance consistently dominate the landscape, with mentions increasing from fewer than five in early years to over 90 and 87, respectively, by 2024. This indicates a sustained emphasis on ensuring the operational consistency and efficiency of UAV systems. Energy-related terms also demonstrate significant growth, reflecting increasing attention to power constraints and energy optimization, particularly as UAV missions become more complex. The frequency of availability and robustness similarly shows steady growth, highlighting a growing interest in system readiness and resilience to disruptions. Notably, mentions of fault-tolerance and safety also rise over the years, though at a comparatively moderate pace. Maintainability appears infrequently, suggesting it remains an underexplored aspect in the current research landscape. Overall, the results illustrate a maturing research focus, where foundational concerns, such as reliability and performance, remain important, while emerging challenges, like energy efficiency, fault handling, and safety, are receiving increased attention as UAV systems are deployed in more demanding and diverse environments.
The temporal distribution of system types reveals notable shifts in focus over the years, as shown in Figure 3(b). The most prominent trend is the steady increase in the attention to edge-connected UAV systems, with mentions rising from only one in 2015 to a peak of 80 by 2024. This indicates the growing importance of edge computing frameworks in supporting real-time processing, reducing latency, and enhancing decision-making capabilities, particularly as UAV applications become increasingly data-intensive. Similarly, research on multi-drone systems has seen a significant rise, increasing from a few mentions in the early years to 71 mentions by 2024. The prominence of multi-drone studies reflects a clear emphasis on scalability, redundancy, and cooperative task execution, which are essential for enhancing the dependability of UAV systems in complex environments. In contrast, single-drone systems maintain a relatively stable but lower frequency, suggesting their continued relevance for focused, cost-effective operations where simpler setups suffice. Interestingly, cloud-connected UAVs show fluctuating yet consistently lower mentions throughout the years, peaking slightly at ten mentions in 2023. This may be attributed to the challenges associated with cloud dependence, such as latency and connectivity reliability, making edge-based solutions more attractive for mission-critical UAV tasks. Lastly, mentions categorized under other system types remain modest, reflecting niche configurations or experimental system architectures tailored to specific use cases. Overall, the results highlight a shift in research toward distributed, collaborative, and real-time capable UAV systems, aligning with broader trends observed in emerging UAV applications and technological advancements.
The yearly distribution of identified dependability threats in UAV systems highlights key vulnerabilities that have attracted sustained research attention, as shown in Figure 3(c). The most frequently addressed threat is network disconnection, which shows a significant upward trend, increasing from only a few mentions in 2015-2016 to 78 mentions in 2024. This highlights the crucial importance of reliable communication links in UAV networks, particularly as these systems increasingly rely on real-time data exchange and edge or cloud connectivity. Similarly, concerns related to hardware faults have steadily increased, rising from occasional mentions in early years to a notable 47 mentions in 2023. This trend highlights the inherent susceptibility of UAV mechanical and electronic components to wear, environmental stress, and operational failures, particularly in demanding or prolonged missions. The software fault category, while receiving fewer mentions overall, exhibits a stable pattern across the years. The relatively lower frequency may suggest that software reliability issues, though important, are perceived as more controllable or well-understood compared to physical and network-related threats. The high and increasing counts in the other threats category—peaking at 72 mentions in 2023—indicate the diverse range of additional challenges faced by UAV systems, such as environmental factors, cyber-attacks, operational errors, and unpredictable mission conditions. This wide range underscores the complexity of ensuring dependability of UAV-based systems in real-world scenarios. In summary, the threat analysis reveals that network reliability and hardware robustness remain the primary concerns, while attention to software vulnerabilities and emerging threats continues to grow, suggesting an ongoing need for holistic and multi-layered dependability solutions.
The temporal analysis of dependability techniques reveals evolving preferences in the tools and techniques employed to enhance system performance and reliability, as shown in Figure 3(d). Modeling and Simulation consistently dominates throughout the years, showing a strong and steady increase from three mentions in 2015 to 66 mentions in 2024. This highlights the importance of simulation-based approaches for testing UAV systems in diverse scenarios, allowing researchers to assess dependability without incurring real-world deployment risks. Optimization Algorithms similarly exhibit significant growth, with mentions rising from two in 2015 to 75 in 2024. This trend underscores the pivotal role of optimization techniques in enhancing UAV efficiency, resource management, and task scheduling. Machine Learning and AI methods display an exponential rise—from a minimal presence in early years to 32 mentions in 2024—indicating increasing integration of intelligent, adaptive systems to enhance UAV autonomy, fault prediction, and decision-making capabilities. Other noteworthy techniques include offloading strategies, which show a steady increase in mentions, emphasizing the need to manage computational loads effectively through edge or cloud resources. Mentions of detection and diagnosis techniques, though lower in frequency not shown in the picture, show gradual growth with single-digit, reflecting growing interest in real-time anomaly detection and system health monitoring. Interestingly, the application of Game Theory, while still emerging, with only five papers mentioned until 2023, has seen a noticeable increase in the last few years, suggesting its growing relevance in strategic decision-making, coordination, and resource negotiation within multi-UAV environments. Overall, the methodology trends indicate a progressive shift from traditional simulation-based validation towards the integration of AI-driven, optimization-centric, and distributed approaches, reflecting the complex operational requirements of modern UAV systems.
The temporal distribution of UAV application domains reveals clear priorities and emerging trends in the research landscape, as shown in Figure 3(e). Among the various application areas, monitoring and surveillance consistently dominate, with mentions increasing significantly over the years. Monitoring-related terms rose from just two mentions in 2015 to a peak of 59 mentions in 2023, while surveillance mentions grew from two in 2015 to 47 in 2023. This reflects the critical role of UAVs in environmental observation, security, and large-scale monitoring tasks, where dependability and operational readiness are paramount. The general application category also shows steady growth, increasing from minimal mentions in early years to 42 mentions by 2024. This indicates the versatility and expanding deployment of UAVs across diverse sectors, including those not confined to specific use cases. In contrast, mentions of rescue missions and maintenance tasks, though lower in absolute numbers, show gradual increases, highlighting the growing recognition of UAVs in safety-critical and infrastructure support roles. Rescue missions, for instance, reached 11 mentions by 2021, reflecting UAVs’ utility in emergency response scenarios. Interestingly, while delivery applications appear less frequently overall, there is a noticeable rise from 0 mentions in early years to occasional mentions in recent years. This suggests an emerging interest in UAV-based logistics, though it remains a relatively underexplored domain compared to surveillance and monitoring. The mentions categorized under other applications remain moderate, pointing to niche or evolving use cases tailored to specific operational needs. In summary, the application analysis highlights the dominance of monitoring and surveillance, while also underscoring the growing opportunities in rescue, maintenance, and delivery domains that warrant further exploration.
5. Dependability threats
As UAVs take on increasingly mission-critical roles, ensuring their dependability becomes essential. This section examines core dependability challenges such as reliability, availability, integrity, and maintainability that directly affect UAV performance under both nominal and adverse conditions. Failures in these aspects can result in mission disruption, data loss, physical damage, or compromised safety, especially in dynamic or high-risk environments. Unlike traditional computing systems, UAVs operate in highly dynamic and often uncontrolled environments. They rely on the seamless integration of multiple subsystems such as sensors, actuators, power modules, communication networks, and embedded software to function properly. A failure in any one of these subsystems can propagate and result in system-wide consequences. Moreover, UAVs are subject to external conditions like weather, terrain complexity, and electromagnetic interference, which further exacerbate the risk landscape. To provide a structured understanding of these challenges, Figure 4 presents a comprehensive taxonomy of dependability threats in UAV systems. This taxonomy categorizes threats into seven major domains: (1) Energy Constraint, involving battery degradation, excessive power draw, and charging limitations; (2) Network Issues, such as signal loss, high latency, bandwidth limitations, congestion, and RF interference; (3) Hardware Failure, covering sensor malfunctions, actuator faults, aging components, and structural breakdowns; (4) Software Failure, including bugs, vulnerabilities, and system crashes; (5) Environmental Impact, focusing on adverse weather and terrain-related difficulties; and (6) Operational Failures, which stem from human error and maintenance oversights.
By systematically identifying and organizing these threats, the taxonomy in Figure 4 serves as a foundational reference for UAV designers, operators, and researchers. It supports more targeted risk assessments, improves fault tolerance planning, and guides the development of resilient UAV systems capable of maintaining dependable operation in real-world deployments. Table 2 complements this taxonomy by categorizing the surveyed papers according to the identified dependability threat types in UAV-based systems, providing further evidence from the literature.
| Category | Subcategory | Related Studies |
| Energy Constraint | Battery Degradation | (Lu et al., 2022), (Hou et al., 2020b), (Yang et al., 2022), (Li and Nie, 2023) |
| Power Consumption | (Falcão et al., 2024), (Kumar and Jannesari, 2023), (Alqudsi et al., 2023) | |
| Charging Challenges | (Liang et al., 2021), (Huang et al., 2019), (Hadi and Ghazizadeh, 2023), (Kabashkin, 2023) | |
| Network Issues | Signal Loss | (Zhang et al., 2022b), (Kramer and Stubberud, 2021), (Li et al., 2020b), (Li and Nie, 2023) |
| High Latency | (Alam et al., 2019), (Feng et al., 2023), (Tang et al., 2023), (Hou et al., 2020a), (Yu et al., 2021), (Tang et al., 2022b) | |
| Data Packet Loss | (Mukherjee et al., 2023), (Wang et al., 2019b), (Nwadiugwu et al., 2021), (Jiang et al., 2019), (Haber et al., 2019), (Bai et al., 2023a), (Zhang et al., 2023c), (Chu et al., 2021), (Mukherjee et al., 2020) | |
| Bandwidth Constraints | (Ali and Ali, 2022), (Zhang et al., 2021a), (Wang et al., 2023b), (Tareq et al., 2018) | |
| Network Congestion | (Picano and Fantacci, 2024), (Nwadiugwu et al., 2021), (Picano and Fantacci, 2024), (She and Ouyang, 2021) | |
| Channel Interference | (Shen et al., 2023b), (Wang et al., 2023a), (Sun et al., 2018), (Zhang et al., 2024d), (Li et al., 2020b) | |
| Hardware Failure | Sensor Malfunction | (Saadaoui et al., 2023), (Hou et al., 2020a), (Shen et al., 2023b), (Ju et al., 2024), (Huang et al., 2024a), (Yuan et al., 2022), (Chen et al., 2020b), (Kabashkin, 2023) |
| Actuator Failure | (Amin and Obermaisser, 2018), (Zhou and Cao, 2018), (Wang et al., 2019b), (Wang et al., 2021d), (Hirota et al., 2020), (Hu and Seiler, 2015) | |
| Component Aging | (Shao et al., 2023), (Aloqaily et al., 2021), (Haber et al., 2019), (Dogea et al., 2023), (Ren et al., 2017), (Guo et al., 2022a), (Zheng et al., 2023), (Wang et al., 2020c), (D’Agati et al., 2024), (Moeyersons et al., 2021), (Xiong et al., 2019), (Li et al., 2019), (Navardi and Mohsenin, 2023), (Zhang, 2020), (Jiang and He, 2023), (Zheng et al., 2021), (Faraci et al., 2023), (Liu et al., 2024c), (Seisa et al., 2022), (Luo et al., 2021), (Haber et al., 2019), (Bao et al., 2023), (Jiang et al., 2019), (Falcão et al., 2023), (Bai et al., 2023b), (Jung et al., 2023a), (Ngoenriang et al., 2022), (Feng et al., 2024), (Geng et al., 2024), (Ning et al., 2024), (De Paula Soares et al., 2023), (Ren et al., 2024), (He et al., 2023), (Gupta and Gupta, 2022), (Chen et al., 2024a), (Picano and Fantacci, 2024), (Wang et al., 2021a),(Li et al., 2023b), (Feng et al., 2024), (Khoramnejad et al., 2024), (Zhou et al., 2024b), (Lei et al., 2023), (Shao et al., 2023), (Hou et al., 2020b), (Yu et al., 2023a), (Zhang et al., 2023c), (Tripathi et al., 2023), (Lorincz et al., 2021), (Pacheco et al., 2021), (Patnayak et al., 2021), (Pamuklu et al., 2023), (Kaleem et al., 2019), (Zeng et al., 2024b), (Adil et al., 2024), (Hang et al., 2020), (Wang et al., 2022c), (Li et al., 2024a) | |
| Mechanical Breakdown | (Luo et al., 2021), (Zhang et al., 2021a), (Russell et al., 2018), (Shen et al., 2023b), (Lv et al., 2024), (Wang et al., 2020b), (Dai et al., 2024), (Choutri et al., 2022), (Zhou et al., 2024a), (Lindqvist et al., 2020), (Kabashkin, 2023), (Lorincz et al., 2021), (Behjat et al., 2019) | |
| Software Failure | System Vulnerabilities | (Pradhan et al., 2015), (Wang et al., 2023a), (Ma et al., 2024), (Vamvakas et al., 2019), (Wang et al., 2023a), (Ren et al., 2017), (Venkatagiri et al., 2018), (Qu et al., 2022), (Chang et al., 2024), (Watanabe and Machida, 2022), (Zhong et al., 2021), (Zuo et al., 2022), (Zhou et al., 2023), (Chen et al., 2022a), (Yang et al., 2019), (Hirota et al., 2020) |
| System Crashes | (Ma et al., 2024), (Amin and Obermaisser, 2018), (Hirota et al., 2020), (Kramer and Stubberud, 2021), (Liu et al., 2019), (Qu et al., 2022), (Chang et al., 2024), (Pan et al., 2024), (Jennings et al., 2022), (Shang et al., 2023), (Taleb et al., 2023), (Zhang et al., 2021c), (Grasso et al., 2022b), (Ritchie et al., 2020), (Lin et al., 2021), (Mo et al., 2022), (Gupta et al., 2021), (Hong et al., 2022), (Zhou et al., 2024c), (Zhong et al., 2021), (Nguyen et al., 2023), (Zhang and Li, 2018), (Kaymaz et al., 2024), (Ahamad and Mishra, 2024), (Goel et al., 2021), (Bai et al., 2023a) | |
| Environmental Impact | Adverse Weather Conditions | (Wang et al., 2021b), (Wang et al., 2021b), (Gallego-Madrid et al., 2020), (Hossain et al., 2017), (Pradhan et al., 2015), (Haber et al., 2021), (Yang et al., 2022), (Lee et al., 2022a), (Sankarasrinivasan et al., 2015), (Barnawi et al., 2021), (Mcilwaine and Rivas Casado, 2021), (Yang et al., 2022), (Fan et al., 2023), (Venkatagiri et al., 2018), (Kumar and Jannesari, 2023) |
| Terrain Challenges | (Dai et al., 2018), (Rokaha et al., 2019) , (Paranjothi et al., 2019), (Bakhtiari et al., 2020), (Gao et al., 2022), (Peng et al., 2023), (Dai et al., 2018), (Kim et al., 2024b), (Faraci et al., 2023), (Mao et al., 2023), (Shang et al., 2023), (Lv et al., 2024), (Wang et al., 2021d), (Hadi and Ghazizadeh, 2023), (Zhou et al., 2024a) | |
| Operational Failures | Human Error | (Xu and Zhai, 2020), (Hou et al., 2020a), (Yu et al., 2021), (Liu et al., 2023c), (Yuan et al., 2022), (Falcão et al., 2023), (Kim et al., 2024b), (Mao et al., 2023), (Faraci et al., 2023), (Thapliyal et al., 2024), (Zhang et al., 2024d), (Xu and Zhai, 2020), (Manzoor et al., 2021), (Jeong et al., 2020), (He et al., 2023), (Hui et al., 2024), (Falcão et al., 2023), (Li et al., 2024e), (Kaleem et al., 2019), (Ozger et al., 2018), (Rao and Naidu, 2024), (Ouyang et al., 2023), (Ju et al., 2024), (Tang et al., 2023), (Amos et al., 2021), (Shen et al., 2023a), (Wang et al., 2020a), (Vamvakas et al., 2019), (Asim et al., 2022) |
| Maintenance Issues | (Guo et al., 2022b), (Jo et al., 2018), (Wu et al., 2024), (Lv et al., 2024), (Yang et al., 2017), (Ren et al., 2017), (Zhang, 2020), (Pradhan et al., 2015), (Lin et al., 2021), (Zhang and Zhang, 2024), (Patnayak et al., 2021), (Shams et al., 2023), (Huang and Wang, 2024) |
5.1. Energy Constraint
Energy constraints critically affect UAV operations, which rely on limited onboard power. The main threats related to this category are presented below, highlighting challenges to ensure safe and efficient UAV functionality.
Battery Degradation: Battery degradation refers to the gradual loss of battery capacity and performance due to repeated charging and discharging cycles. In UAVs, lithium-polymer (LiPo) batteries are the most commonly used due to their high energy density (Lu et al., 2022). However, over time, these batteries undergo chemical aging, reducing the amount of charge they can hold (Hou et al., 2020b). This results in shorter flight durations (Yang et al., 2022), inconsistent power output, and potential mid-flight failures (Li and Nie, 2023). Factors that accelerate degradation include overcharging, deep discharges, and exposure to high temperatures. For instance, a UAV that initially offers 30 minutes of flight time may reduce to just 15 minutes after 100 charge cycles. Real-time monitoring and battery health management systems can help detect and mitigate degradation risks.
Power Consumption: Power consumption in UAVs is influenced by multiple subsystems, including propulsion, communication, payload sensors, and onboard computing. Propulsion, particularly in multi-rotor systems, is the largest power consumer, especially during aggressive maneuvers or hovering. High-definition video transmission and real-time processing for autonomous navigation also draw significant energy. Managing power consumption requires efficient hardware selection, software optimization, and flight path planning (Falcão et al., 2024). For example, flying at a steady altitude with minimal changes in direction can conserve power, while rapid acceleration and deceleration will deplete the battery faster. In addition, virtualization overhead in MEC-enabled UAVs increases computational demands, leading to higher power consumption, which challenges the ability to maintain node availability and fulfill URLLC (ultra-reliable low-latency communication) requirements (Falcão et al., 2024). Environmental factors, such as strong wind disturbances, also force UAVs to consume additional energy for stabilization and trajectory correction, further impacting overall efficiency (Kumar and Jannesari, 2023). To address these issues, recent studies propose two energy-aware optimization schemes: one that minimizes flight cycles by optimizing sensor wake-up schedules, UAV trajectories, and time slots, and another that maximizes energy efficiency through the joint optimization of data transmission and propulsion power consumption (Alqudsi et al., 2023).
Charging Challenges: Charging challenges refer to the difficulties in quickly and safely recharging UAV batteries in operational environments. In field conditions, access to stable power sources can be limited, and charging LiPo batteries requires specialized chargers with voltage and current regulation. Charging times are often long, reducing operational readiness (Liang et al., 2021). Furthermore, improper charging can lead to overheating or battery swelling, posing safety hazards. Fast-charging solutions and battery swapping systems are being developed to reduce downtime, but they come with added weight and logistical complexity. For instance, in remote surveillance missions, UAV teams may need to carry generators or solar panels for recharging, adding to operational overhead. Recent studies have proposed optimizing UAV positions by considering recharging point accessibility and building avoidance to support energy management in real-world deployments (Huang et al., 2019). A recent framework integrates intelligent reflecting surfaces (IRS), UAVs, and laser-based wireless charging to enhance energy harvesting efficiency, offering a new approach to reduce charging delays and support continuous UAV operation (Hadi and Ghazizadeh, 2023). Additionally, service continuity challenges are tackled by modeling UAV replacement during charging periods to maintain network availability (Kabashkin, 2023).
5.2. Network Issues
UAVs require stable and secure communication links. Below, we present the main threats in this category, which impact performance, safety, and mission reliability.
Signal Loss: Signal loss is a critical issue for remotely piloted or semi-autonomous UAVs, resulting in the interruption of control commands and telemetry feedback. Causes include exceeding the communication range (Zhang et al., 2022b), line-of-sight obstructions (e.g., buildings, hills), antenna misalignment, hardware malfunctions, and network disconnection due to unstable communication links (Kramer and Stubberud, 2021). When signal loss occurs, most UAVs trigger fail-safe behaviors such as returning to home or hovering in place . However, in GPS-denied environments or during complex missions, signal loss can result in mission failure or crashes. For example, a drone flying behind a tall building may lose the link with the operator and enter an unintended hover, exhausting its battery and eventually crashing. Recent studies have also highlighted that high Doppler shifts in UAV communication can cause signal fading and interference, degrading communication reliability. This issue is especially critical in high-mobility scenarios and may impact applications including but not limited to mobile edge computing (MEC) (Li et al., 2020b). Additionally, network disconnections between UAVs and control centers, often resulting from unstable links or system faults, further contribute to signal interruptions during flight missions (Li and Nie, 2023).
High Latency: High latency refers to delays between command transmission and UAV response, which is especially problematic in real-time operations like inspection or FPV (First-Person View) control. Latency can be caused by long-range communication links (e.g., satellite), network congestion, or processing delays. In addition, limited onboard processing capacity often forces UAVs to offload computational tasks to the cloud, introducing further communication delays and increasing latency (Alam et al., 2019). Conventional UAV image processing methods, which involve merging images and the use of ground control points (GCPs), can also introduce high latency and positioning errors if not managed carefully (Feng et al., 2023). Moreover, the integration of UAVs and electric vehicles (EVs) in edge computing environments faces stringent latency challenges, especially under dynamic workloads and varying task demands from users (Tang et al., 2023). Network disconnections between UAVs and user equipment have also been identified as a cause of increased latency in computational offloading, further impacting system reliability (Hou et al., 2020a). Additionally, intermittent connectivity and complex route optimization in UAV-enabled MEC systems contribute to high latency under dynamic network conditions (Yu et al., 2021; Tang et al., 2022b). For instance, a 300 ms delay might not affect automated mapping missions but can make manual control erratic and dangerous. Reducing latency involves optimizing data routing, using low-latency protocols, and deploying edge-computing systems onboard the UAV.
Data Packet Loss: Packet loss occurs when one or more transmitted data packets fail to reach their destination, leading to gaps in communication. In UAVs, this can result in missing telemetry, sensor data, or delayed command execution. Causes include RF interference, poor signal quality, and buffer overflow in the UAV or ground station. In high-mobility networks such as FANETs (Flying Ad Hoc Networks), the likelihood of packet loss and intermittent connections increases due to dynamic topology changes and unstable links (Mukherjee et al., 2023). Additionally, quadrotor UAV systems are vulnerable to network issues such as random delays and packet loss, which can destabilize real-time control and disrupt reliable wireless communication (Wang et al., 2019b). Enhancements at the MAC protocol level, such as prioritizing nodes based on proximity, have been proposed to reduce network delay and improve robustness against packet loss (Nwadiugwu et al., 2021). Poor packet scheduling and resource allocation in multi-agent UAV systems can lead to higher packet loss, especially when network conditions change (Jiang et al., 2019). Network disconnections between UAVs, ground stations, satellites, and IoT devices are also identified as major contributors to data loss and increased latency, impacting the reliability of computational offloading and communication services (Haber et al., 2019; Bai et al., 2023a; Zhang et al., 2023c; Chu et al., 2021). High packet loss can degrade mission accuracy. In IoDT applications, high message latency and packet loss can reduce quality of service (QoS), especially in dynamic environments (Mukherjee et al., 2020). Limited energy and memory in edge devices make these challenges worse. For example, a UAV sending image data back to a ground station may deliver incomplete or corrupted files, requiring retransmission and causing delays. Protocols with error correction and adaptive retransmission strategies help mitigate these issues.
Bandwidth Constraints: Bandwidth limitations affect the volume and speed of data transmission between the UAV and ground control. High-bandwidth applications like HD video streaming or large-scale sensor data collection require stable and fast connections. In multi-UAV operations or urban environments, limited bandwidth can lead to reduced data quality or interrupted connections (Ali and Ali, 2022; Wang et al., 2023b). For example, a surveillance drone attempting to stream 4K footage over a congested 2.4GHz channel may suffer frame drops or lag. To address bandwidth challenges, frameworks like Fed-UAV implement federated learning on edge devices, reducing the amount of data transmitted to central servers and helping to minimize bandwidth usage (Zhang et al., 2021a). Another method to reduce bandwidth problems is to use a normalized link throughput strategy, which helps lower interference and improve data transmission (Ali and Ali, 2022). SDN-based approaches are also used to handle network disconnections. Managing limited SBS resources and bandwidth helps ensure fast, reliable communication for AVs without relying too much on the cloud (Tareq et al., 2018).
Network Congestion: Network congestion happens when multiple UAVs or devices compete for the same communication resources, causing delays and data collisions. This is common in swarm operations, public events, or urban areas. Congestion leads to increased latency, packet loss, and sometimes total communication failure (Picano and Fantacci, 2024). Lack of effective node prioritization in UAV networks can worsen congestion by increasing competition for limited communication channels (Nwadiugwu et al., 2021). Heavy traffic and strict QoS demands on UAV and ground edge nodes can cause congestion, delay communications, and lead to missed deadlines for important tasks (Picano and Fantacci, 2024; She and Ouyang, 2021). For instance, in a drone light show involving dozens of UAVs, congestion could cause synchronization errors, impacting performance and safety. Solutions involve dynamic channel allocation, time-division multiplexing, and efficient communication protocols that adapt to network load.
Channel Interference: Channel interference refers to the disruption of UAV communication by overlapping signals from other RF sources such as Wi-Fi routers, mobile phones, or other drones. Interference reduces signal clarity, increases error rates, and can cause unexpected disconnections (Shen et al., 2023b). Deliberate channel access attacks (CAA) and fast-varying channel conditions in dynamic environments, such as vehicular networks, both contribute to communication interference and instability in UAV networks (Wang et al., 2023a; Sun et al., 2018; Li et al., 2020b). Channel uncertainty caused by dynamic air-ground communication in UAV networks also increases the risk of interference and affects the reliability of maintaining continuous connections, particularly for applications like the Metaverse (Zhang et al., 2024d). UAVs operating in crowded RF environments need robust modulation techniques and interference avoidance strategies. For example, a UAV inspecting infrastructure near a busy telecom tower may suffer from sporadic link quality due to interference. Frequency hopping and spectrum sensing technologies are commonly employed to reduce this risk.
5.3. Hardware failure
UAVs rely on reliable physical components to maintain stable and safe operation. Below, we present the main threats in this category, which impact flight control, system durability, and mission success
Sensor Malfunction: Sensor malfunctions can occur in components such as GPS, IMU, altimeters, or cameras. These devices are critical for navigation, stabilization, and mission-specific tasks. Faulty or drifting sensors provide inaccurate data to the flight controller, leading to instability or incorrect responses (Saadaoui et al., 2023). For example, a GPS module reporting an incorrect position may cause the drone to veer off course, while an altimeter failure could result in improper altitude hold. Environmental factors such as magnetic interference or moisture can also contribute to sensor errors. UAV systems’ sensor malfunctions can disrupt critical services such as mobile edge computing, vehicular communications, or fog computing operations, impacting overall system reliability (Hou et al., 2020a; Shen et al., 2023b; Ju et al., 2024; Huang et al., 2024a; Yuan et al., 2022; Chen et al., 2020b). Environmental conditions, such as bad weather or collisions, further increase the risk of sensor malfunctions, requiring fault management strategies like deploying additional UAVs to replace malfunctioning ones and maintain service continuity (Kabashkin, 2023). To address these risks, system designs increasingly incorporate redundancy, regular calibration procedures, health monitoring, and fault-tolerant offloading strategies.
Actuator Failure: Actuators, such as servos and motor controllers, translate flight control commands into mechanical motion. Failure in these components can result in reduced maneuverability or total loss of control (Amin and Obermaisser, 2018). A jammed servo may lock a control surface on a fixed-wing UAV, while a faulty motor controller can shut down a rotor on a quadcopter. Main causes include overheating, mechanical wear, electrical faults, or broader hardware issues affecting actuators during flight operations (Zhou and Cao, 2018; Wang et al., 2019b; Amin and Obermaisser, 2018). Some systems address actuator failures by enabling rapid re-planning in response to failures, allowing continued operation through dynamic adjustments (Hirota et al., 2020). Low-cost UAVs are especially vulnerable due to limited actuator redundancy, making accurate fault detection and control switching essential; recent approaches use reliability frameworks with pivotal decomposition and new fault detection and isolation (FDI) metrics to handle actuator failures without relying on extra hardware (Hu and Seiler, 2015). Quadrotor UAV control systems are particularly vulnerable to actuator failures when combined with environmental disturbances and network issues like random delays and packet loss, which can threaten stability and disrupt real-time control (Wang et al., 2021d). Ensuring component quality, regular inspection, and incorporating fail-safes like motor redundancy in quadcopters can mitigate such risks.
Component Aging: With repeated usage, UAV components degrade due to mechanical fatigue (Shao et al., 2023), corrosion, and thermal cycling. Bearings in motors, solder joints on PCBs, and electrical connectors are particularly vulnerable. Aging components can reduce efficiency, increase energy consumption, or suddenly fail mid-flight. In addition, long-term stress and environmental exposure can gradually weaken critical UAV hardware, increasing the risk of unexpected failures and service disruptions (Aloqaily et al., 2021; Haber et al., 2019; Dogea et al., 2023; Ren et al., 2017; Guo et al., 2022a; Zheng et al., 2023; Wang et al., 2020c; D’Agati et al., 2024; Moeyersons et al., 2021; Xiong et al., 2019). Such aging-related failures, combined with network disconnections between aircraft and ground stations, may result in increased latency or data loss, impacting the reliability of data transmission and service continuity. Efficient energy management is also critical, as aging components may draw more power, reducing operational time. Moreover, aging systems can be more prone to software faults, where bugs or vulnerabilities emerge over time, requiring robust maintenance and development practices to uphold dependability (Ren et al., 2017).
Mechanical Breakdown: Structural components like frames, propellers, and landing gear are exposed to significant stress from takeoffs, landings, and potential collisions. Mechanical fatigue and wear over time can lead to cracks and fractures, particularly in components like landing gear and propeller mounts (Luo et al., 2021; Zhang et al., 2021a; Russell et al., 2018). Stress accumulation from repeated missions and vibrations contributes to deformation and weakening of airframes and connectors (Shen et al., 2023b; Lv et al., 2024; Wang et al., 2020b). Mid-air collisions or impacts during landing are common triggers for abrupt structural failures, especially in multi-UAV or urban operations where maneuvering space is limited (Dai et al., 2024; Choutri et al., 2022; Zhou et al., 2024a; Lindqvist et al., 2020). Additionally, poor material quality or inadequate structural design increases the risk of breakage under operational loads (Kabashkin, 2023; Lorincz et al., 2021; Behjat et al., 2019). A broken propeller mid-flight, for example, can result in immediate loss of control, particularly in systems lacking redundancy. Mechanical breakdowns are often preventable through visual inspections, structural stress testing, and the use of high-quality materials during design and assembly.
5.4. Software failure
UAVs depend on robust and secure software systems for navigation and control. Below, we present the main threats in this category, which affect system stability, security, and operational continuity.
System Vulnerabilities: System vulnerabilities in UAVs refer to weaknesses in the software architecture, coding logic, or configuration settings that can be exploited by attackers or cause unintended behavior. These vulnerabilities may include unvalidated inputs, buffer overflows, and insecure memory handling, which expose the system to risks like memory corruption and code injection (Pradhan et al., 2015; Wang et al., 2023a; Ma et al., 2024). UAVs operating in sensitive environments such as surveillance or defense are especially at risk. For example, attackers may exploit unpatched firmware or default access credentials to hijack UAV functions or inject malicious commands (Vamvakas et al., 2019). Some studies have also highlighted vulnerabilities in channel access mechanisms, where channel access attacks (CAA) can degrade the freshness and reliability of information by disrupting data transmission through protocol exploitation (Wang et al., 2023a). Additionally, system-level software faults, including bugs, logic errors, insecure configurations, integration issues within dynamic cyber-physical environments, or soft errors from environmental factors like radiation, can compromise functionality and security, disrupting service delivery and threatening mission continuity (Ren et al., 2017; Pradhan et al., 2015; Ma et al., 2024; Venkatagiri et al., 2018; Qu et al., 2022; Chang et al., 2024). Software aging is another critical concern, as prolonged operation may introduce subtle faults such as memory leaks or degraded performance, which can eventually lead to system failure or ineffective execution of key functions like task offloading (Watanabe and Machida, 2022). Furthermore, complex detection algorithms, such as those based on YOLOv5s, are vulnerable to configuration errors and mispredictions, which may result in high false alarm rates and reduced reliability of UAV-based detection systems (Zhong et al., 2021; Zuo et al., 2022; Zhou et al., 2023). Deep learning-based flight control systems may also encounter software faults, such as incorrect parameterization or model misbehavior in neural schedulers or DNNs, potentially leading to unsafe or inaccurate navigation decisions (Chen et al., 2022a). Monocular quadrotors also suffer from software limitations in depth estimation, where the absence of 3D information leads to unreliable obstacle avoidance, especially in unfamiliar or unstructured environments (Yang et al., 2019). Some UAV systems implement fault-tolerant mechanisms that enable rapid re-planning in response to actuator or sensor faults and allow seamless path switching through precomputed recovery scenarios in the event of unexpected software errors (Hirota et al., 2020). An attacker could also intercept unencrypted telemetry data and manipulate it to alter flight paths or collect confidential information. Regular software audits, secure coding practices, and real-time intrusion detection systems are essential to mitigate these risks.
System Crashes: UAV system crashes can occur due to software bugs, memory leaks, unexpected sensor inputs, or processing resources overload (Ma et al., 2024; Amin and Obermaisser, 2018; Hirota et al., 2020). These crashes may freeze the onboard computer, trigger automatic reboot, shut down critical flight functions or cause navigation-related faults (Kramer and Stubberud, 2021). For example, a software loop that does not terminate might consume all of the system memory, leading to a crash during flight. If recovery systems are not well implemented, the UAV may fall or crash uncontrollably. Edge-enabled UAVs are also vulnerable to system crashes if compute offloading mechanisms fail or if the drone is left without sufficient onboard fallback, potentially halting autonomous functions during critical missions (Liu et al., 2019). In multi-UAV edge orchestration frameworks, such as those used for collaborative video analytics, poorly optimized task scheduling or simultaneous high-load processing may overwhelm system resources, resulting in unresponsive drones or mid-mission system reboots (Qu et al., 2022). Misconfigured control logic at edge terminals, especially when not detected by testing, can result in severe flight instabilities or control system crashes, as demonstrated in fuzzing-based evaluations of UAV safety risks (Chang et al., 2024). Deep learning based detection modules like AIMED-Net, when deployed on resource-constrained edge platforms, may experience inference failures or overload under real-time constraints, potentially leading to unresponsive systems or control crashes during UAV operation. Comparative evaluations of onboard processors further highlight that under heavy computational loads, low-tier processors can become bottlenecks, increasing the risk of midmission system instability or crash (Pan et al., 2024; Jennings et al., 2022). Detection algorithms based on YOLOv5s can also suffer from configuration errors or inference instability, and in some cases, software crashes during execution, further increasing the risk of unresponsiveness in UAV detection systems(Shang et al., 2023; Taleb et al., 2023; Zhang et al., 2021c; Grasso et al., 2022b; Ritchie et al., 2020; Lin et al., 2021; Mo et al., 2022; Gupta et al., 2021; Hong et al., 2022). Other high-performance visual networks, such as VDTNet, may also introduce instability during edge-based detection and tracking tasks when deployed on resource-limited UAV platforms, potentially resulting in dropped detection, inference delays, or runtime crashes during intruder pursuit missions (Zhou et al., 2024c). System crashes can also stem from environmental hardware faults, energy depletion, unreliable communication links, or faults in AI-based detection and GPS/BDS-based attitude determination systems, which may cause incorrect orientation estimates and lead to UAV instability or crashes (Zhong et al., 2021; Nguyen et al., 2023; Zhang and Li, 2018; Kaymaz et al., 2024; Ahamad and Mishra, 2024). Additionally, inaccurate estimation of stability and control derivatives from simulated flight data can result in poorly tuned controllers, increasing the risk of instability or system crash during real-world UAV operations (Goel et al., 2021). A detailed fault tree of such failures in edge-based UAV orchestration has also been discussed in recent studies (Bai et al., 2023a). Reliability testing, hardware-in-the-loop simulations, and watchdog timers help manage the risk of software crashes.
5.5. Environmental impact
UAVs operate in complex and often harsh environments that can severely affect their performance and mission success. Below, we present the main environmental challenges that impact communication reliability, navigation accuracy, and overall system robustness.
Adverse Weather Conditions: Environmental factors such as extremes of wind, rain, snow, fog, temperature, and dynamic air traffic can dramatically affect UAV performance and cause network disconnections that disrupt communication and data flow(Wang et al., 2021b, b; Gallego-Madrid et al., 2020; Hossain et al., 2017; Pradhan et al., 2015; Haber et al., 2021). For example, strong winds can destabilize flight, while rain can interfere with electronics and sensors that are not adequately protected (Yang et al., 2022; Lee et al., 2022a; Sankarasrinivasan et al., 2015). Cold weather reduces battery efficiency, limiting flight time, and icy conditions may add weight or damage propellers. Environmental factors also affect thermal sensor accuracy (Barnawi et al., 2021). UAVs used for delivery or inspection must be ruggedized or equipped with weather-resilient sensors capable of adjusting flight behavior based on real-time input. Similarly, environmental factors such as sunlight glare, waves, and varying water color pose challenges for detection, while limited data availability further risks accuracy (Mcilwaine and Rivas Casado, 2021). Wind gusts may also cause sudden attitude changes, while electromagnetic interference, particularly near industrial zones or high voltage lines, can corrupt communication signals or disrupt GNSS-based navigation (Yang et al., 2022). Cloud cover and adverse conditions can cause data gaps or errors in mapping, reducing the reliability of timely agricultural monitoring (Fan et al., 2023). In this context, weather forecasting and adaptive flight planning systems are essential to ensure mission reliability and operational continuity. Additionally, UAVs operating at high altitudes are vulnerable to soft errors caused by radiation exposure, which can induce silent data corruptions in memory or logic circuits, affecting onboard computation and processes such as video summarization. These transient faults may go undetected, leading to reliability degradation during critical missions. To mitigate these risks, systems often require redundancy mechanisms, error-correcting codes, or radiation-hardened hardware components (Venkatagiri et al., 2018). Studies have shown that lightweight deep learning models such as MobileNet strike an effective balance between inference performance and power consumption, making them suitable for real-time trail detection in UAVs operating in adverse conditions (Kumar and Jannesari, 2023).
Terrain Challenges: Terrain features such as mountains, forests, urban environments, or bodies of water can obstruct signals or limit UAV maneuverability, directly affecting operational reliability (Dai et al., 2018). GPS signals may be reflected or blocked in canyons or dense urban areas, causing multipath errors and reduced navigational accuracy. Maintaining centimeter-level positioning accuracy becomes even more difficult under such conditions, particularly when using low-cost RTK-GNSS base stations (Rokaha et al., 2019). Tall buildings create radio control dead zones, while navigating tight spaces increases collision risks. In VANET-assisted UAV systems, urban obstacles such as buildings can disrupt radio waves, leading to frequent disconnections and failed message delivery (Paranjothi et al., 2019; Bakhtiari et al., 2020). UAVs face coverage challenges due to occlusions in geometrically complex areas like urban or mountainous terrains, which can degrade image quality (Gao et al., 2022; Peng et al., 2023). Limited onboard energy further necessitates efficient path planning to maximize mission effectiveness (Dai et al., 2018). Additionally, detection systems deployed in remote or challenging terrains often face network disconnections, which can delay alerts and responses critical for low-altitude security operations (Kim et al., 2024b; Faraci et al., 2023). Computational limits and channel obstacles hinder UAV task offloading, especially for dependent tasks modeled as directed acyclic graphs (Mao et al., 2023). These challenges necessitate robust collision avoidance, terrain-following sensors like LiDAR, and autonomous systems to maintain reliable UAV operation in complex environments (Shang et al., 2023). In addition to signal-related issues, hardware faults caused by environmental stress or mechanical wear in rugged terrains may lead to individual UAV failures during missions (Lv et al., 2024). Furthermore, network disconnections in obstructed environments degrade coordination among UAVs, potentially disrupting formation maintenance and task execution (Wang et al., 2021d). Obstacles at ground level can also interrupt UAV-to-IoT communication, while limited battery capacity and hardware faults may disrupt edge computing services; laser-based power transfer has been proposed to stabilize such operations (Hadi and Ghazizadeh, 2023). These challenges necessitate robust collision avoidance, terrain-following sensors like LiDAR, and autonomous systems to maintain reliable UAV operation in complex environments. Multi-UAV path planning also becomes difficult in uncertain, cluttered terrains where real-time obstacle avoidance and inter-UAV cooperation are critical for collision prevention and mission success (Zhou et al., 2024a).
5.6. Operational failures
UAV missions depend heavily on correct human operation and effective maintenance practices. Below, we present the main operational failure threats that affect system safety, reliability, and mission continuity.
Human Error:
Human error is a leading cause of UAV accidents, often arising during mission planning, parameter input (e.g., altitude, geofencing), incorrect mode selection, or delayed responses in emergencies (Xu and Zhai, 2020). Typical scenarios include unintentionally disabling obstacle avoidance or launching missions with insufficient battery charge, which can result in system loss, property damage, or injury. Regular training and procedural checklists help mitigate such errors (Hou et al., 2020a). Recent studies highlight a range of human-involved or human-impacting operational vulnerabilities. In MEC-enabled UAV systems, operator-induced mistakes or insufficient monitoring can degrade system performance under tight latency and energy constraints (Yu et al., 2021; Liu et al., 2023c; Yuan et al., 2022; Falcão et al., 2023). Human misjudgment during UAV placement, route optimization, and offloading decisions may cause connectivity failures and task delays, especially in dynamic and uncertain network conditions (Kim et al., 2024b; Mao et al., 2023; Faraci et al., 2023; Thapliyal et al., 2024). Inaccurate assessment of battery status or communication strength also contributes to mission failure, particularly when UAVs serve dual roles (e.g., computing and relaying) (Zhang et al., 2024d; Xu and Zhai, 2020; Manzoor et al., 2021). Scenarios involving limited pilot observation further complicate UAV status awareness, increasing the likelihood of unsafe maneuvers or late interventions (Jeong et al., 2020). In high-demand or post-disaster environments, human error in managing energy and computing resources may result in service outages or imbalanced workloads (He et al., 2023; Hui et al., 2024; Falcão et al., 2023; Li et al., 2024e; Kaleem et al., 2019). Moreover, inadequate responses to user-level interference (e.g., overloading UAV bandwidth) or security threats (e.g., eavesdropping) may stem from policy gaps or oversight limitations (Ozger et al., 2018; Rao and Naidu, 2024; Ouyang et al., 2023; Ju et al., 2024; Tang et al., 2023; Amos et al., 2021; Shen et al., 2023a). Finally, system-level miscoordination due to delayed or misinformed human actions, such as poor handoff management or overutilized UAVs, leads to reduced QoE and compromised safety (Wang et al., 2020a; Vamvakas et al., 2019; Asim et al., 2022). These scenarios underscore the importance of designing robust, human-aware systems, which include real-time feedback, operator training, and semi-autonomous error prevention mechanisms.
Maintenance Issues: Inadequate or skipped maintenance can lead to undetected issues such as loose connectors, worn-out rotors, degraded solder joints, or battery defects. Over time, these minor faults may escalate into critical in-flight failures, such as mid-air blade breakage. Ensuring dependability of UAV-based systems, therefore, demands regular inspections, maintenance documentation, and onboard diagnostic tools. Maintenance concerns in drone swarms involve node-level failures, communication, structural, or mission-related issues that degrade collective reliability; complex network modeling is used to evaluate their impact (Guo et al., 2022b). UAV inspection systems face hardware wear and safety hazards, especially when image resolution or network stability limits detection reliability (Jo et al., 2018; Wu et al., 2024). Environmental factors such as weather and mechanical degradation introduce hardware faults that disrupt coordination and task execution, particularly in dynamic terrains (Lv et al., 2024). Effective maintenance planning hinges on accurate failure forecasting. Time series models like Holt-Winters improve predictive accuracy, reducing downtime and safety risks (Yang et al., 2017). Engine and component malfunctions in UAVs or aircraft, coupled with disconnections in diagnostic links, lead to increased latency and unreliable maintenance actions (Ren et al., 2017; Zhang, 2020). Software bugs and vulnerabilities further necessitate continuous maintenance in cyber-physical and IIoT-based UAV systems (Pradhan et al., 2015; Lin et al., 2021). System performance can also be affected by faults in RFID tags, wearable computing devices, or embedded processors, where network disconnections and energy limits hinder tracking and diagnostic accuracy (Zhang and Zhang, 2024; Patnayak et al., 2021). In UAV-assisted robotic inspections, long-distance signal loss and energy constraints complicate real-time data processing and HD video transmission (Shams et al., 2023). For offshore wind farm monitoring, limited battery life and long-range travel demand optimized maintenance routes and inspection scheduling (Huang and Wang, 2024).
6. Dependability Techniques
In the rapidly evolving field of UAVs, a wide array of methodologies and techniques has emerged to optimize performance, address challenges, and enhance the effectiveness of UAV systems in various applications. This section outlines the key dependability techniques that have been extensively explored in the literature, categorizing them into four principal categories as shown in Figure 5. Each category represents a unique set of strategies and methods that contribute to the improvement of UAV capabilities. Within each of these categories, various subcategories and specific techniques have been developed and refined to address particular challenges, such as minimizing task completion time, ensuring reliable data transmission, or optimizing UAV energy consumption. These techniques not only focus on improving specific aspects of UAV performance but also enable more adaptable systems capable of responding to dynamic environments and real-time decision-making. As we examine each category in detail, we also explore how these techniques relate to each other and highlight trends that have emerged in recent research.
To systematically analyze the dependability techniques in UAV-enabled computing systems, the selected papers were categorized into four major groups derived from the foundational classification proposed by Avizienis et al. (2004): Fault Prevention, Fault Tolerance, Fault Removal, and Fault Forecasting. Each group consists of several subcategories, which reflect the specific methods or strategies employed in the reviewed papers.
6.1. Fault Prevention
Fault prevention includes techniques that aim to eliminate or reduce the likelihood of faults before they occur. These techniques are typically applied during the design and development stages of a system to ensure that faults are less likely to occur or be triggered during runtime. In the context of UAV-enabled environments, fault prevention is particularly important due to limited computing and energy resources, as well as the safety-critical nature of many UAV missions. The reviewed literature reveals a growing emphasis on prevention strategies that improve reliability and safety by addressing architectural, operational, and environmental vulnerabilities. Table 3 summarizes the subcategories identified in fault prevention and the corresponding related studies.
| Technique | Subcategory | Related Studies |
| Fault Prevention | Energy & Resource-Aware Design | (Ramos-Ramos et al., 2024), (Zhou et al., 2022a), (Wei et al., 2022), (Jeong et al., 2020), (Zhang and Bai, 2024), (Fatima et al., 2024), (Maatouk et al., 2024), (Xu et al., 2020), (Suganya et al., 2024), (Peng et al., 2024), (Palossi et al., 2019), (Rahul et al., 2021), (Wang et al., 2022b), (Asim et al., 2022), (Zhang et al., 2023a), (Alioua et al., 2018), (Zhang et al., 2025), (Yang LI, 2024), (Selim et al., 2021), (Xue et al., 2023a), (Tang et al., 2024a), (Wang et al., 2020d), (Laroui et al., 2021), (Lin et al., 2021), (Zhang et al., 2021b), (Raj and Simmhan, 2023), (Shao et al., 2023), (Seisa et al., 2022), (Hossain et al., 2017), (Lu et al., 2024), (Grasso and Schembra, 2019), (Dou et al., 2024), (Ma and Sun, 2022), (Qureshi et al., 2024), (Grasso et al., 2022a), (Gupta et al., 2021), (Wang et al., 2019a), (Chen et al., 2020a), (Awada et al., 2023), (Navardi and Mohsenin, 2023), (Mart´ınez-Sánchez et al., 2023), (Vélez et al., 2024), (Faraci et al., 2023), (Zeng et al., 2024b), (Shi et al., 2023), (Hong et al., 2022), (Grasso et al., 2022b), (Baktayan et al., 2022), (Hang et al., 2020), (Tsung et al., 2022), (Cheng et al., 2023), (Asheralieva and Niyato, 2020), (Geng et al., 2024), (Shams et al., 2023), (Gao et al., 2019), (Akbari et al., 2024), (Li et al., 2024a), (Hadjkouider et al., 2023), (Zheng et al., 2024), (Karam et al., 2024), (Xiang et al., 2024), (Venkatagiri et al., 2018), (Liu et al., 2019), (Callegaro et al., 2021), (Khan et al., 2022), (Li et al., 2020c), (Zhao et al., 2022), (Xu et al., 2024a), (Asheralieva and Niyato, 2019), (Yu et al., 2022a), (Huang et al., 2024b), (Apostolopoulos et al., 2023), (Jung et al., 2023a), (Zhang et al., 2018), (Wang et al., 2023b), (Haber et al., 2021), (Pandey et al., 2021a), (Wang et al., 2024b), (Dai et al., 2024), (Alam et al., 2019), (Subburaj et al., 2023), (Fatima et al., 2024), (Yang et al., 2020), (Gu et al., 2022), (Liu et al., 2023c), (Falcão et al., 2024), (Wang et al., 2020c), (Hou et al., 2020b), (Lu et al., 2022), (Yu et al., 2021), (Kim and Kim, 2021), (Manzoor et al., 2021), (Zhang et al., 2024d), (Thapliyal et al., 2024), (Awada et al., 2022), (Li et al., 2020a), (Li et al., 2024e), (Tang et al., 2024c), (Luo et al., 2023), (Lin et al., 2022), (Ghosh and Kuila, 2023), (Islam et al., 2019), (Ye et al., 2020), (Yu et al., 2022b), (Ju et al., 2024), (Ouyang et al., 2023), (Rao and Naidu, 2024), (Awada et al., 2023), (Zhang et al., 2024b), (Rokaha et al., 2019), (Hao et al., 2024), (Tang et al., 2022a), (Zhang et al., 2023d), (Pan et al., 2022), (Xu et al., 2022), (Liu et al., 2023b), (Haber et al., 2019), (Wang et al., 2024a), (Liang et al., 2021), (Tang et al., 2024b), (Zeng et al., 2023), (Wang et al., 2021a), (Hou et al., 2020a), (Chu et al., 2021), (Tang et al., 2022b), (Ju et al., 2024), (Chen et al., 2020c), (Yuan et al., 2022), (Falcão et al., 2023), (Liu et al., 2024b), (Hossain et al., 2022), (Chen et al., 2020b), (Xue et al., 2023c), (Hadi and Ghazizadeh, 2023), (Gao et al., 2020), (Kumar and Jannesari, 2023), (Pandey et al., 2021b), (Jeong et al., 2020), (He et al., 2023), (Han et al., 2021), (Dai et al., 2022), (Zhang et al., 2024a), (Wang et al., 2021e), (Zhang et al., 2022b), (Huang and Wang, 2024), (Zhang et al., 2022a), (Wei et al., 2021), (Qu et al., 2022), (Chai et al., 2023), (Wang et al., 2018), (Zhang et al., 2023e), (Zhang et al., 2024e), (Asim and Abd El-Latif, 2023), (Ouyang et al., 2024), (Seong et al., 2024), (Yu et al., 2023b), (Awada et al., 2022), (Geng et al., 2024), (Awada et al., 2021), (Zeng et al., 2024b), (Zhang et al., 2023c), (Hou et al., 2020a), (Khoramnejad et al., 2024), (Chen et al., 2024a), (Wang et al., 2020c), (Zheng et al., 2023), (Chu et al., 2021), (Li et al., 2024a), (Qureshi et al., 2024), (Hou et al., 2020b), (Wang et al., 2024b), (Tian et al., 2024), (Zhang et al., 2018), (Bezziane et al., 2024), (Chen et al., 2021a), (Basharat et al., 2023), (Chen et al., 2022a), (Zou et al., 2022), (D’Agati et al., 2024), (Zhao et al., 2024), (Su et al., 2022), (Thapliyal et al., 2024), (Gao et al., 2019), (Ma and Sun, 2022), (Akbari et al., 2024), (Xiong et al., 2019), (Al-habob et al., 2024), (Sun et al., 2020), (Navardi and Mohsenin, 2023), (Faraci et al., 2023), (Seisa et al., 2022), (Luo et al., 2021), (Haber et al., 2019), (Bao et al., 2023), (Bai et al., 2023b), (Jung et al., 2023a), (Ngoenriang et al., 2022), (Kong et al., 2022c), (Ning et al., 2024), (He et al., 2023), (Gupta and Gupta, 2022), (Picano and Fantacci, 2024), (Wang et al., 2021a), (Moeyersons et al., 2021), (Kong et al., 2022b), (Shao et al., 2023), (Pamuklu et al., 2023), (Kaleem et al., 2019), (Lin et al., 2021), (Garge and Balakrishna, 2018), (Hang et al., 2020), (Tendolkar et al., 2021), (Hui et al., 2024), (Xue et al., 2023b), (Gao et al., 2024), (Manzini et al., 2023), (Liu et al., 2024b), (Ju et al., 2023), (Pang et al., 2024) |
| Robust Trajectory & Collision-Avoidance Planning | (Lee et al., 2022b), (Krishnan et al., 2021), (Alqudsi et al., 2023), (Wang et al., 2022b), (Hu et al., 2024), (Asim et al., 2022), (Zhang et al., 2023a), (Xu et al., 2024c), (Singh et al., 2023), (Souli et al., 2023), (Yang LI, 2024), (Li et al., 2024c), (Kaymaz et al., 2024), (Zeng et al., 2024a), (Hirota et al., 2020), (Gao et al., 2021), (Chen et al., 2022a), (Zheng et al., 2024), (Liu et al., 2019), (Huang et al., 2024b), (Wang et al., 2023b), (Choutri et al., 2022), (Dai et al., 2024), (Subburaj et al., 2023), (Liu et al., 2023c), (Wang et al., 2020c), (Hou et al., 2020b), (Lu et al., 2022), (Yu et al., 2021), (Luo et al., 2021), (Tripathi et al., 2023), (Tang et al., 2024c), (Lee et al., 2022a), (Luo et al., 2023), (Dai et al., 2018), (Ye et al., 2020), (Behjat et al., 2019), (Koerkamp et al., 2019), (She and Ouyang, 2021), (Huang et al., 2019), (Ju et al., 2024), (Awada et al., 2023), (Li et al., 2018), (Deng et al., 2022), (Zhou et al., 2024a), (Baccarelli et al., 2024), (Li et al., 2024d), (Kim et al., 2024a), (Shen et al., 2023b), (Huang et al., 2024a), (Xue et al., 2023c), (Peng et al., 2022), (Lv et al., 2024), (Qayyum et al., 2021), (Yang et al., 2022), (Lorincz et al., 2021), (Wang et al., 2021e), (Yang et al., 2019) (Lindqvist et al., 2020), (Ye et al., 2022), (Guo et al., 2022a), (Bekkouche et al., 2018), (Tripathi et al., 2023), (De Paula Soares et al., 2023), (Zhan and Xu, 2024), (Zhou et al., 2022b), (Li et al., 2024e), (Gao et al., 2021), (Hirota et al., 2020), (Wu et al., 2023), (Zheng et al., 2024), (Jiang et al., 2019), (Ngoenriang et al., 2022), (Kong et al., 2022c), (Ning et al., 2024), (Wang et al., 2024d), (Marcos and Utete, 2021) |
6.1.1. Energy & Resource-Aware Design:
Energy & Resource-Aware Design focuses on strategies that proactively manage power consumption and computational workload to ensure stable UAV operations, particularly under limited energy and processing capacity. This approach is essential in UAV-based computing systems, where energy depletion or overload can lead to mission failure or system shutdown. This technique primarily addresses energy exhaustion, system overload, and thermal-related hardware degradation, which are critical dependability threats in UAV deployments. By carefully allocating resources, adjusting processing rates, or offloading tasks to nearby nodes or edge servers, UAVs can operate more sustainably and with greater resilience.
For instance, energy-aware offloading frameworks are proposed to reduce UAV energy consumption by dynamically assigning tasks to ground servers or other UAVs based on current battery levels and computational demands (Ramos-Ramos et al., 2024; Zhou et al., 2022a; Fatima et al., 2024; Wang et al., 2022b; Yang LI, 2024; Suganya et al., 2024; Peng et al., 2024; Zhang et al., 2025; Wang et al., 2020d; Seisa et al., 2022; Chen et al., 2020a; Faraci et al., 2023; Asheralieva and Niyato, 2020; Li et al., 2024a; Liu et al., 2019; Callegaro et al., 2021; Jung et al., 2023a; Haber et al., 2021; Yang et al., 2020; Awada et al., 2022). Additionally, many works integrate power control and bandwidth allocation schemes to adaptively regulate UAV transmissions and processing loads, thereby lowering overall energy usage while maintaining service quality (Zhang and Bai, 2024; Xu et al., 2020; Selim et al., 2021; Tang et al., 2024a; Hossain et al., 2017; Hang et al., 2020; Cheng et al., 2023; Gao et al., 2019; Xiang et al., 2024; Zhao et al., 2022; Zhang et al., 2018; Wang et al., 2023b; Liu et al., 2023c; Kim and Kim, 2021). These strategies directly contribute to enhancing system availability and reliability by preventing unexpected shutdowns and maintaining functional operation under varying load conditions.
6.1.2. Robust Trajectory & Collision-Avoidance Planning:
Robust Trajectory & Collision-Avoidance Planning refers to the proactive design of UAV flight paths that minimize the risk of collision, unsafe proximity, or environmental hazards. This technique is crucial for ensuring operational safety and fault prevention in autonomous UAV missions, especially in dynamic or congested airspaces. This technique directly mitigating threats related to mechanical breakdowns, terrain challenges, adverse weather conditions, and human error. Studies in this area typically employ trajectory optimization algorithms and collision-avoidance schemes to maintain navigational integrity and system continuity.
Multiple studies have proposed robust trajectory planning and avoidance frameworks that balance safety with performance. For example, model predictive control, velocity obstacles, and tabu search methods are commonly used to generate safe flight paths in constrained environments (Lee et al., 2022b; Wang et al., 2022b; Xu et al., 2024c; Bekkouche et al., 2018; Wu et al., 2023). Reinforcement learning and DRL-based approaches further enhance adaptivity, allowing UAVs to dynamically adjust paths in response to moving obstacles or real-time conditions (Hu et al., 2024; Behjat et al., 2019; Gao et al., 2021; Kong et al., 2022c; Ning et al., 2024). Some works also incorporate probabilistic models or risk-aware strategies to handle uncertainties such as wind gusts, localization errors, or multi-agent interactions (Zeng et al., 2024a; Liu et al., 2019; Luo et al., 2021; Zhan and Xu, 2024).
In high-density or mission-critical scenarios like disaster recovery or automated surveillance, multi-UAV coordination algorithms are introduced to prevent in-air collisions while minimizing delays or energy usage (Singh et al., 2023; Lee et al., 2022a; Li et al., 2018; Jiang et al., 2019). These techniques contribute directly to dependability metrics such as safety and reliability, enabling UAV systems to maintain continuous operation even in volatile or adversarial environments.
6.2. Fault tolerance
Fault tolerance includes techniques that enable UAV-based computing systems to continue functioning correctly even when faults occur during operation. These techniques are essential in dynamic and resource-constrained UAV environments where failures in hardware, software, or communication components may be inevitable. By incorporating mechanisms such as error detection and recovery, fault tolerance ensures system availability, reliability, and robustness in the presence of unexpected disruptions. The reviewed literature highlights various approaches that allow UAV systems to autonomously detect and recover from faults to maintain dependable mission execution. Table 4 summarizes the subcategories identified in fault tolerance and the corresponding related studies.
| Technique | Subcategory | Related Studies |
| Fault Tolerance | Error Detection | (Ahamad and Mishra, 2023), (Cheng et al., 2024), (Vemulapalli and Mesapam, 2021), (Gokulakrishnan et al., 2023), (Ma et al., 2024), (Saadaoui et al., 2023), (Kostoglotov et al., 2024), (Zhang and Li, 2018), (Zhou et al., 2024c), (Amin and Obermaisser, 2018), (Jeong et al., 2017), (Wang et al., 2021b), (Mo et al., 2022), (Chang et al., 2024), (Alam et al., 2019), (Wu et al., 2024), (Balamuralidhar et al., 2021), (Behjat et al., 2019), (Bakhtiari et al., 2020), (Rokaha et al., 2019), (Nwadiugwu et al., 2021), (Zhong et al., 2021), (Jeong et al., 2016), (Shang et al., 2023), (Pan et al., 2024), (Li et al., 2024b), (Zhou et al., 2023), (Zhang and Zhang, 2024), (Zheng et al., 2021), (Yuwei and Aijun, 2020) |
| Recovery | (Farrukh and West, 2023), (Luo et al., 2022), (Bobronnikov et al., 2022), (Alkinani et al., 2021), (Zuo et al., 2022), (Wang et al., 2021d), (Zhou and Cao, 2018), (Callegaro et al., 2021), (Choutri et al., 2022), (Wang et al., 2024b), (Hou et al., 2020b), (Luo et al., 2021), (Machida and Andrade, 2021), (Kim and Kim, 2021), (Li et al., 2024e), (Machida et al., 2023), (Yu et al., 2022b), (Ju et al., 2024), (Awada et al., 2023), (Zhang et al., 2024b), (Machida et al., 2023), (Petrenko et al., 2023), (Pan et al., 2022), (Watanabe and Machida, 2022), (Wang et al., 2019b), (Li et al., 2023a), (Damigos et al., 2023), (Machida and Andrade, 2021) | |
| Redundant Architecture | (Ma et al., 2024), (Luo et al., 2020), (Farrukh and West, 2023), (La Salandra et al., 2024), (Ren et al., 2017), (Kilbourne et al., 2021), (Nguyen et al., 2023), (Patnayak et al., 2021), (Awada et al., 2021), (Zhang et al., 2021c), (Gallego-Madrid et al., 2020), (Kaleem et al., 2019), (Bian et al., 2023), (Dogea et al., 2023), (Karam et al., 2024), (Aloqaily et al., 2021), (Liu et al., 2019), (Callegaro et al., 2021), (Khan et al., 2022), (Zhao et al., 2022), (Jung et al., 2023a), (Sharma et al., 2019), (Pacheco et al., 2021), (Hou et al., 2020b), (Awada et al., 2022), (Ali and Ali, 2022), (Kim, 2024), (Petrenko et al., 2023), (Sun et al., 2018), (El Sayed et al., 2023), (Sara et al., 2016), (Chen and Ma, 2024), (Barnawi et al., 2021), (Wang et al., 2020a), (Javanmardi et al., 2024), (Panwar et al., 2023), (Shen et al., 2023a), (Naouri et al., 2024), (Tang et al., 2023), (Pustokhina et al., 2021), (Taleb et al., 2023), (Patnayak et al., 2021), (Lorincz et al., 2021), (Karam et al., 2024), (Lei et al., 2023), (Li et al., 2023b), (Haber et al., 2021), (Li and Nie, 2023), (Pradhan et al., 2015) | |
| Network Topology Design | (Xu et al., 2021), (Guruprasad et al., 2024), (Zhang et al., 2023b), (Vamvakas et al., 2019), (Wang et al., 2024c), (Chen et al., 2022b), (Fujii et al., 2017), (Shao et al., 2023), (Mart´ınez-Sánchez et al., 2023), (Wang et al., 2021c), (Dao et al., 2021), (Ozger et al., 2018), (A. Alissa et al., 2023), (Tan et al., 2020), (Su, 2021), (Khan et al., 2021), (Wang et al., 2022c), (Zhou et al., 2024b), (Kong et al., 2022a), (Chen et al., 2024b), (Aloqaily et al., 2021), (Khan et al., 2022), (Li et al., 2020c), (Haber et al., 2021), (Sharma et al., 2019), (Gu et al., 2022), (Thapliyal et al., 2024), (Li et al., 2020b), (Zhang et al., 2021a), (Jung et al., 2023b), (Mukherjee et al., 2020), (Paranjothi et al., 2019), (Peng et al., 2023), (Nwadiugwu et al., 2021), (Mukherjee et al., 2023), (Mao et al., 2023), (Zhang et al., 2023b), (Yue et al., 2018), (Yadav and Baranwal, 2023), (Wang et al., 2020b), (Zhang et al., 2024c), (Amos et al., 2021), (Khanna et al., 2019), (Wang et al., 2023a), (Wang et al., 2022c), (Zhou et al., 2024b), (Liu et al., 2024c), (Liu et al., 2023a), (Taleb et al., 2023), (Gu et al., 2022), (Wang et al., 2019a), (Sharma et al., 2019), (Ren et al., 2024), (Shao et al., 2023), (Song et al., 2022), (Li et al., 2020c), (Ozger et al., 2018) |
6.2.1. Error Detection:
Error detection techniques are designed to identify faults during the operation of systems before they propagate into more severe failures. In UAV-based systems, error detection plays a crucial role in maintaining system reliability and safety, particularly under adverse conditions such as sensor degradation, software anomalies, or data transmission errors. These techniques often involve real-time monitoring of system states, anomaly detection models, and sensor consistency checks. They directly address dependability threats such as software faults, hardware faults, and communication errors, enabling timely mitigation actions.
Several studies employ diverse error detection mechanisms to enhance UAV system dependability. Signal integrity monitoring and consistency verification techniques are used to detect sensor or actuator faults during autonomous missions (Ahamad and Mishra, 2023; Vemulapalli and Mesapam, 2021; Gokulakrishnan et al., 2023). Software-based detection methods, such as model-driven diagnosis or runtime monitoring, are applied to identify software anomalies and functional deviations (Cheng et al., 2024; Ma et al., 2024; Jeong et al., 2017). Machine learning and statistical models are leveraged to recognize abnormal patterns from sensor data, aiding in early detection of both hardware degradation and external disruptions (Saadaoui et al., 2023; Kostoglotov et al., 2024; Alam et al., 2019). These error detection strategies contribute directly to system reliability, safety, and robustness by enabling early warning mechanisms and facilitating fault recovery processes.
6.2.2. Recovery
Fault recovery techniques aim to restore the normal operation of systems after a fault has occurred. In resource-constrained UAV environments, recovery strategies must be lightweight, adaptive, and capable of executing in real-time to avoid mission failure or system downtime. These techniques can involve system reboot, task re-execution, redundancy switching, or computational task migration to backup nodes. Fault recovery specifically addresses dependability threats such as software crashes, network disconnections, and hardware failures, helping maintain availability and resilience during unforeseen disruptions.
Many studies have explored recovery techniques to improve UAV system availability and minimize downtime after faults. Lightweight task migration and redundancy-based switching strategies are frequently employed to reallocate computing workloads in response to partial system failures or degraded node performance (Luo et al., 2022; Zuo et al., 2022; Luo et al., 2021; Awada et al., 2023; Pan et al., 2022). In scenarios where UAVs operate in harsh or disconnected environments, distributed task re-execution or fallback mechanisms are introduced to ensure mission continuity (Callegaro et al., 2021; Choutri et al., 2022; Hou et al., 2020b; Machida et al., 2023, 2023). Energy-aware recovery approaches are also proposed, balancing the trade-off between recovery success and remaining UAV resources (Wang et al., 2024b; Yu et al., 2022b; Zhang et al., 2024b; Watanabe and Machida, 2022).
6.2.3. Redundant Architecture
Redundant architecture refers to fault tolerance strategies where multiple instances of hardware, software components, or communication paths are integrated into UAV-based systems to maintain service continuity in the presence of faults. Rather than preventing faults from occurring, these techniques aim to ensure that system functionality is preserved despite failures in some components. In UAV-based systems, redundancy is employed across communication links, computation nodes, sensors, and actuator subsystems. For example, dual communication modules and duplicated computing resources are used to tolerate communication failures and processing overloads without interrupting mission-critical services (Ma et al., 2024; Farrukh and West, 2023; Kilbourne et al., 2021; Awada et al., 2021; Karam et al., 2024).
Multi-UAV swarms with overlapping coverage enable persistent sensing and data transmission even when individual units become inoperative (Patnayak et al., 2021; Zhang et al., 2021c; Bian et al., 2023; Jung et al., 2023a; Kim, 2024). Backup control modules and redundant edge/fog nodes are also integrated to support recovery and seamless task offloading in the event of partial system degradation or network instability (Kaleem et al., 2019; Dogea et al., 2023; Callegaro et al., 2021; Pacheco et al., 2021; Petrenko et al., 2023; El Sayed et al., 2023; Sara et al., 2016). Additionally, hybrid cloud-edge frameworks offer parallel computational paths to guarantee task execution continuity under localized failures or load fluctuations (Luo et al., 2020; La Salandra et al., 2024; Khan et al., 2022; Chen and Ma, 2024; Wang et al., 2020a; Shen et al., 2023a; Tang et al., 2023). These fault-tolerant architectures directly contribute to dependability attributes such as availability, reliability, and resilience by enabling uninterrupted operations and graceful degradation in the presence of faults, rather than complete system failure.
6.2.4. Network Topology Design
Network Topology Design refers to fault tolerance strategies that focus on maintaining reliable and resilient communication within systems, especially in environments where signal disruption, disconnection, or routing instability is common. Rather than preventing communication faults outright, these techniques aim to sustain service availability and network functionality even when individual links or nodes fail during operation. In UAV-based systems, dynamic topologies such as flying ad hoc networks (FANETs), mesh structures, and mobile relays are often employed to mitigate link degradation and maintain multi-hop connectivity. For instance, adaptive topology control and link prediction mechanisms are used to anticipate and respond to potential disconnections due to UAV mobility or energy depletion (Xu et al., 2021; Guruprasad et al., 2024; Zhang et al., 2023b; Vamvakas et al., 2019; Chen et al., 2022b; Wang et al., 2021c; Dao et al., 2021; Ozger et al., 2018; Su, 2021; Wang et al., 2022c).
UAV-aided reconfiguration is another popular approach, where drones are repositioned to act as temporary relays or mobile base stations to bridge coverage gaps and reduce the risk of total communication failure (Zhou et al., 2024b; Kong et al., 2022a; Zhang et al., 2021a; Mukherjee et al., 2020; Nwadiugwu et al., 2021; Mukherjee et al., 2023; Yue et al., 2018; Wang et al., 2020b; Zhou et al., 2024b; Wang et al., 2019a). These topology-aware fault tolerance mechanisms directly contribute to availability, connectivity, and system resilience, ensuring that UAV missions such as remote surveillance, disaster response, or public safety operations can continue despite disruptions in the communication infrastructure.
6.3. Fault Removal
Fault removal includes techniques designed to identify and eliminate existing faults in systems, typically during the development, integration, or testing phases. Unlike fault prevention, which aims to avoid the introduction of faults, fault removal focuses on detecting and correcting latent faults before they manifest in the operational environment. In UAV-enabled systems, this process is critical due to the high stakes of operational failure, especially in missions involving surveillance, delivery, or disaster response. Common approaches include static and dynamic analysis, software debugging, formal verification, and simulation-based testing. These techniques enhance reliability, safety, and robustness by ensuring the system behaves as intended under various conditions. As shown in Table 5, the reviewed studies that fall under this category demonstrate the growing emphasis on systematic fault detection and elimination to increase overall system dependability.
| Technique | Subcategory | Related Studies |
| Fault Removal | Testing & Verification | (Yoshimoto et al., 2020), (Kumar et al., 2021), (Jung et al., 2018), (Goel et al., 2021), (Gao et al., 2022), (Lu et al., 2023), (Ritchie et al., 2020), (Chang et al., 2024), (Wu et al., 2024), (Balamuralidhar et al., 2021), (Koerkamp et al., 2019), (Fan et al., 2023), (Sharma et al., 2021), (Russell et al., 2018), (Feng et al., 2023), (Hines and Bekemeyer, 2023), (Bowley et al., 2019), (Sankarasrinivasan et al., 2015), (Mcilwaine and Rivas Casado, 2021), (Zhang et al., 2024f), (Yu et al., 2023a), (Jiang and He, 2023), (Krishnan et al., 2021), (Sabino et al., 2024), (Kostoglotov et al., 2024), (Bouzidi et al., 2020), (Kilbourne et al., 2021) |
| Diagnosis | (Xu and Zhai, 2020), (Venkatagiri et al., 2018), (Deng et al., 2021), (Jo et al., 2018), (Amin and Obermaisser, 2018), (Ramasamy et al., 2024), (Zhang and Zhang, 2024), (Wang et al., 2021b), (Wang et al., 2022a), (Wang et al., 2024c), (Kostoglotov et al., 2024), (Bobronnikov et al., 2022), (Bouzidi et al., 2020), (Amin and Obermaisser, 2018) | |
| Correction | (Gu et al., 2022), (Luo et al., 2022), (Bobronnikov et al., 2022) | |
| Non-regression Verification |
6.3.1. Testing and Verification:
Testing and Verification techniques are essential for uncovering latent faults in UAV-based computing systems before deployment. Their primary objective is to ensure that the system meets design specifications and behaves correctly across a variety of expected and unexpected conditions. These techniques primarily address software faults and hardware faults by identifying inconsistencies, logic errors, or integration issues during development. Moreover, they significantly contribute to safety and reliability, thereby reducing the likelihood of mission-critical failures in real-time UAV operations. Numerous studies have employed testing and verification strategies to validate the functionality of UAV-based systems under complex and uncertain conditions. These techniques for UAV-based systems can be broadly categorized into fuzzing-based testing, simulation and modeling, formal verification, dataset-based evaluation, and edge-integrated real-time testing.
Fuzzing-based methods such as LDFuzzer aim to identify abnormal flight behaviors caused by configuration errors, using genetic algorithm-guided input mutation and lightweight test oracles to ensure efficient detection under constrained UAV resources (Chang et al., 2024). Simulation-driven approaches validate UAV dynamics or network performance under controlled conditions, as seen in flight stability estimation with neural networks (Goel et al., 2021), AFDX network simulation using OPNET (optimized network engineering tool) (Jiang and He, 2023), and autonomous image processing pipeline development for real-time geolocation correction (Feng et al., 2023). On the other hand, formal modeling techniques ensure protocol-level correctness in time-critical UAV tasks, exemplified by UPPAAL-based verification of serial communication in autonomous UAV-based systems (Jung et al., 2018). Dataset-based validation focuses on the robustness of visual perception models through large-scale annotated datasets, including P-DESTRE for pedestrian re-identification (Kumar et al., 2021), JellyNet for jellyfish bloom detection (Mcilwaine and Rivas Casado, 2021), and MultEYE for multi-task aerial traffic analysis (Balamuralidhar et al., 2021). Finally, edge-integrated real-time testing is increasingly applied in UAV deployments, where detection and monitoring systems run directly on UAVs or embedded platforms such as super-resolution-based defect detection (Wu et al., 2024), helmet compliance monitoring (Sharma et al., 2021), onboard intelligent inspection platforms (Yu et al., 2023a), and near-real-time object identification pipelines in agricultural monitoring (Feng et al., 2023). These studies demonstrate how rigorous testing and verification processes serve as foundational mechanisms for identifying and removing faults, thereby significantly strengthening the reliability and safety of UAV-based systems.
6.3.2. Diagnosis:
Diagnosis techniques focus on identifying the root cause of faults or abnormal behaviors in UAV-based computing systems after they have occurred. These methods aim to accurately localize faults, enabling targeted mitigation, correction, or recovery actions. In the context of UAV-enabled networks, diagnosis plays a vital role in maintaining operational continuity, especially when systems operate autonomously in unpredictable environments.
Diagnosis primarily addresses software faults, hardware faults, and network disconnection, where faults may manifest as unexpected delays, abnormal UAV motion, or degraded communication quality. Several studies utilize behavioral and model-based diagnosis to identify operator-induced errors by comparing them with expected control models during simulated missions, providing insights into human-machine interaction errors and their safety implications (Xu and Zhai, 2020; Amin and Obermaisser, 2018). Secondly, cloud and edge-enabled anomaly detection frameworks are increasingly employed to enhance scalability and responsiveness, particularly for large-scale UAV applications such as power line inspection and predictive maintenance. These include distributed isolation forest algorithms and embedded monitoring tools that enable timely fault isolation while minimizing computational burden (Ramasamy et al., 2024; Zhang and Zhang, 2024; Wang et al., 2022a).
On the other hand, intelligent agents and software-based diagnostic systems support modular and autonomous operation by integrating distributed reasoning or software approximation techniques. For example, agent-based building inspection platforms and approximation-aware video summarization systems leverage real-time assessment capabilities to maintain operational resilience and energy efficiency (Venkatagiri et al., 2018; Jo et al., 2018). Meanwhile, cooperative and network-level diagnosis approaches utilize drone-assisted sensing and edge intelligence to detect anomalies within broader communication networks, including space-air-ground infrastructures, thereby ensuring systemic fault awareness and coordination (Deng et al., 2021; Wang et al., 2021b).
6.3.3. Correction:
Correction techniques aim to modify system behavior or state to eliminate or mitigate the impact of detected faults, ensuring that systems continue to operate safely and effectively. Unlike recovery, which focuses on returning a system to a previous known good state, correction involves making targeted adjustments that are often based on diagnostic results. This directly addresses the underlying fault or its manifestations. In UAV-enabled environments, correction techniques are particularly crucial in addressing software faults, sensor anomalies, and configuration errors, as these issues, if left unresolved, can compromise system reliability, robustness, and safety.
These techniques are typically implemented through adaptive control, parameter reconfiguration, or algorithm-level adjustments. For instance, A study that presents a correction mechanism that adapts offloading and communication strategies when faced with security or reliability threats during UAV operations (Khan et al., 2022). Similarly, the smart onboard information and computing system (SOICS) framework enables onboard parameter correction and task reallocation under dynamic UAV mission loads and hardware constraints (Bobronnikov et al., 2022). These adaptive correction schemes are particularly effective when UAVs operate under resource limitations or uncertain environments. In another direction, a study proposes a model-driven online adaptation mechanism that predicts requirement violations based on environmental conditions and proactively selects optimal corrective strategies for UAV delivery and surveillance missions (Luo et al., 2022).
6.3.4. Non-regression Verification:
Non-regression verification ensures that newly applied system fixes or updates do not accidentally introduce new faults or degrade previously verified functionality. In UAV-based computing systems, where mission-critical software updates, reconfigurations, or adaptations are common, this verification step is essential to enforce reliability and security after system modifications. This technique is particularly relevant in dynamic UAV environments, where frequent software adaptations, task reassignments, or system upgrades occur, and even small regressions can lead to critical failures or mission disruptions. Non-regression verification ensures that fixes or corrections applied during fault removal do not negatively impact system behavior in other parts of the system or under different operating conditions.
Despite its critical importance for maintainability and long-term reliability, none of the reviewed papers explicitly implement or evaluate non-regression verification mechanisms. This absence suggests a significant gap in the current body of research about dependability of UAV-based systems, where emphasis has largely been placed on pre-deployment testing, runtime error detection, or recovery mechanisms, but not on validating system integrity post-modification.
6.4. Fault Forecasting
Fault forecasting refers to techniques that aim to predict the future occurrence or impact of faults. Rather than preventing faults or correcting them after they occur, fault forecasting focuses on evaluating the likelihood, timing, or consequences of potential faults. This enables proactive dependability planning and dynamic resource management in UAV environments where uncertainty and system variability are inherent. In the context of UAV-based systems, fault forecasting is particularly valuable for supporting predictive maintenance, service continuity planning, and early failure detection in constrained and distributed infrastructures. Techniques under this category often rely on probabilistic models, historical performance logs, or learning-based analytics to anticipate degradation trends and inform preemptive countermeasures. Table 6 summarizes the subcategories identified under fault forecasting and highlights the related studies that incorporate predictive mechanisms to enhance system reliability, availability, or performance stability in UAV-based systems.
| Technique | Subcategory | Related Studies |
| Fault Prevention | Ordinal Evaluation | (Bouzidi et al., 2020), (Guo et al., 2022b), (Hossain et al., 2017), (Manzini et al., 2023), (Fesenko et al., 2023), (Venkatagiri et al., 2018), (Kramer and Stubberud, 2021), (Cao et al., 2021) |
| Probabilistic Evaluation | (Sabino et al., 2024), (Ngoenriang et al., 2022). (Goel et al., 2021), (Tahat et al., 2020), (Ji et al., 2024), (Apostolopoulos et al., 2023), (Bai et al., 2023a), (Brito et al., 2021), (Falcão et al., 2024), (Machida and Andrade, 2021), (Kim and Kim, 2021), (Manzoor et al., 2021), (Thapliyal et al., 2024), (Machida et al., 2023), (Dai et al., 2018), (She and Ouyang, 2021), (Machida et al., 2023), (Chen et al., 2021a), (Yang et al., 2017), (Falcão et al., 2023), (Watanabe and Machida, 2022), (Schäfer et al., 2018), (Jeong et al., 2020), (Picano and Fantacci, 2024), (Kabashkin, 2023), (Yao and Ansari, 2018), (She et al., 2018), (Silva et al., 2024), (Hu and Seiler, 2015), (Falcão et al., 2023), (Li et al., 2019), (Bai et al., 2023a), (Venkatagiri et al., 2018), (Zhang, 2020) |
6.4.1. Ordinal Evaluation:
Ordinal Evaluation techniques in fault forecasting are designed to qualitatively or semiquantitatively assess the likelihood or severity of faults in UAV-based systems. Rather than producing precise probabilistic output, these methods rank system states or events based on relative levels of risk, reliability degradation, or fault impact. This approach is particularly useful in UAV environments where data may be incomplete or uncertain, and where full statistical modeling is infeasible due to resource constraints. This technique contributes to dependability by enhancing fault awareness and supporting preventive maintenance and operational readiness, particularly under uncertain or evolving conditions. For example, ordinal-based risk frameworks are employed to classify mission-critical parameters, such as link stability and sensor health, enabling early warnings before failures occur (Bouzidi et al., 2020; Guo et al., 2022b; Hossain et al., 2017).
6.4.2. Probabilistic Evaluation:
Probabilistic Evaluation techniques aim to quantitatively estimate the likelihood of faults or system failures based on statistical models and historical or runtime data. In UAV-based computing systems, these techniques are critical for understanding how component reliability, network conditions, or operational environments contribute to the probability of fault occurrences. Common approaches include Markov models, Bayesian inference, and stochastic Petri nets, which model system states and transitions under uncertainty. Such methods are especially valuable in resource-constrained aerial platforms where proactive fault handling must be supported by data-driven decision-making. These techniques help quantify expected system availability, failure rates, and downtime, enabling UAV designers and operators to make risk-aware trade-offs between performance, energy efficiency, and dependability.
In the context of dependability of UAV-based systems, probabilistic evaluation directly supports reliability analysis, availability estimation, and maintenance planning, offering a mathematical foundation for system robustness under varying workloads and fault scenarios. For example, studies that apply probabilistic models to assess component failure rates and communication link interruptions under uncertain conditions (Sabino et al., 2024; Ngoenriang et al., 2022; Tahat et al., 2020). Other studies integrate Markov models and stochastic processes to estimate system uptime and mission success rates (Ji et al., 2024; Apostolopoulos et al., 2023; Bai et al., 2023a). Bayesian networks and Monte Carlo simulations are used to analyze the effects of uncertain environmental conditions or resource variability on UAV operations (Falcão et al., 2024; Kim and Kim, 2021; Manzoor et al., 2021).
7. Future research directions
While substantial progress has been made in enhancing the dependability of UAV-based computing and networking systems, our review reveals several critical gaps and emerging challenges that demand further investigation. This section outlines key directions for future research, as detailed below.
7.1. LLM/AI agent-based approaches
The integration of Large Language Models (LLMs) and AI agents into UAV systems holds transformative opportunities for autonomy and adaptability, yet also introduces novel and critical dependability challenges (Javaid et al., 2024; Chen et al., 2023; Wassim et al., 2024). LLMs facilitate natural language interaction, enabling operators to issue high-level commands using plain language instructions. AI agents powered by LLMs further expand autonomy through multimodal decision-making capabilities. For instance, models like Claude 3.5 (Anthropic, 2024) point to a future where AI agents could synthesize data from cameras, LiDAR, and environmental sensors to dynamically adapt flight paths and intelligently prioritize tasks in complex or cluttered environments. Additionally, LLMs can process unstructured data, such as weather updates, to adapt flight paths in real-time, while AI agents can predict dependability threats like battery depletion or network congestion, enabling proactive fault detection and recovery. Examples include semantic fusion of sensor streams, context-aware mission reconfiguration, or resilient edge-based decision-making when cloud connectivity is degraded. Combined with learning-based planning or reinforcement learning, AI agents can improve UAV swarm coordination and energy-efficient operation under evolving mission constraints.
However, integrating LLMs and AI agents into UAV systems also introduces critical dependability concerns. First, their high computational demands strain UAVs’ limited resources, accelerating battery depletion and risking mission failures (Huang et al., 2024c). While edge-cloud offloading and model compression can mitigate this, latency-sensitive tasks such as obstacle avoidance may still require hybrid architectures that combine LLM-driven strategies with traditional control loops. Secondly, the non-deterministic and non-transparent nature of LLMs may result in unpredictable or erroneous outputs (“hallucinations”), especially in safety-critical applications (Xu et al., 2024b). Ensuring robustness requires redundancy and validation layers, such as cross-checking LLM outputs against rule-based systems or applying uncertainty-aware thresholds.
To ensure dependability, system designs must address energy-performance trade-offs and fail-safe mechanisms. Architectures that combine symbolic safety layers with neural models may offer a hybrid pathway toward dependable intelligent UAV systems. As AI agents become more integrated into aerial platforms, a systematic dependability framework will be essential to ensure trust and safety in LLM-enabled UAV autonomy.
7.2. Constellation of satellites
Satellite connectivity offers significant breakthroughs for UAV-based systems by enabling global coverage, allowing operations in remote areas like oceans or disaster zones where terrestrial networks are unavailable. Satellite links also provide resilient communication during ground network outages, ensuring mission continuity in emergencies (Zhu et al., 2024). It supports real-time swarm coordination across large regions, enhances navigation in GPS-denied environments using alternative satellite systems, and provides high bandwidth for data-intensive tasks like HD video streaming, as enabled by low Earth orbit (LEO) constellations like Starlink (Izhikevich et al., 2024; Chaudhry and Yanikomeroglu, 2021).
However, challenges include high latency with geostationary satellites (500–700 ms), which can delay critical decisions (Vankka, 2013), and increased energy demands from satellite hardware (Esho et al., 2024), risking battery depletion and reducing flight duration. Additional concerns include link intermittency due to weather or obstructions, vulnerability to jamming or spoofing attacks, and the high cost of commercial satellite services.
To address these, reliability considerations involve energy-efficient designs, such as low-power transceivers and burst transmissions, as well as hybrid communication strategies that combine satellite and terrestrial links (Wang et al., 2019c) with fallback mechanisms. Robust signal resilience could be realised through adaptive modulation, latency-aware decision-making, and security measures. While satellite-enabled UAVs offer unprecedented operational range and flexibility, their deployment requires careful architectural design and robust fault-tolerance mechanisms to maintain dependable performance in dynamic and contested environments.
7.3. B5G/6G network integration
As UAVs are increasingly deployed in time-critical and data-intensive applications, such as real-time surveillance, smart city monitoring, and emergency response, the limitations of 4G and early 5G infrastructures become more pronounced. B5G/6G technologies promise ultra-reliable low-latency communication (URLLC), massive machine-type communication (mMTC), and high spectral efficiency, which are essential for handling high-bandwidth tasks like video streaming, edge AI inference, and swarm coordination among UAVs (Qadir et al., 2023).
Future research should explore the design of UAV-based systems that are tailored to B5G/6G environments (Shrestha et al., 2021). In addition, joint optimization of communication and computing, including seamless handovers and UAV trajectory planning with connectivity guarantees, will be crucial to maintain QoS under highly dynamic aerial conditions. Integration with terahertz communication, intelligent reflecting surfaces (IRS), and network function virtualization (NFV) can further extend the coverage, reduce energy consumption, and improve resilience of UAV networks (Chen et al., 2021b).
7.4. DRL and collaborative learning
Deep Reinforcement Learning (DRL) has emerged as a powerful tool for enabling autonomous decision-making in UAV-based computing systems, especially in dynamic and uncertain environments (Liu et al., 2024a). It facilitates real-time task offloading, trajectory optimization, and resource management by learning optimal strategies through continuous interaction with the environment. However, most current DRL-based approaches are designed for single-agent systems, which limits their scalability and adaptability in multi-UAV deployments.
To address this, future research should focus on multi-agent DRL (MADRL) and collaborative learning frameworks, where UAVs learn not only from their own experiences but also from the behaviors and outcomes of other UAVs in the network (Feng et al., 2024). Such approaches enable distributed intelligence, improve energy and spectrum efficiency, and enhance the fault-tolerance of the overall system.
7.5. Urban air mobility and smart city integration
Urban Air Mobility (UAM) is rapidly evolving as a key frontier in UAV deployment, aiming to integrate aerial vehicles into smart city infrastructures for tasks ranging from traffic monitoring and infrastructure inspection to cargo transport and air taxi services. Unlike traditional UAV applications, which are confined to isolated missions, UAM envisions the continuous and coordinated operation of multiple drones within densely populated urban environments. This shift introduces complex challenges in airspace management, collision avoidance, noise pollution, communication reliability, and public safety (Pak et al., 2024; Yan et al., 2024). To ensure dependable UAM operations, future research must focus on scalable air traffic control algorithms, dynamic geofencing, and AI-driven decision-making to manage interactions among drones, manned aircraft, and urban infrastructure (Asmer et al., 2024; Wu et al., 2019). Dependability in dense urban environments—often characterized by “urban canyons” with high RF interference and GPS signal blockage—necessitates robust communication protocols and failover mechanisms (Becker and Schalk, 2024; Salehi et al., 2022).
Furthermore, the convergence of UAM and ground-based Internet of Things systems introduces additional integration challenges and requirements for interoperability and synchronized data flows (Yan et al., 2024). This integration requires a system-of-systems design philosophy, where the performance, safety, and fault tolerance of each subsystem are ensured under dynamic, uncertain, and often adversarial operating conditions. Achieving dependable UAM operations in future smart cities will require joint progress in autonomous UAV system design, urban communication infrastructure, policy standardization, and multi-modal transportation integration.
7.6. Precision agriculture and environmental sustainability
UAVs are becoming indispensable assets in precision agriculture, enabling real-time, high-resolution monitoring of soil conditions, crop health, water distribution, and pest outbreaks through multispectral, hyperspectral, and thermal sensors (Narzari et al., 2025; Makam et al., [n. d.]; Moghimi et al., 2020). These sensing capabilities support targeted and data-driven agricultural interventions that help minimize resource waste and reduce environmental impact (Sa et al., 2017). Despite these benefits, achieving dependable UAV operations in agricultural environments remains a critical challenge. Such environments often expose UAVs to harsh stressors like wind, dust, and humidity, which can impair sensor performance and flight stability. Additionally, infrastructure limitations—notably unreliable connectivity and restricted energy availability in remote farmlands—impede consistent UAV operation (Hossen et al., 2021). These constraints are exacerbated during long-duration missions, which place considerable strain on battery resources and complicate logistics for maintenance and recharging.
Future studies should focus on reinforcing system robustness, ensuring energy autonomy, and advancing intelligent coordination to guarantee that UAV systems can operate reliably and sustainably under the real-world constraints of agricultural landscapes.
7.7. Operational safety
As UAVs are increasingly deployed in mission-critical and autonomous contexts, traditional safety assessments centered on hardware faults or communication breakdowns have become insufficient to capture the growing complexity of operational risks. New classes of threats are emerging, driven by AI-based decision-making, rapidly changing environments, and intricate human-system interactions (Tsamados et al., 2024; Khatiri et al., 2025). One of the most pressing concerns is the lack of transparency in AI decision-making processes. The “black-box” nature of deep learning models makes it challenging to interpret or predict UAV behavior, particularly in unstructured or high-stakes scenarios (Willers et al., 2020). This opacity undermines the ability to certify and validate safety guarantees, raising questions about accountability and trustworthiness. Additionally, the integration of autonomous capabilities introduces critical issues in human-machine coordination. Achieving reliable and intuitive collaboration between UAV systems and human operators remains a major challenge, as many current frameworks do not adequately support real-time co-adaptation or mutual understanding (Tsamados et al., 2024).
To address this challenge, Explainable AI (XAI) approaches may help increase transparency by making AI-driven behaviors more understandable, thus supporting better human oversight and trust in autonomous decisions (Willers et al., 2020). Ensuring explainability will be essential to advancing the safety and dependability of future aerial platforms.
7.8. Maintainability challenges
As UAVs evolve from short-term trials to long-term deployments, maintainability becomes a key component of system dependability. Challenges such as delayed fault detection, limited upgrade paths, and complex servicing in the field can lead to mission failure and prolonged downtime, especially in harsh or remote environments where logistics are limited and hardware-software integration complicates repairs (Hashim, 2024; Vyasa and Xu, 2024). UAVs often operate in environments that pose significant challenges to maintainability. These include exposure to harsh weather conditions, limited access to maintenance facilities, and the complexity of integrating diverse hardware and software components. Such factors can lead to increased wear and tear, unexpected failures, and difficulties in performing timely repairs or upgrades. Additionally, the lack of standardized diagnostic tools and protocols can hinder effective fault detection and resolution, further compromising the reliability and availability of UAV systems (Wang et al., 2025).
To address these challenges, maintainability-focused UAV designs can incorporate modular architectures that simplify hardware replacement and software updates, enabling rapid recovery during field operations. For instance, integrating modular components allows for easier swapping of faulty parts, reducing downtime and simplifying the maintenance process (Perno et al., 2023). Moreover, future studies also need to focus more on prognostics, real-time health monitoring, and predictive maintenance to ensure the long-term dependability of UAV-based systems.
8. Threats to validity
As with this survey’s findings and conclusions, several threats to validity may influence the completeness, accuracy, and generalizability of the results.
Research Question Formulation
The formulation of research questions fundamentally shapes the scope and focus of the survey. Although the RQs were carefully formulated to cover key aspects of dependability of UAV-based systems, any ambiguity or overly narrow framing may have led to the exclusion of relevant studies or overemphasis/underrepresented on certain dimensions, potentially skewing the synthesized outcomes.
Search Strategy and Coverage
We followed a systematic protocol, but the search process may not have captured all relevant studies due to database-specific limitations and language restrictions. Although a comprehensive set of search strings was developed, minor adjustments were made to tailor the queries for specific databases, which could introduce inconsistencies. Consequently, it is possible that some relevant articles were inadvertently omitted due to differences in metadata standards or keyword variability.
Selection and Data Extraction Bias
During the study selection and data extraction phases, subjective judgment was involved in interpreting the relevance and classification of articles. Despite multiple authors independently participating in screening and resolving discrepancies based on predefined inclusion and exclusion criteria, individual interpretations of a study’s relevance or focus could have led to classification errors or overlooking subtle distinctions. After individual selection, any disagreements were settled by consensus discussions.
Temporal Bias
Our review focuses on literature published between 2015 and mid-2024. While this window captures the most recent advancements, there is a possibility that very recent studies published after our cutoff date or in press may not be included.
Publication Bias
As our review primarily considered peer-reviewed journal and conference articles, other high-quality research reported in non-traditional venues (e.g., white papers, technical reports, preprints) may have been excluded, potentially omitting novel but less formalized developments in the field. Similarly, regional research disparities (e.g., heavier emphasis on 5G in Asian vs. European studies) might skew the perceived prevalence of certain methodologies.
9. Conclusion
In this survey, we conducted a systematic literature review to investigate the state of research on the dependability of UAV-based networks and computing systems. By applying rigorous inclusion and exclusion criteria, we identified and analyzed 458 relevant publications from the past decade. Our review highlights the growing academic and practical attention to dependability of UAV-based systems, driven by their increasing deployment in mission-critical applications. We systematically examined the dependability threats that challenge UAV operations, while also evaluating a range of mitigation techniques. Furthermore, we provided an outlook on future research directions, proposing advancements such as integrated dependability frameworks, AI-driven predictive models, and 6G-enabled networking to address emerging challenges and enhance the reliability of UAV systems.
Acknowledgements.
This work was supported by JST SPRING Grant Number JPMJBS2414, and partly supported by a grant from the Telecommunications Advancement Foundation.References
- (1)
- A. Alissa et al. (2023) Khalid A. Alissa, Mohammed Maray, Areej A. Malibari, Sana Alazwari, Hamed Alqahtani, Mohamed K. Nour, Marwa Obbaya, Mohamed A. Shamseldin, and Mesfer Al Duhayyim. 2023. Optimal deep learning model enabled secure UAV classification for industry 4.0. Comput. Mater. Contin. 74, 3 (2023), 5349–5367.
- Adil et al. (2024) Muhammad Adil, Houbing Song, Mian Ahmad Jan, Muhammad Khurram Khan, Xiangjian He, Ahmed Farouk, and Zhanpeng Jin. 2024. UAV-assisted IoT applications, QoS requirements and challenges with future research directions. Comput. Surveys 56, 10 (2024), 1–35.
- Ahamad and Mishra (2023) Rayees Ahamad and Kamta Nath Mishra. 2023. Knowledge discovery of suspicious objects using hybrid approach with video clips and UAV images in distributed environments: a novel approach. Wirel. Netw. 29, 8 (Nov. 2023), 3393–3416.
- Ahamad and Mishra (2024) Rayees Ahamad and Kamta Nath Mishra. 2024. Hybrid approach for suspicious object surveillance using video clips and UAV images in cloud-IoT-based computing environment. Cluster Comput. 27, 1 (Feb. 2024), 761–785.
- Ahmed et al. (2025) Manzoor Ahmed, Aized Amin Soofi, Feroz Khan, Salman Raza, Wali Ullah Khan, Lina Su, Fang Xu, and Zhu Han. 2025. Toward a Sustainable Low-Altitude Economy: A Survey of Energy-Efficient RIS-UAV Networks. arXiv preprint arXiv:2504.02162 (2025).
- Akbari et al. (2024) Mohammad Akbari, Aisha Syed, W Sean Kennedy, and Melike Erol-Kantarci. 2024. AoI-aware energy-efficient SFC in UAV-aided smart agriculture using asynchronous federated learning. IEEE Open J. Commun. Soc. 5 (2024), 1222–1242.
- Al-habob et al. (2024) Ahmed A Al-habob, Jianqiang Lin, Octavia A Dobre, and Yindi Jing. 2024. Min-max latency minimization for energy-constrained multi-UAV mobile edge computing. IEEE Trans. Netw. Sci. Eng. 11, 5 (Sept. 2024), 4577–4590.
- Alam et al. (2019) Md Shahzad Alam, Natesha B. V., Ashwin T. S., and Ram Mohana Reddy Guddeti. 2019. UAV based cost-effective real-time abnormal event detection using edge computing. Multimed. Tools Appl. 78, 24 (Dec. 2019), 35119–35134.
- Ali and Ali (2022) Zainab H Ali and Hesham A Ali. 2022. EEOMA: End-to-end oriented management architecture for 6G-enabled drone communications. Peer Peer Netw. Appl. 15, 2 (March 2022), 1232–1254.
- Alioua et al. (2018) Ahmed Alioua, Sidi-Mohammed Senouci, Samira Moussaoui, Hichem Sedjelmaci, and Mohamed-Ayoub Messous. 2018. Efficient data processing in software-defined UAV-assisted vehicular networks: A sequential game approach. Wirel. Pers. Commun. 101, 4 (Aug. 2018), 2255–2286.
- Alkinani et al. (2021) Monagi H Alkinani, Abdulwahab Ali Almazroi, N Z Jhanjhi, and Navid Ali Khan. 2021. 5G and IoT based reporting and accident detection (RAD) system to deliver first aid box using unmanned aerial vehicle. Sensors (Basel) 21, 20 (Oct. 2021), 6905.
- Aloqaily et al. (2021) Moayad Aloqaily, Ouns Bouachir, Azzedine Boukerche, and Ismaeel Al Ridhawi. 2021. Design Guidelines for Blockchain-Assisted 5G-UAV Networks. IEEE Netw. 35, 1 (Jan. 2021), 64–71.
- Alqudsi et al. (2023) Yunes Sh Alqudsi, Radhwan A A Saleh, Murat Makaraci, and H Metin Ertunç. 2023. Enhancing aerial robots performance through robust hybrid control and metaheuristic optimization of controller parameters. Neural Comput. Appl. (Nov. 2023).
- Amin and Obermaisser (2018) Sarah Amin and Roman Obermaisser. 2018. A time-triggered scheduling algorithm for active diagnosis in heterogeneous distributed systems. In 2018 IEEE International Conference on Computational Science and Engineering (CSE) (Bucharest). IEEE.
- Amos et al. (2021) Peprah Amos, Pei Li, Wei Wu, and Baoyun Wang. 2021. Computation efficiency maximization for secure UAV-enabled mobile edge computing networks. Phys. Commun. 46, 101284 (June 2021), 101284.
- Anthropic (2024) Anthropic. 2024. Claude 3.5 [Large language model]. https://www.anthropic.com/claude
- Apostolopoulos et al. (2023) Pavlos Athanasios Apostolopoulos, Georgios Fragkos, Eirini Eleni Tsiropoulou, and Symeon Papavassiliou. 2023. Data offloading in UAV-assisted multi-access edge computing systems under resource uncertainty. IEEE Trans. Mob. Comput. 22, 1 (Jan. 2023), 175–190.
- Asheralieva and Niyato (2019) Alia Asheralieva and Dusit Niyato. 2019. Hierarchical game-theoretic and reinforcement learning framework for computational offloading in UAV-enabled mobile edge computing networks with multiple service providers. IEEE Internet Things J. 6, 5 (Oct. 2019), 8753–8769.
- Asheralieva and Niyato (2020) Alia Asheralieva and Dusit Niyato. 2020. Distributed dynamic resource management and pricing in the IoT systems with blockchain-as-a-service and UAV-enabled mobile edge computing. IEEE Internet Things J. 7, 3 (March 2020), 1974–1993.
- Asim and Abd El-Latif (2023) Muhammad Asim and Ahmed A Abd El-Latif. 2023. Intelligent computational methods for multi-unmanned aerial vehicle-enabled autonomous mobile edge computing systems. ISA Trans. 132 (Jan. 2023), 5–15.
- Asim et al. (2022) Muhammad Asim, Wali Khan Mashwani, Habib Shah, and Samir Brahim Belhaouari. 2022. An evolutionary trajectory planning algorithm for multi-UAV-assisted MEC system. Soft Comput. 26, 16 (Aug. 2022), 7479–7492.
- Asmer et al. (2024) Lukas Asmer, Roman Jaksche, Henry Pak, and Petra Kokus. 2024. A city-centric approach to estimate and evaluate global Urban Air Mobility demand. CEAS Aeronautical Journal (2024), 1–16.
- Avizienis et al. (2004) Algirdas Avizienis, J-C Laprie, Brian Randell, and Carl Landwehr. 2004. Basic concepts and taxonomy of dependable and secure computing. IEEE transactions on dependable and secure computing 1, 1 (2004), 11–33.
- Awada et al. (2021) Uchechukwu Awada, Jiankang Zhang, Sheng Chen, and Shuangzhi Li. 2021. Air-to-air collaborative learning: A multi-task orchestration in federated aerial computing. In 2021 IEEE 14th International Conference on Cloud Computing (CLOUD) (Chicago, IL, USA). IEEE.
- Awada et al. (2022) Uchechukwu Awada, Jiankang Zhang, Sheng Chen, and Shuangzhi Li. 2022. AirEdge: A dependency-aware multi-task orchestration in federated aerial computing. IEEE Trans. Veh. Technol. 71, 1 (Jan. 2022), 805–819.
- Awada et al. (2023) Uchechukwu Awada, Jiankang Zhang, Sheng Chen, Shuangzhi Li, and Shouyi Yang. 2023. EdgeDrones: Co-scheduling of drones for multi-location aerial computing missions. J. Netw. Comput. Appl. 215, 103632 (June 2023), 103632.
- Baccarelli et al. (2024) Enzo Baccarelli, Michele Scarpiniti, and Alireza Momenzadeh. 2024. Energy-minimizing 3D circular trajectory optimization of rotary-wing UAV under probabilistic path-loss in constrained hotspot environments. Veh. Commun. 46, 100730 (April 2024), 100730.
- Bai et al. (2023b) Haohao Bai, Zhineng Luo, Meihui Chen, and Senke Gui. 2023b. Energy saving research on auxiliary task unloading and resource allocation of UAV based on edge computing. In 2023 International Conference on Communications, Computing and Artificial Intelligence (CCCAI) (Shanghai, China). IEEE, 103–108.
- Bai et al. (2023a) Jing Bai, Xiaolin Chang, Ricardo J Rodríguez, Kishor S Trivedi, and Shupan Li. 2023a. Towards UAV-based MEC service chain resilience evaluation: A quantitative modeling approach. IEEE Trans. Veh. Technol. 72, 4 (April 2023), 5181–5194.
- Bakhtiari et al. (2020) Zohreh Bakhtiari, Rozita Jamili Oskouei, Mona Soleymani, and Akhtar Hussain Jalbani. 2020. Presenting an effective method to detect and track the broken path in VANET using UAVs. Wirel. Commun. Mob. Comput. 2020 (Aug. 2020), 1–12.
- Baktayan et al. (2022) Asrar Ahmed Baktayan, Ibrahim Ahmed Al-Baltah, and Abdul Azim Abd Ghani. 2022. Intelligent pricing model for task offloading in unmanned aerial vehicle mounted mobile edge computing for vehicular network. J. Commun. Softw. Syst. 18, 2 (2022), 111–123.
- Balamuralidhar et al. (2021) Navaneeth Balamuralidhar, Sofia Tilon, and Francesco Nex. 2021. MultEYE: Monitoring system for real-time vehicle detection, tracking and speed estimation from UAV imagery on edge-computing platforms. Remote Sens. (Basel) 13, 4 (Feb. 2021), 573.
- Bao et al. (2023) Lingyan Bao, Jia Luo, Yuyu Hao, Rongqian Zhang, Xianqi Zhang, Yunchun Zhang, and Mingxiong Zhao. 2023. Cruise duration minimization for UAV-and-basestation hybrid assisted thermal-aware MEC networks. In 2023 IEEE Wireless Communications and Networking Conference (WCNC) (Glasgow, United Kingdom). IEEE, 1–6.
- Barnawi et al. (2021) Ahmed Barnawi, Prateek Chhikara, Rajkumar Tekchandani, Neeraj Kumar, and Bander Alzahrani. 2021. Artificial intelligence-enabled Internet of Things-based system for COVID-19 screening using aerial thermal imaging. Future Gener. Comput. Syst. 124 (Nov. 2021), 119–132.
- Basharat et al. (2023) Mehak Basharat, Muhammad Naeem, and Alagan Anpalagan. 2023. Latency and cost minimization for task offloading with energy harvesting in UAV-MEC network. In 2023 IEEE Symposium on Wireless Technology & Applications (ISWTA) (Kuala Lumpur, Malaysia). IEEE, 103–108.
- Becker and Schalk (2024) Dennis Becker and Lukas Marcel Schalk. 2024. Toward robust and efficient communications for urban air mobility. CEAS Aeronautical Journal (2024), 1–28.
- Behjat et al. (2019) Amir Behjat, Steve Paul, and Souma Chowdhury. 2019. Learning reciprocal actions for cooperative collision avoidance in quadrotor unmanned aerial vehicles. Rob. Auton. Syst. 121, 103270 (Nov. 2019), 103270.
- Bekkouche et al. (2018) Oussama Bekkouche, Tarik Taleb, and Miloud Bagaa. 2018. UAVs traffic control based on multi-access edge computing. In 2018 IEEE Global Communications Conference (GLOBECOM) (Abu Dhabi, United Arab Emirates). IEEE.
- Bezziane et al. (2024) Mohamed Ben Bezziane, Siham Hasan, Bouziane Brik, Fathi Eltayeeb Abukhres, Ali Algaddafi, Amina Ben Bezziane, Ahmed Korichi, and Mohamed Redouane Kafi. 2024. Game theory-based UAV-cloud for service selection architecture in flying ad hoc networks. IEEE Open J. Veh. Technol. 5 (2024), 1692–1711.
- Bian et al. (2023) Sizhen Bian, Lukas Schulthess, Georg Rutishauser, Alfio Di Mauro, Luca Benini, and Michele Magno. 2023. ColibriUAV: An ultra-fast, energy-efficient neuromorphic edge processing UAV-platform with event-based and frame-based cameras. In 2023 9th International Workshop on Advances in Sensors and Interfaces (IWASI) (Monopoli (Bari), Italy). IEEE, 287–292.
- Bobronnikov et al. (2022) V T Bobronnikov, A V Starkov, N V Udalova, and I A Belousov. 2022. Smart onboard information and computing system for pilotless drones. Russ. Eng. Res. 42, 6 (June 2022), 630–633.
- Bouzidi et al. (2020) Zohra Bouzidi, Labib Sadek Terrissa, Noureddine Zerhouni, and Soheyb Ayad. 2020. QoS of cloud prognostic system: application to aircraft engines fleet. Eur. J Ind. Eng. 14, 1 (2020), 34.
- Bowley et al. (2019) Connor Bowley, Marshall Mattingly, Andrew Barnas, Susan Ellis-Felege, and Travis Desell. 2019. An analysis of altitude, citizen science and a convolutional neural network feedback loop on object detection in Unmanned Aerial Systems. J. Comput. Sci. 34 (May 2019), 102–116.
- Brito et al. (2021) Carlos Brito, Leonardo Silva, Gustavo Callou, Tuan Anh Nguyen, Dugki Min, Jae-Woo Lee, and Francisco Airton Silva. 2021. Offloading data through unmanned aerial vehicles: A dependability evaluation. Electronics (Basel) 10, 16 (Aug. 2021), 1916.
- Callegaro et al. (2021) Davide Callegaro, Marco Levorato, and Francesco Restuccia. 2021. SeReMAS: Self-resilient mobile autonomous systems through predictive edge computing. In 2021 18th Annual IEEE International Conference on Sensing, Communication, and Networking (SECON) (Rome, Italy). IEEE.
- Cao et al. (2021) Jin Cao, Chunhui Liu, Zanliang Dong, and Pei Wang. 2021. Effectiveness evaluation of UAV Ad hoc network in complex task environment. In 2021 International Conference on Communications, Information System and Computer Engineering (CISCE) (Beijing, China). IEEE.
- Chai et al. (2023) Jiahui Chai, Yun Meng, Wei Wang, Yue Lyu, Haochuan Yue, and Xinyi Liu. 2023. Task scheduling of computation-intensive graph jobs in UAV-assisted hybrid vehicular networks. Veh. Commun. 42, 100630 (Aug. 2023), 100630.
- Chang et al. (2024) Zhiwei Chang, Hanfeng Zhang, Yan Jia, Sihan Xu, Tong Li, and Zheli Liu. 2024. Low-cost fuzzing drone control system for configuration errors threatening flight safety in edge terminals. Comput. Commun. 220 (April 2024), 138–148.
- Chaudhry and Yanikomeroglu (2021) Aizaz U Chaudhry and Halim Yanikomeroglu. 2021. Laser intersatellite links in a starlink constellation: A classification and analysis. IEEE vehicular technology magazine 16, 2 (2021), 48–56.
- Chen et al. (2023) Guojun Chen, Xiaojing Yu, Neiwen Ling, and Lin Zhong. 2023. Typefly: Flying drones with large language model. arXiv preprint arXiv:2312.14950 (2023).
- Chen et al. (2022a) Haowei Chen, Liekang Zeng, Xiaoxi Zhang, and Xu Chen. 2022a. AdaDrone: Quality of navigation based neural adaptive scheduling for edge-assisted drones. In 2022 IEEE 42nd International Conference on Distributed Computing Systems (ICDCS) (Bologna, Italy). IEEE.
- Chen et al. (2022b) Lunyuan Chen, Rui Zhao, Ke He, Zichao Zhao, and Liseng Fan. 2022b. Intelligent ubiquitous computing for future UAV-enabled MEC network systems. Cluster Comput. 25, 4 (Aug. 2022), 2417–2427.
- Chen et al. (2024b) Peipei Chen, Lailong Luo, Deke Guo, Xueshan Luo, Xinyi Li, and Yuchen Sun. 2024b. Secure task offloading for rural area surveillance based on UAV-UGV collaborations. IEEE Trans. Veh. Technol. 73, 1 (Jan. 2024), 923–937.
- Chen et al. (2020a) Ping-Yang Chen, Jun-Wei Hsieh, Munkhjargal Gochoo, Ming-Ching Chang, Chien-Yao Wang, Yong-Sheng Chen, and Hong-Yuan Mark Liao. 2020a. Drone-based vehicle flow estimation and its application to traffic conflict hotspot detection at intersections. In 2020 IEEE International Conference on Image Processing (ICIP) (Abu Dhabi, United Arab Emirates). IEEE.
- Chen and Ma (2024) Xuefeng Chen and Rui Ma. 2024. Intelligent UAV planning for task-offloading with limited buffer and multiple computing servers. Phys. Commun. 62, 102240 (Feb. 2024), 102240.
- Chen et al. (2021a) Yali Chen, Bo Ai, Yong Niu, Hongliang Zhang, and Zhu Han. 2021a. Energy-constrained computation offloading in space-air-ground integrated networks using distributionally robust optimization. IEEE Trans. Veh. Technol. 70, 11 (Nov. 2021), 12113–12125.
- Chen et al. (2024a) Yali Chen, Min Liu, Bo Ai, Yuwei Wang, and Sheng Sun. 2024a. Adaptive bitrate video caching in UAV-assisted MEC networks based on distributionally robust optimization. IEEE Trans. Mob. Comput. (2024), 1–14.
- Chen et al. (2021b) Zhi Chen, Boyu Ning, Chong Han, Zhongbao Tian, and Shaoqian Li. 2021b. Intelligent Reflecting Surface Assisted Terahertz Communications toward 6G. arXiv:2104.02897 [cs.IT] https://arxiv.org/abs/2104.02897
- Chen et al. (2020b) Zhixiong Chen, Nan Xiao, and Dongsheng Han. 2020b. A multilevel mobile fog computing offloading model based on UAV-assisted and heterogeneous network. Wirel. Commun. Mob. Comput. 2020 (July 2020), 1–11.
- Chen et al. (2020c) Zhixiong Chen, Nan Xiao, and Dongsheng Han. 2020c. Multilevel task offloading and resource optimization of edge computing networks considering UAV relay and green energy. Appl. Sci. (Basel) 10, 7 (April 2020), 2592.
- Cheng et al. (2023) Ming Cheng, Canlin Zhu, Min Lin, Jun-Bo Wang, and Wei-Ping Zhu. 2023. An O-MAPPO scheme for joint computation offloading and resources allocation in UAV assisted MEC systems. Comput. Commun. 208 (Aug. 2023), 190–199.
- Cheng et al. (2024) Qing Cheng, Yazhe Wang, Wenjian He, and Yu Bai. 2024. Lightweight air-to-air unmanned aerial vehicle target detection model. Sci. Rep. 14, 1 (Jan. 2024), 2609.
- Choutri et al. (2022) Kheireddine Choutri, Mohand Lagha, Souham Meshoul, and Samiha Fadloun. 2022. Path planning and formation control for UAV-enabled Mobile Edge Computing network. Sensors (Basel) 22, 19 (Sept. 2022), 7243.
- Chu et al. (2021) Liju Chu, Xuejun Li, Jia Xu, Azadeh Ghari Neiat, and Xiao Liu. 2021. A holistic service provision strategy for drone-as-a-service in MEC-based UAV delivery. In 2021 IEEE International Conference on Web Services (ICWS) (Chicago, IL, USA). IEEE.
- Cyrille et al. (2024) Nkotto Akono Cyrille, Ndzana Jean Calvin, Mengata Mengounou Ghislain, and Sosso Mayi Olivier Thierry. 2024. Critical review on integrated real-time energy management strategy and digital twin applied to hybrid electric UAVs. Electrical Engineering (2024), 1–23.
- D’Agati et al. (2024) Luca D’Agati, Francesco Longo, Giovanni Merlino, Antonio Puliafito, and Giuseppe Tricomi. 2024. Enhancing UAV operational efficiency through cloud computing and autopilot system integration. In 2024 IEEE International Conference on Smart Computing (SMARTCOMP) (Osaka, Japan). IEEE.
- Dai et al. (2018) Rui Dai, Sneha Fotedar, Mohammadreza Radmanesh, and Manish Kumar. 2018. Quality-aware UAV coverage and path planning in geometrically complex environments. Ad Hoc Netw. 73 (May 2018), 95–105.
- Dai et al. (2024) Yu Dai, Jiaming Fu, Zhen Gao, and Lei Yang. 2024. Research on joint optimization of task offloading and UAV trajectory in mobile edge computing considering communication cost based on safe reinforcement learning. Appl. Sci. (Basel) 14, 6 (March 2024), 2635.
- Dai et al. (2022) Zhiqiang Dai, Gaochao Xu, Ziqi Liu, Jiaqi Ge, and Wei Wang. 2022. Energy saving strategy of UAV in MEC based on deep reinforcement learning. Future Internet 14, 8 (July 2022), 226.
- Damigos et al. (2023) Gerasimos Damigos, Achilleas Santi Seisa, Sumeet Gajanan Satpute, Tore Lindgren, and George Nikolakopoulos. 2023. A resilient framework for 5G-edge-connected UAVs based on switching edge-MPC and onboard-PID control. In 2023 IEEE 32nd International Symposium on Industrial Electronics (ISIE) (Helsinki, Finland). IEEE.
- Dao et al. (2021) Van-Lan Dao, Le-Nam Hoang, Svetlana Girs, and Elisabeth Uhlemann. 2021. Defeating jamming using outage performance aware joint power allocation and access point placement in uplink pairwise NOMA. IEEE Open J. Commun. Soc. 2 (2021), 1957–1979.
- De Paula Soares et al. (2023) Lucas De Paula Soares, Fabíola M C De Oliveira, Carlos A Kamienski, and Luiz F Bittencourt. 2023. Drone edge management system (DREMS): Sequencing drone takeoff and landing. In 2023 10th International Conference on Future Internet of Things and Cloud (FiCloud) (Marrakesh, Morocco). IEEE.
- Deng et al. (2022) Liping Deng, Hong Jiang, He Xiao, Qiuyun Zhang, Ying Luo, Chun Wu, and Changqing Ye. 2022. Completion time minimization for multi-antenna UAV-enabled data collection in uncorrelated Rician fading. Veh. Commun. 37, 100501 (Oct. 2022), 100501.
- Deng et al. (2021) Xiaoheng Deng, Yajun Liu, Congxu Zhu, and Honggang Zhang. 2021. Air–Ground Surveillance Sensor Network based on edge computing for target tracking. Comput. Commun. 166 (Jan. 2021), 254–261.
- Dogea et al. (2023) Ramona Dogea, Xiu T Yan, and Richard Millar. 2023. Implementation of an edge-fog-cloud computing IoT architecture in aircraft components. MRS Commun. 13, 3 (May 2023), 416–424.
- Dou et al. (2024) Wanchun Dou, Bowen Liu, Jirun Duan, Fei Dai, Lianyong Qi, and Xiaolong Xu. 2024. SeeMe: An intelligent edge server selection method for location‐aware business task computing over IIoT. Softw. Pract. Exp. 54, 10 (Oct. 2024), 1939–1956.
- El Sayed et al. (2023) O H El Sayed, S M Youssef, and O M Ismail. 2023. Swarm unmanned aerial vehicles (uavs)-based Fog Computing platform supporting internet of things applications. ISPRS Ann. Photogramm. Remote Sens. Spat. Inf. Sci. X-1/W1-2023 (Dec. 2023), 1041–1048.
- Esho et al. (2024) Adeola Ona-Olapo Esho, Tosin Daniel Iluyomade, Tosin Michael Olatunde, and Osayi Philip Igbinenikaro. 2024. A comprehensive review of energy-efficient design in satellite communication systems. International Journal of Engineering Research Updates 6, 02 (2024), 013–025.
- Falcão et al. (2023) Marcos Falcão, Caio Souza, Andson Balieiro, and Kelvin Dias. 2023. Dynamic resource allocation for URLLC in UAV-enabled multi-access edge computing. In 2023 Joint European Conference on Networks and Communications & 6G Summit (EuCNC/6G Summit) (Gothenburg, Sweden). IEEE.
- Falcão et al. (2024) Marcos Falcão, Caio Bruno Souza, Andson Balieiro, and Kelvin Dias. 2024. Resource allocation for UAV-enabled multi-access edge computing. J. Supercomput. 80, 15 (Oct. 2024), 22770–22802.
- Fan et al. (2023) Xiang Fan, Zhipan Wang, Hua Zhang, Huan Liu, Zhuoyi Jiang, and Xianghe Liu. 2023. Large-scale rice mapping based on Google earth engine and multi-source remote sensing images. J. Ind. Soc. Remote Sens. 51, 1 (Jan. 2023), 93–102.
- Faraci et al. (2023) Giuseppe Faraci, Santi Agatino Rizzo, and Giovanni Schembra. 2023. Green edge intelligence for smart management of a FANET in disaster-recovery scenarios. IEEE Trans. Veh. Technol. 72, 3 (March 2023), 3819–3831.
- Farrukh and West (2023) Anam Farrukh and Richard West. 2023. FlyOS: rethinking integrated modular avionics for autonomous multicopters. Real-Time Syst. 59, 2 (June 2023), 256–301.
- Fatima et al. (2024) Nida Fatima, Paresh Saxena, and Giovanni Giambene. 2024. Deep reinforcement learning based computation offloading for xURLLC services with UAV-assisted IoT-based multi-access edge computing system. Wirel. Netw. 30, 9 (Dec. 2024), 7275–7291.
- Feng et al. (2023) Aijing Feng, Chin Nee Vong, Jing Zhou, Lance S Conway, Jianfeng Zhou, Earl D Vories, Kenneth A Sudduth, and Newell R Kitchen. 2023. Developing an image processing pipeline to improve the position accuracy of single UAV images. Comput. Electron. Agric. 206, 107650 (March 2023), 107650.
- Feng et al. (2024) Xiao Feng, Ru Gao, Fujie Tang, and Jianjun Luo. 2024. Multi-UAV Collaborative Reconnaissance Based on Multi Agent Deep Reinforcement Learning. 65–70. doi:10.1109/BigDIA63733.2024.10808323
- Fesenko et al. (2023) Herman Fesenko, Oleg Illiashenko, Vyacheslav Kharchenko, Ihor Kliushnikov, Olga Morozova, Anatoliy Sachenko, and Stanislav Skorobohatko. 2023. Flying sensor and edge network-based Advanced Air Mobility systems: Reliability analysis and applications for urban monitoring. Drones 7, 7 (June 2023), 409.
- Fujii et al. (2017) Satoshi Fujii, Kazutami Arimoto, Kohei Hiranaka, Seichio Miyagi, Motoki Kyan, Shouichi Tanifuji, Tansuriyavong Suriyon, Zacharie Mbaitiga, Noriaki Yoshikawa, and Kensaku Kinoshita. 2017. Integration of drones’ communication into an ITS network. In 2017 International Conference on Intelligent Informatics and Biomedical Sciences (ICIIBMS) (Okinawa). IEEE.
- Gallego-Madrid et al. (2020) Jorge Gallego-Madrid, Alejandro Molina-Zarca, Ramon Sanchez-Iborra, Jorge Bernal-Bernabe, José Santa, Pedro Miguel Ruiz, and Antonio F Skarmeta-Gómez. 2020. Enhancing extensive and remote LoRa deployments through MEC-powered drone gateways. Sensors (Basel) 20, 15 (July 2020), 4109.
- Gao et al. (2020) Ang Gao, Tianli Geng, Yansu Hu, Wei Liang, and Weijun Duan. 2020. Decentralized continuous game for task offloading in UAV cloud. In 2020 29th Wireless and Optical Communications Conference (WOCC) (Newark, NJ, USA). IEEE.
- Gao et al. (2019) Ang Gao, Yansu Hu, Wei Liang, Yizhi Lin, Lixin Li, and Xu Li. 2019. A QoE-oriented scheduling scheme for energy-efficient computation offloading in UAV cloud system. IEEE Access 7 (2019), 68656–68668.
- Gao et al. (2021) Ang Gao, Qi Wang, Wei Liang, and Zhiguo Ding. 2021. Game combined multi-agent reinforcement learning approach for UAV assisted offloading. IEEE Trans. Veh. Technol. 70, 12 (Dec. 2021), 12888–12901.
- Gao et al. (2024) Jiakun Gao, Xiaolong Xu, Lianyong Qi, Wanchun Dou, Xiaoyu Xia, and Xiaokang Zhou. 2024. Distributed computation offloading and power control for UAV-enabled Internet of Medical Things. ACM Trans. Internet Technol. (March 2024).
- Gao et al. (2022) Peng Gao, Kangbeen Lee, Lukas Wiku Kuswidiyanto, Seung-Hwa Yu, Kai Hu, Gaotian Liang, Yufeng Chen, Weixing Wang, Fei Liao, Yu Seok Jeong, Moon-Seok Jeon, Inchan Choi, and Xiongzhe Han. 2022. Dynamic beehive detection and tracking system based on YOLO V5 and unmanned aerial vehicle. J. Biosyst. Eng. 47, 4 (Dec. 2022), 510–520.
- Garge and Balakrishna (2018) Gopi Krishna Garge and Chitra Balakrishna. 2018. Unmanned aerial vehicles (UAVs) as on-demand QoS enabler for multimedia applications in smart cities. In 2018 International Conference on Innovation and Intelligence for Informatics, Computing, and Technologies (3ICT) (Sakhier, Bahrain). IEEE.
- Geng et al. (2024) Liwei Geng, Hongbo Zhao, and Changming Zou. 2024. Stackelberg game-based computation offloading and pricing in UAV assisted vehicular networks. IEEE Trans. Reliab. (2024), 1–15.
- Ghosh and Kuila (2023) Santanu Ghosh and Pratyay Kuila. 2023. Efficient offloading in disaster-affected areas using unmanned aerial vehicle-assisted mobile edge computing: A gravitational search algorithm-based approach. Int. J. Disaster Risk Reduct. 97, 104067 (Oct. 2023), 104067.
- Goel et al. (2021) Muskan Goel, Rajoshi Sarkar, and S Singh. 2021. Estimation of stability and control derivatives using simulated flight data. In 2021 11th International Conference on Cloud Computing, Data Science & Engineering (Confluence) (Noida, India). IEEE.
- Gokulakrishnan et al. (2023) S Gokulakrishnan, Muhammad Aslam Jarwar, Mohammed Hasan Ali, M M Kamruzzaman, Iyapparaja Meenakshisundaram, Mustafa Musa Jaber, and R Lakshmana Kumar. 2023. Maliciously roaming person’s detection around hospital surface using intelligent cloud-edge based federated learning. J. Comb. Optim. 45, 1 (Jan. 2023).
- Grasso et al. (2022a) Christian Grasso, Raoul Raftopoulos, and Giovanni Schembra. 2022a. Slicing a FANET for heterogeneous delay-constrained applications. Comput. Commun. 195 (Nov. 2022), 362–375.
- Grasso et al. (2022b) Christian Grasso, Raoul Raftopoulos, Giovanni Schembra, and Salvatore Serrano. 2022b. H-HOME: A learning framework of federated FANETs to provide edge computing to future delay-constrained IoT systems. Comput. Netw. 219, 109449 (Dec. 2022), 109449.
- Grasso and Schembra (2019) Christian Grasso and Giovanni Schembra. 2019. A fleet of MEC UAVs to extend a 5G network slice for video monitoring with low-latency constraints. J. Sens. Actuator Netw. 8, 1 (Jan. 2019), 3.
- Gu et al. (2022) Xiaohui Gu, Guoan Zhang, Mingxing Wang, Wei Duan, Miaowen Wen, and Pin-Han Ho. 2022. UAV-aided energy-efficient edge computing networks: Security offloading optimization. IEEE Internet Things J. 9, 6 (March 2022), 4245–4258.
- Guo et al. (2022b) Jinlong Guo, Lizhi Wang, and Xiaohong Wang. 2022b. A group maintenance method of drone swarm considering system mission reliability. Drones 6, 10 (Sept. 2022), 269.
- Guo et al. (2022a) Zhiqi Guo, Jinzhou Deng, Tiao Tan, and Ming Zhao. 2022a. Trajectory optimization algorithm of UAV based on pointer network. In 2022 IEEE 4th International Conference on Civil Aviation Safety and Information Technology (ICCASIT) (Dali, China). IEEE.
- Gupta and Gupta (2022) Akshita Gupta and Sachin Kumar Gupta. 2022. Intelligent collaboration of multi-agent flying UAV-fog networking for better QoS. In 2022 International Conference on Electrical, Computer, Communications and Mechatronics Engineering (ICECCME) (Maldives, Maldives). IEEE.
- Gupta et al. (2021) Rajesh Gupta, Dakshita Reebadiya, and Sudeep Tanwar. 2021. 6G-enabled edge intelligence for ultra -reliable low latency applications: Vision and mission. Comput. Stand. Interfaces 77, 103521 (Aug. 2021), 103521.
- Guruprasad et al. (2024) Kapilesh Guruprasad, Dhanush Sridhar, Gokula Kannan Venkatesan Poovazhaki, and Mary Anita Rajam Viswasam. 2024. A priority aware local mutual exclusion algorithm for flying ad hoc networks. Computing 106, 3 (March 2024), 981–1012.
- Haber et al. (2019) Elie El Haber, Hyame Assem Alameddine, Chadi Assi, and Sanaa Sharafeddine. 2019. A reliability-aware computation offloading solution via UAV-mounted cloudlets. In 2019 IEEE 8th International Conference on Cloud Networking (CloudNet) (Coimbra, Portugal). IEEE.
- Haber et al. (2021) Elie El Haber, Hyame Assem Alameddine, Chadi Assi, and Sanaa Sharafeddine. 2021. UAV-aided ultra-reliable low-latency computation offloading in future IoT networks. IEEE Trans. Commun. 69, 10 (Oct. 2021), 6838–6851.
- Hadi and Ghazizadeh (2023) Majid Hadi and Reza Ghazizadeh. 2023. Computation bits enhancement for IRS-assisted multi-UAV wireless powered mobile edge computing systems. Veh. Commun. 43, 100656 (Oct. 2023), 100656.
- Hadjkouider et al. (2023) Abdessalam Mohammed Hadjkouider, Chaker Abdelaziz Kerrache, Ahmed Korichi, Yesin Sahraoui, and Carlos T Calafate. 2023. Stackelberg game approach for service selection in UAV networks. Sensors (Basel) 23, 9 (April 2023).
- Han et al. (2021) Pengcheng Han, Chenglie Du, Jinchao Chen, Fuyuan Ling, and Xiaoyan Du. 2021. Cost and makespan scheduling of workflows in clouds using list multiobjective optimization technique. J. Syst. Arch. 112, 101837 (Jan. 2021), 101837.
- Hang et al. (2020) Bingqing Hang, Biling Zhang, Li Wang, Jingjing Wang, Yong Ren, and Zhu Han. 2020. A user association policy for UAV-aided time-varying vehicular networks with MEC. In 2020 IEEE Wireless Communications and Networking Conference (WCNC) (Seoul, Korea (South)). IEEE.
- Hao et al. (2024) Jialin Hao, Rola Naja, and Djamal Zeghlache. 2024. Adaptive federated reinforcement learning for critical realtime communications in UAV assisted vehicular networks. Comput. Netw. 247, 110456 (June 2024), 110456.
- Hashim (2024) Hashim A Hashim. 2024. Advances in UAV avionics systems architecture, classification and integration: A comprehensive review and future perspectives. Results in Engineering (2024), 103786.
- He et al. (2023) Saifei He, Ming Cheng, Yijin Pan, Min Lin, and Wei-Ping Zhu. 2023. Distributed access and offloading scheme for multiple UAVs assisted MEC networks. In 2023 IEEE 98th Vehicular Technology Conference (VTC2023-Fall) (Hong Kong, Hong Kong). IEEE, 1–5.
- Heidari et al. (2023) Arash Heidari, Nima Jafari Navimipour, Mehmet Unal, and Guodao Zhang. 2023. Machine learning applications in internet-of-drones: Systematic review, recent deployments, and open issues. Comput. Surveys 55, 12 (2023), 1–45.
- Hines and Bekemeyer (2023) Derrick Hines and Philipp Bekemeyer. 2023. Graph neural networks for the prediction of aircraft surface pressure distributions. Aerosp. Sci. Technol. 137, 108268 (June 2023), 108268.
- Hirota et al. (2020) Yusuke Hirota, Ittetsu Taniguchi, and Takao Onoye. 2020. Parallelization of local path planning for high reliable autonomous drones. In 2020 International SoC Design Conference (ISOCC) (Yeosu, Korea (South)). IEEE.
- Hong et al. (2022) Qingqing Hong, Ling Jiang, Zhenghua Zhang, Shu Ji, Chen Gu, Wei Mao, Wenxi Li, Tao Liu, Bin Li, and Changwei Tan. 2022. A lightweight model for wheat ear Fusarium head blight detection based on RGB images. Remote Sens. (Basel) 14, 14 (July 2022), 3481.
- Hossain et al. (2017) Murad Hossain, Sameer Alam, and Hussein Abbass. 2017. A dynamic multi-commodity flow optimization algorithm for estimating airport network capacity. In Lecture Notes in Electrical Engineering. Springer Japan, Tokyo, 205–220.
- Hossain et al. (2022) Md Sahabul Hossain, Fisayo Sangoleye, Oshan Poudyal, and Eirini Eleni Tsiropoulou. 2022. Network economics-enabled edge computing in UAV-assisted public safety systems. In 2022 18th International Conference on Distributed Computing in Sensor Systems (DCOSS) (Marina del Rey, Los Angeles, CA, USA). IEEE.
- Hossen et al. (2021) Md Abir Hossen, Prasoon K Diwakar, and Shankarachary Ragi. 2021. Total nitrogen estimation in agricultural soils via aerial multispectral imaging and LIBS. Scientific Reports 11, 1 (2021), 12693.
- Hou et al. (2020b) Xiangwang Hou, Zhiyuan Ren, Jingjing Wang, Shuya Zheng, Wenchi Cheng, and Hailin Zhang. 2020b. Distributed fog computing for latency and reliability guaranteed swarm of drones. IEEE Access 8 (2020), 7117–7130.
- Hou et al. (2020a) Xiangwang Hou, Zhiyuan Ren, Jingjing Wang, Shuya Zheng, and Hailin Zhang. 2020a. Latency and reliability oriented collaborative optimization for multi-UAV aided mobile edge computing system. In IEEE INFOCOM 2020 - IEEE Conference on Computer Communications Workshops (INFOCOM WKSHPS) (Toronto, ON, Canada). IEEE.
- Hu and Seiler (2015) Bin Hu and Peter Seiler. 2015. Pivotal decomposition for reliability analysis of fault tolerant control systems on unmanned aerial vehicles. Reliab. Eng. Syst. Saf. 140 (Aug. 2015), 130–141.
- Hu et al. (2024) Hao Hu, Yulin Chen, Bao Peng, Zhenjun Li, Qibao Wu, Ziran Lin, and Xiaolei Wang. 2024. Cooperative positioning of UAV internet of things based on optimization algorithm. Wirel. Netw. 30, 5 (July 2024), 4495–4505.
- Huang et al. (2019) Hailong Huang, Andrey V Savkin, Ming Ding, and Mohamed Ali Kaafar. 2019. Optimized deployment of drone base station to improve user experience in cellular networks. J. Netw. Comput. Appl. 144 (Oct. 2019), 49–58.
- Huang et al. (2024b) Meng Huang, Hanming Li, Yina Zhou, Ting Ma, Jinshan Su, and Haibo Zhou. 2024b. A UAV aided lightweight target information collection and detection approach. Peer Peer Netw. Appl. 17, 3 (May 2024), 1667–1681.
- Huang and Wang (2024) Xianfei Huang and Gaocai Wang. 2024. Optimization for total energy consumption of drone inspection based on distance-constrained capacitated vehicle routing problem: A study in wind farm. Expert Syst. Appl. 255, 124880 (Dec. 2024), 124880.
- Huang et al. (2024c) Yingbing Huang, Lily Jiaxin Wan, Hanchen Ye, Manvi Jha, Jinghua Wang, Yuhong Li, Xiaofan Zhang, and Deming Chen. 2024c. New solutions on LLM acceleration, optimization, and application. In Proceedings of the 61st ACM/IEEE Design Automation Conference. 1–4.
- Huang et al. (2024a) Zhenqi Huang, Zhufang Kuang, Siyu Lin, Fen Hou, and Anfeng Liu. 2024a. Energy-efficient joint trajectory and reflecting design in IRS-enabled UAV edge computing. IEEE Internet Things J. 11, 12 (June 2024), 21872–21884.
- Huda and Moh (2022) SM Asiful Huda and Sangman Moh. 2022. Survey on computation offloading in UAV-Enabled mobile edge computing. Journal of Network and Computer Applications 201 (2022), 103341.
- Hui et al. (2024) Min Hui, Jian Chen, Long Yang, Lu Lv, Hai Jiang, and Naofal Al-Dhahir. 2024. UAV-assisted mobile edge computing: Optimal design of UAV altitude and task offloading. IEEE Trans. Wirel. Commun. 23, 10 (Oct. 2024), 13633–13647.
- Islam et al. (2019) Nahina Islam, Kandeepan Sithamparanathan, Karina Gomez Chavez, James Scott, and Hamid Eltom. 2019. Energy efficient and delay aware ternary-state transceivers for aerial base stations. Digit. Commun. Netw. 5, 1 (Feb. 2019), 40–50.
- Izhikevich et al. (2024) Liz Izhikevich, Reese Enghardt, Te-Yuan Huang, and Renata Teixeira. 2024. A Global Perspective on the Past, Present, and Future of Video Streaming over Starlink. Proceedings of the ACM on Measurement and Analysis of Computing Systems 8, 3 (2024), 1–22.
- Javaid et al. (2024) Shumaila Javaid, Hamza Fahim, Bin He, and Nasir Saeed. 2024. Large language models for uavs: Current state and pathways to the future. IEEE Open Journal of Vehicular Technology (2024).
- Javanmardi et al. (2024) Saeed Javanmardi, Georgia Sakellari, Mohammad Shojafar, and Antonio Caruso. 2024. Why it does not work? Metaheuristic task allocation approaches in Fog-enabled Internet of Drones. Simul. Model. Pract. Theory 133, 102913 (May 2024), 102913.
- Jennings et al. (2022) Joseph Jennings, Jean Pan, Lucas Overbey, and Jamie Lyle. 2022. Comparison of onboard processors for real-time target identification in unmanned aircraft systems. In Real-Time Image Processing and Deep Learning 2022 (Orlando, United States), Nasser Kehtarnavaz and Matthias F Carlsohn (Eds.). SPIE.
- Jeong et al. (2016) Hyeok-June Jeong, Jeong Dan Choi, and Young-Guk Ha. 2016. Vision based displacement detection for stabilized UAV control on cloud server. Mob. Inf. Syst. 2016 (2016), 1–11.
- Jeong et al. (2020) Hyeok-June Jeong, Suh-Yong Choi, Sung-Su Jang, and Young-Guk Ha. 2020. Probability machine-learning-based communication and operation optimization for cloud-based UAVs. J. Supercomput. 76, 10 (Oct. 2020), 8101–8117.
- Jeong et al. (2017) H.-J Jeong, M.-J Lee, C.E. Lee, S.-H Kim, and Young-Guk Ha. 2017. Machine learning-based real-time anomaly detection for unmanned aerial vehicles with a cloud server. Journal of Internet Technology 18 (01 2017), 823–832. doi:10.6138/JIT.2017.18.4.20160130b
- Ji et al. (2024) Baofeng Ji, Ying Wang, Weixing Wang, Shahid Mumtaz, and Charalampos Tsimenidis. 2024. Outage analysis of optimal UAV cooperation with IRS via energy harvesting enhancement assisted computational offloading. Comput. Model. Eng. Sci. 138, 2 (2024), 1885–1905.
- Jiang and He (2023) Lei Jiang and Jinsong He. 2023. AFDX network system simulation design based on OPNET. In 2023 3rd International Conference on Electronic Information Engineering and Computer (EIECT) (Shenzhen, China). IEEE.
- Jiang et al. (2019) Zhi Jiang, Tingting Yang, Lin Zhou, Yuqing Yuan, and Hailong Feng. 2019. Maritime search and rescue networking based on multi-agent cooperative communication. J. Commun. Inf. Netw. 4, 1 (March 2019), 42–53.
- Jo et al. (2018) Jun Jo, Zahra Jadidi, and Bela Stantic. 2018. A drone-based building inspection system using software-agents. In Studies in Computational Intelligence. Springer International Publishing, Cham, 115–121.
- Ju et al. (2023) Tao Ju, Linjuan Li, Jiuyuan Huo, and Qinan Li. 2023. Fine-grained task offloading for UAV-assisted mobile edge computing. In 2023 9th International Conference on Computer and Communications (ICCC) (Chengdu, China). IEEE.
- Ju et al. (2024) Tao Ju, Linjuan Li, Shuai Liu, and Yu Zhang. 2024. A multi-UAV assisted task offloading and path optimization for mobile edge computing via multi-agent deep reinforcement learning. J. Netw. Comput. Appl. 229, 103919 (Sept. 2024), 103919.
- Jung et al. (2018) Hyeok-June Jung, Kyeong-Sik Park, Cheol-Jin Kim, and Young-Guk Ha. 2018. Formal modeling and verification of serial communication for autonomous vehicles. In 2018 IEEE International Conference on Big Data and Smart Computing (BigComp) (Shanghai). IEEE.
- Jung et al. (2023a) Sooyeob Jung, Seongah Jeong, Jinkyu Kang, and Joonhyuk Kang. 2023a. Marine IoT systems with space–air–sea integrated networks: Hybrid LEO and UAV edge computing. IEEE Internet Things J. 10, 23 (Dec. 2023), 20498–20510.
- Jung et al. (2023b) Soyi Jung, Jae-Hyun Kim, David Mohaisen, and Joongheon Kim. 2023b. Truthful and performance-optimal computation outsourcing for aerial surveillance platforms via learning-based auction. Comput. Netw. 225, 109651 (April 2023), 109651.
- Kabashkin (2023) Igor Kabashkin. 2023. Availability of services in wireless sensor network with aerial base station placement. J. Sens. Actuator Netw. 12, 3 (May 2023), 39.
- Kaleem et al. (2019) Zeeshan Kaleem, Muhammad Yousaf, Aamir Qamar, Ayaz Ahmad, Trung Q Duong, Wan Choi, and Abbas Jamalipour. 2019. UAV-empowered disaster-resilient edge architecture for delay-sensitive communication. IEEE Netw. 33, 6 (Nov. 2019), 124–132.
- Karam et al. (2024) Sana Nasim Karam, Kashif Bilal, Junaid Shuja, Latif U Khan, Muhammad Bilal, and Muhammad Khurram Khan. 2024. Intelligent IoT- and UAV-Assisted Architecture for Pipeline Monitoring in OGI. IT Prof. 26, 3 (May 2024), 46–54.
- Kaymaz et al. (2024) Mehmetcan Kaymaz, Recep Ayzit, Onur Akgün, Kamil Canberk Atik, Mustafa Erdem, Baris Yalcin, Gürkan Cetin, and Nazım Kemal Ure. 2024. Trading-off safety with agility using deep pose error estimation and reinforcement learning for perception-driven UAV motion planning. J. Intell. Robot. Syst. 110, 2 (March 2024).
- Khan et al. (2022) Abdullah Ayub Khan, Asif Ali Laghari, Thippa Reddy Gadekallu, Zaffar Ahmed Shaikh, Abdul Rehman Javed, Mamoon Rashid, Vania V Estrela, and Alexey Mikhaylov. 2022. A drone-based data management and optimization using metaheuristic algorithms and blockchain smart contracts in a secure fog environment. Comput. Electr. Eng. 102, 108234 (Sept. 2022), 108234.
- Khan et al. (2021) Muhammad Asghar Khan, Tanweer Ahmad Cheema, Insaf Ullah, Fazal Noor, Sajjad Ur Rehman, Ateeq Ur Rehman, and Muhammad Adnan Aziz. 2021. A dual-mode medium access control mechanism for UAV-enabled Intelligent Transportation System. Mob. Inf. Syst. 2021 (Feb. 2021), 1–13.
- Khanna et al. (2019) Ashish Khanna, Joel J P C Rodrigues, Naman Gupta, Abhishek Swaroop, Deepak Gupta, Kashif Saleem, and Victor Hugo C de Albuquerque. 2019. A mutual exclusion algorithm for flying Ad Hoc networks. Comput. Electr. Eng. 76 (June 2019), 82–93.
- Khatiri et al. (2025) Sajad Khatiri, Fatemeh Mohammadi Amin, Sebastiano Panichella, and Paolo Tonella. 2025. When Uncertainty Leads to Unsafety: Empirical Insights into the Role of Uncertainty in Unmanned Aerial Vehicle Safety. arXiv preprint arXiv:2501.08908 (2025).
- Khoramnejad et al. (2024) Fahime Khoramnejad, Aisha Syed, W Sean Kennedy, and Melike Erol-Kantarci. 2024. Energy and delay aware general task dependent offloading in UAV-aided smart farms. IEEE Trans. Netw. Serv. Manag. 21, 5 (Oct. 2024), 5033–5048.
- Kilbourne et al. (2021) Todd Kilbourne, Frederick Wieland, Miro Lehky, Bryan Trainum, and Matt Underwood. 2021. Toward a cloud-based flight management system. In 2021 IEEE/AIAA 40th Digital Avionics Systems Conference (DASC) (San Antonio, TX, USA). IEEE.
- Kim et al. (2024b) Doyoung Kim, Seongah Jeong, and Jinkyu Kang. 2024b. Energy-efficient secure offloading system designed via UAV-mounted intelligent reflecting surface for resilience enhancement. IEEE Internet Things J. 11, 3 (Feb. 2024), 3768–3778.
- Kim et al. (2024a) Minsoo Kim, Jalel Ben-Othman, and Hyunbum Kim. 2024a. Reconfigurable UAV-aided 3D sustainable surveillance in classified air-spaces. Veh. Commun. 45, 100728 (Feb. 2024), 100728.
- Kim (2024) Sungwook Kim. 2024. Hierarchical aerial offload computing algorithm based on the Stackelberg-evolutionary game model. Comput. Netw. 245, 110348 (May 2024), 110348.
- Kim and Kim (2021) Seung-Yeon Kim and Yi-Kang Kim. 2021. An energy efficient UAV-based edge computing system with reliability guarantee for mobile ground nodes. Sensors (Basel) 21, 24 (Dec. 2021), 8264.
- Koerkamp et al. (2019) N W Klein Koerkamp, C Borst, Max Mulder, and M M van Paassen. 2019. Supporting humans in solving multi-UAV dynamic vehicle routing problems. IFAC-PapersOnLine 52, 19 (2019), 359–364.
- Kong et al. (2022a) Lingjun Kong, Bing Chen, and Feng Hu. 2022a. Blockchain-assisted adaptive reconfiguration method for trusted UAV network. Electronics (Basel) 11, 16 (Aug. 2022), 2549.
- Kong et al. (2022b) Peng Kong, Bo Li, Yonghao Wang, Xiao Huang, Kaibo Shi, Dong Ma, Bo Ran, and Jitao Huang. 2022b. Multi-UAV cooperative computational delay and energy consumption modeling and DDPG optimization. In 2022 IEEE 8th International Conference on Computer and Communications (ICCC) (Chengdu, China). IEEE.
- Kong et al. (2022c) Peng Kong, Bo Li, Yonghao Wang, Xiao Huang, Kaibo Shi, Dong Ma, Bo Ran, and Jitao Huang. 2022c. Multi-UAV cooperative computational latency modeling and DDPG optimization. In 2022 IEEE 22nd International Conference on Communication Technology (ICCT) (Nanjing, China). IEEE.
- Kostoglotov et al. (2024) Andrey A Kostoglotov, Vladimir O Zekhtser, Anton S Penkov, and Sergey V Lazarenko. 2024. Synthesis of a measurement procedure for estimating the orientation of a small unmanned aerial vehicle under changes in the status of measurement results. Meas. Tech. 66, 12 (March 2024), 897–905.
- Kramer and Stubberud (2021) Kathleen A Kramer and Stephen C Stubberud. 2021. Determining impact of navigation errors on mission capabilities. In 2021 IEEE/AIAA 40th Digital Avionics Systems Conference (DASC) (San Antonio, TX, USA). IEEE.
- Krishnan et al. (2021) Srivatsan Krishnan, Behzad Boroujerdian, William Fu, Aleksandra Faust, and Vijay Janapa Reddi. 2021. Air Learning: a deep reinforcement learning gym for autonomous aerial robot visual navigation. Mach. Learn. 110, 9 (Sept. 2021), 2501–2540.
- Kumar and Jannesari (2023) Chandan Kumar and Ali Jannesari. 2023. Optimal deep learning model for UAVs: A case study. In 2023 26th International Symposium on Wireless Personal Multimedia Communications (WPMC) (Tampa, FL, USA). IEEE, 51–57.
- Kumar et al. (2021) S V Aruna Kumar, Ehsan Yaghoubi, Abhijit Das, B S Harish, and Hugo Proenca. 2021. The P-DESTRE: A fully annotated dataset for pedestrian detection, tracking, and short/long-term re-identification from aerial devices. IEEE Trans. Inf. Forensics Secur. 16 (2021), 1696–1708.
- La Salandra et al. (2024) Marco La Salandra, Stefano Nicotri, Giacinto Donvito, Alessandro Italiano, Rosa Colacicco, Giorgia Miniello, Isabella Lapietra, Rodolfo Roseto, Pierfrancesco Dellino, and Domenico Capolongo. 2024. A paradigm shift in processing large UAV image datasets for emergency management of natural hazards. Int. J. Appl. Earth Obs. Geoinf. 132, 103996 (Aug. 2024), 103996.
- Laroui et al. (2021) Mohammed Laroui, Hatem Ibn-Khedher, Hassine Moungla, and Hossam Afifi. 2021. Autonomous UAV aided vehicular edge computing for service offering. In 2021 IEEE Global Communications Conference (GLOBECOM) (Madrid, Spain). IEEE.
- Lee et al. (2022a) Chao-Yang Lee, Huan-Jung Lin, Ming-Yuan Yeh, and Jer Ling. 2022a. Effective remote sensing from the Internet of drones through flying control with lightweight multitask learning. Appl. Sci. (Basel) 12, 9 (May 2022), 4657.
- Lee et al. (2022b) Minjae Lee, Sung Gyun Shin, Seungsoo Jang, Woosung Cho, Sungkyum Kim, Sangsoo Han, Chanho Choi, Jooyeon Kim, Youngmin Kim, and Song Hyun Kim. 2022b. Visual-based landing guidance system of UAV with deep learning technique for environments of visual-detection impairment. Int. J. Control Autom. Syst. 20, 5 (May 2022), 1735–1744.
- Lei et al. (2023) Yi Lei, Xiaochong Tong, Chunping Qiu, Yuekun Sun, Jiayi Tang, Congzhou Guo, and He Li. 2023. Onboard data management approach based on a discrete grid system for multi-UAV cooperative image localization. IEEE Trans. Geosci. Remote Sens. 61 (2023), 1–17.
- Li et al. (2024e) Bin Li, Rongrong Yang, Lei Liu, Junyi Wang, Ning Zhang, and Mianxiong Dong. 2024e. Robust Computation Offloading and Trajectory Optimization for Multi-UAV-Assisted MEC: A Multiagent DRL Approach. IEEE Internet Things J. 11, 3 (Feb. 2024), 4775–4786.
- Li et al. (2020b) Chengfang Li, Xiaojie Fang, Xiaokang Zhou, Lin Mei, and Xuejun Sha. 2020b. Enhanced signalling provisioning for UAV‐enabled MEC: A GWFRFT‐based energy‐spreading transmission approach. IET Commun. 14, 15 (Sept. 2020), 2524–2531.
- Li et al. (2024a) Chunlin Li, Kun Jiang, Yong Zhang, Lincheng Jiang, Youlong Luo, and Shaohua Wan. 2024a. Deep Reinforcement Learning-based mining task offloading scheme for intelligent connected vehicles in UAV-aided MEC. ACM Trans. Des. Automat. Electron. Syst. 29, 3 (May 2024), 1–29.
- Li et al. (2020a) Jiangfan Li, Xiaofeng Cao, Deke Guo, Junjie Xie, and Honghui Chen. 2020a. Task scheduling with UAV-assisted vehicular cloud for road detection in highway scenario. IEEE Internet Things J. 7, 8 (Aug. 2020), 7702–7713.
- Li and Nie (2023) Jun Li and Yaojia Nie. 2023. Design of flight test mission evaluation system based on microservice. In Proceedings of the 2023 7th International Conference on Electronic Information Technology and Computer Engineering (Xiamen China). ACM, New York, NY, USA.
- Li et al. (2020c) Meng Li, F Richard Yu, Pengbo Si, Ruizhe Yang, Zhuwei Wang, and Yanhua Zhang. 2020c. UAV-assisted data transmission in blockchain-enabled M2M communications with mobile edge computing. IEEE Netw. 34, 6 (Nov. 2020), 242–249.
- Li et al. (2024c) Qinlin Li, Dan Xing, M A Ilyas, and Nazhatul Hafizah Kamarudin. 2024c. Autonomous balance control method of unmanned aerial vehicle with manipulator based on artificial intelligence algorithm. Soft Comput. 28, 23-24 (Dec. 2024), 13897–13907.
- Li et al. (2018) Xianfeng Li, Jie Chen, and Jianfeng Li. 2018. FATES: A framework with adaptive track-explore strategy for moving targets search by a FANET. In 2018 IEEE Intl Conf on Parallel & Distributed Processing with Applications, Ubiquitous Computing & Communications, Big Data & Cloud Computing, Social Computing & Networking, Sustainable Computing & Communications (ISPA/IUCC/BDCloud/SocialCom/SustainCom) (Melbourne, Australia). IEEE.
- Li et al. (2023a) Xuehua Li, Yu Fang, Chunyu Pan, Yuanxin Cai, and Mingyu Zhou. 2023a. Resource scheduling for UAV-assisted failure-prone MEC in Industrial Internet. Drones 7, 4 (April 2023), 259.
- Li et al. (2024b) Xiang Li, Peng Xiao, Di Tang, Xingguo Li, Qixu Wang, and Dajiang Chen. 2024b. UAVs-assisted QoS guarantee scheme of IoT applications for reliable mobile edge computing. Comput. Commun. 223 (July 2024), 55–67.
- Li et al. (2019) Yanchu Li, Shufang Li, Qingqing Ding, and Chao Wang. 2019. Research on reliability of ad hoc network communication in UAVs formations. In 2019 IEEE 6th International Symposium on Electromagnetic Compatibility (ISEMC) (Nanjing, China). IEEE.
- Li et al. (2023b) Yuke Li, Xianguo Yu, Xiangke Wang, and Jiayi Zheng. 2023b. Design of UAV swarm simulation platform based on ROS1 and ROS2 hybrid architecture. In 2023 42nd Chinese Control Conference (CCC) (Tianjin, China). IEEE.
- Li et al. (2024d) Zewu Li, Chen Xu, Zhanpeng Zhang, and Runze Wu. 2024d. Deep reinforcement learning based trajectory design and resource allocation for task-aware multi-UAV enabled MEC networks. Comput. Commun. 213 (Jan. 2024), 88–98.
- Liang et al. (2021) Shuang Liang, Zhiyi Fang, Geng Sun, Chi Lin, Jiahui Li, Songyang Li, and Aimin Wang. 2021. Charging UAV deployment for improving charging performance of wireless rechargeable sensor networks via joint optimization approach. Comput. Netw. 201, 108573 (Dec. 2021), 108573.
- Lin et al. (2022) Jie Lin, Lin Huang, Hanlin Zhang, Xinyu Yang, and Peng Zhao. 2022. A novel lyapunov based dynamic resource allocation for UAVs-assisted edge computing. Comput. Netw. 205, 108710 (March 2022), 108710.
- Lin et al. (2021) Xinhua Lin, Jing Zhang, Lin Xiang, and Xiaohu Ge. 2021. Energy consumption optimization for UAV assisted private blockchain-based IIoT networks. In 2021 IEEE 94th Vehicular Technology Conference (VTC2021-Fall) (Norman, OK, USA). IEEE.
- Lindqvist et al. (2020) Björn Lindqvist, Sina Sharif Mansouri, Pantelis Sopasakis, and George Nikolakopoulos. 2020. Collision avoidance for multiple micro aerial vehicles using fast centralized nonlinear model predictive control. IFAC-PapersOnLine 53, 2 (2020), 9303–9309.
- Liu et al. (2024a) Jianhua Liu, Peng Xie, Jiajia Liu, and Xiaoguang Tu. 2024a. Task Offloading and Trajectory Optimization in UAV Networks: A Deep Reinforcement Learning Method Based on SAC and A-Star. CMES - Computer Modeling in Engineering and Sciences 141, 2 (2024), 1243–1273. doi:10.32604/cmes.2024.054002
- Liu et al. (2023b) Jianxin Liu, Zhiguo Xu, and Zhigang Wen. 2023b. Joint data transmission and trajectory optimization in UAV-enabled wireless powered mobile edge learning systems. IEEE Trans. Veh. Technol. 72, 9 (Sept. 2023), 11617–11630.
- Liu et al. (2019) Kaikai Liu, Shivam Chauhan, Revathy Devaraj, Sneha Shahi, and Unnikrishnan Sreekumar. 2019. Enabling autonomous unmanned aerial systems via edge computing. In 2019 IEEE International Conference on Service-Oriented System Engineering (SOSE) (San Francisco East Bay, CA, USA). IEEE.
- Liu et al. (2023a) Lei Liu, Jie Wang, and Bingxin Wang. 2023a. Analysis of IMT in the 1765-1785 MHz band for the coexistence of civil UAVs with broadband TDD private network systems. In 2023 International Wireless Communications and Mobile Computing (IWCMC) (Marrakesh, Morocco). IEEE.
- Liu et al. (2024c) Shuai Liu, Helin Yang, Liang Xiao, Mengting Zheng, Huabing Lu, and Zehui Xiong. 2024c. Learning-based resource management optimization for UAV-assisted MEC against jamming. IEEE Trans. Commun. 72, 8 (Aug. 2024), 4873–4886.
- Liu et al. (2024b) Yinan Liu, Chao Yang, Xin Chen, and Fengyan Wu. 2024b. Joint hybrid caching and replacement scheme for UAV-assisted vehicular edge computing networks. IEEE Trans. Intell. Veh. 9, 1 (Jan. 2024), 866–878.
- Liu et al. (2023c) Zhihui Liu, Qiwei Zhang, and Yi Su. 2023c. PPO-based joint optimization for UAV-assisted edge computing networks. Appl. Sci. (Basel) 13, 23 (Nov. 2023), 12828.
- Lorincz et al. (2021) Josip Lorincz, Adnan Tahirovic, and Biljana Risteska Stojkoska. 2021. A novel real-time unmanned aerial vehicles-based disaster management framework. In 2021 29th Telecommunications Forum (TELFOR) (Belgrade, Serbia). IEEE.
- Lu et al. (2024) Jiaheng Lu, Zhenzhe Qu, Anfeng Liu, Shaobo Zhang, and Neal N Xiong. 2024. MLM-WR: A swarm-intelligence-based cloud–edge–terminal collaboration data collection scheme in the era of AIoT. IEEE Internet Things J. 11, 1 (Jan. 2024), 243–255.
- Lu et al. (2022) Yangguang Lu, Xin Chen, Fengjun Zhao, and Ying Chen. 2022. Energy efficient deployment and task offloading for UAV-assisted mobile edge computing. In Lecture Notes in Computer Science. Springer International Publishing, Cham, 421–435.
- Lu et al. (2023) Yongkang Lu, Wei Liu, Yang Zhang, Junqing Li, Shouquan Sun, Jiacheng Cui, and Ting Zeng. 2023. Accurate construction method of sensor measurement network system for in-situ manufacturing of large aerospace component. IEEE Trans. Instrum. Meas. 72 (2023), 1–11.
- Luo et al. (2023) Quyuan Luo, Tom H Luan, Weisong Shi, and Pingzhi Fan. 2023. Deep reinforcement learning based computation offloading and trajectory planning for multi-UAV cooperative target search. IEEE J. Sel. Areas Commun. 41, 2 (Feb. 2023), 504–520.
- Luo et al. (2020) Shixun Luo, Zhongshan Zhang, Shuai Wang, Shuo Zhang, Jibo Dai, Xiangyuan Bu, and Jianping An. 2020. Network for hypersonic UCAV swarms. Sci. China Inf. Sci. 63, 4 (April 2020).
- Luo et al. (2021) Yizhe Luo, Wenrui Ding, and Baochang Zhang. 2021. Optimization of task scheduling and dynamic service strategy for multi-UAV-enabled mobile-edge computing system. IEEE Trans. Cogn. Commun. Netw. 7, 3 (Sept. 2021), 970–984.
- Luo et al. (2022) Yixing Luo, Yuan Zhou, Haiyan Zhao, Zhi Jin, Tianwei Zhang, Yang Liu, Danny Barthaud, and Yijun Yu. 2022. Online adaptation for autonomous unmanned systems driven by requirements satisfaction model. Softw. Syst. Model. 21, 4 (Aug. 2022), 1295–1319.
- Lv et al. (2024) Tianyi Lv, Qingyuan Xia, and Qiwen Zheng. 2024. Intelligent UAV swarm planning based on undirected graph model. In Lecture Notes in Computer Science. Springer Nature Singapore, Singapore, 270–283.
- Ma et al. (2024) Chengyan Ma, Ning Xi, Di Lu, Yebo Feng, and Jianfeng Ma. 2024. CToMP: a cycle-task-oriented memory protection scheme for unmanned systems. Sci. China Inf. Sci. 67, 6 (June 2024).
- Ma and Sun (2022) Zhi Ma and Songlin Sun. 2022. Research on vehicle-to-road collaboration and end-to-end collaboration for multimedia services in the internet of vehicles. IEEE Access 10 (2022), 18146–18155.
- Maatouk et al. (2024) Nesrine Maatouk, Asma Ben Letaifa, and Abderrezak Rachedi. 2024. AI empowered data offloading in MEC enabled UAV networks. Multimedia Tools and Applications (July 2024). doi:10.1007/s11042-024-19859-5
- Machida and Andrade (2021) Fumio Machida and Ermeson Andrade. 2021. PA-offload: Performability-aware adaptive fog offloading for drone image processing. In 2021 IEEE 5th International Conference on Fog and Edge Computing (ICFEC) (Melbourne, Australia). IEEE.
- Machida et al. (2023) Fumio Machida, Qingyang Zhang, and Ermeson Andrade. 2023. Performability analysis of adaptive drone computation offloading with fog computing. Future Gener. Comput. Syst. 145 (Aug. 2023), 121–135.
- Makam et al. ([n. d.]) S Makam, BK Komatineni, SS Meena, and U Meena. [n. d.]. Unmanned aerial vehicles (UAVs): an adoptable technology for precise and smart farming. Discov Internet Things. 2024; 4: 12.
- Manzini et al. (2023) Thomas Manzini, Robin Murphy, David Merrick, and Justin Adams. 2023. Wireless network demands of data products from small uncrewed aerial systems at hurricane Ian. In 2023 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS) (Detroit, MI, USA). IEEE, 9941–9946.
- Manzoor et al. (2021) Aunas Manzoor, Kitae Kim, Shashi Raj Pandey, S M Ahsan Kazmi, Nguyen H Tran, Walid Saad, and Choong Seon Hong. 2021. Ruin theory for energy-efficient resource allocation in UAV-assisted cellular networks. IEEE Trans. Commun. 69, 6 (June 2021), 3943–3956.
- Mao et al. (2023) Weihao Mao, Ke Xiong, Yang Lu, Pingyi Fan, and Zhiguo Ding. 2023. Energy consumption minimization in secure multi-antenna UAV-assisted MEC networks with channel uncertainty. IEEE Trans. Wirel. Commun. 22, 11 (Nov. 2023), 7185–7200.
- Marcos and Utete (2021) Juliana T C Marcos and Simukai W Utete. 2021. Animal tracking within a formation of drones. In 2021 IEEE 24th International Conference on Information Fusion (FUSION) (Sun City, South Africa). IEEE.
- Mart´ınez-Sánchez et al. (2023) J Martínez-Sánchez, D Conde Morales, S Novoa, and P Arias. 2023. An Unmanned Aerial System for on-the-edge traffic monitoring in work-zone approach area. ISPRS - Int. Arch. Photogramm. Remote Sens. Spat. Inf. Sci. XLVIII-1/W2-2023 (Dec. 2023), 605–610.
- Mcilwaine and Rivas Casado (2021) Ben Mcilwaine and Mónica Rivas Casado. 2021. JellyNet: The convolutional neural network jellyfish bloom detector. Int. J. Appl. Earth Obs. Geoinf. 97, 102279 (May 2021), 102279.
- Mekdad et al. (2023) Yassine Mekdad, Ahmet Aris, Leonardo Babun, Abdeslam El Fergougui, Mauro Conti, Riccardo Lazzeretti, and A Selcuk Uluagac. 2023. A survey on security and privacy issues of UAVs. Computer networks 224 (2023), 109626.
- Miclea and Sanislav (2011) Liviu Miclea and Teodora Sanislav. 2011. About dependability in cyber-physical systems. In 2011 9th East-West Design & Test Symposium (EWDTS). IEEE, 17–21.
- Mo et al. (2022) Wen Mo, Wei Liu, Guosheng Huang, Naixue Xiong, Anfeng Liu, and Shaobo Zhang. 2022. A cloud-assisted reliable trust computing scheme for data collection in internet of things. IEEE Trans. Industr. Inform. 18, 7 (July 2022), 4969–4980.
- Moeyersons et al. (2021) Jerico Moeyersons, Martijn Gevaert, Karl-Erik Réculé, Bruno Volckaert, and Filip De Turck. 2021. UAVs-as-a-Service: Cloud-based Remote Application Management for Drones. In 2021 IFIP/IEEE International Symposium on Integrated Network Management (IM). 926–931.
- Moghimi et al. (2020) Ali Moghimi, Ce Yang, and James A Anderson. 2020. Aerial hyperspectral imagery and deep neural networks for high-throughput yield phenotyping in wheat. Computers and Electronics in Agriculture 172 (2020), 105299.
- Mukherjee et al. (2020) Amartya Mukherjee, Nilanjan Dey, and Debashis De. 2020. EdgeDrone: QoS aware MQTT middleware for mobile edge computing in opportunistic Internet of Drone Things. Comput. Commun. 152 (Feb. 2020), 93–108.
- Mukherjee et al. (2023) Amartya Mukherjee, Nilanjan Dey, Atreyee Mondal, Debashis De, and Rubén González Crespo. 2023. iSocialDrone: QoS aware MQTT middleware for social internet of drone things in 6G-SDN slice. Soft Comput. 27, 8 (April 2023), 5119–5135.
- Naouri et al. (2024) Abdenacer Naouri, Huansheng Ning, Nabil Abdelkader Nouri, Amar Khelloufi, Abdelkarim Ben Sada, Salim Naouri, Attia Qammar, and Sahraoui Dhelim. 2024. Maximizing UAV fog deployment efficiency for critical rescue operations: A multi-objective optimization approach. Future Gener. Comput. Syst. 159 (Oct. 2024), 255–271.
- Narzari et al. (2025) Rumi Narzari, Burhan U Choudhury, Gaurav Singhal, and Karun K Choudhary. 2025. A critical review of how UAVs can transform precision agriculture in the realm of Agroecology. Discover Soil 2, 1 (2025), 28.
- Navardi and Mohsenin (2023) Mozhgan Navardi and Tinoosh Mohsenin. 2023. MLAE2: Metareasoning for latency-aware energy-efficient autonomous nano-drones. In 2023 IEEE International Symposium on Circuits and Systems (ISCAS) (Monterey, CA, USA). IEEE, 1–5.
- Ngoenriang et al. (2022) Napat Ngoenriang, Stephen John Turner, Dusit Niyato, and Sucha Supittayapornpong. 2022. Joint UAV placement and data delivery in aerial inspection under uncertainties. IEEE Internet Things J. 9, 9 (May 2022), 6389–6403.
- Nguyen et al. (2023) Tri Nguyen, Risto Katila, and Tuan Nguyen Gia. 2023. An advanced Internet-of-Drones System with Blockchain for improving quality of service of Search and Rescue: A feasibility study. Future Gener. Comput. Syst. 140 (March 2023), 36–52.
- Ning et al. (2023) Zhaolong Ning, Hao Hu, Xiaojie Wang, Lei Guo, Song Guo, Guoyin Wang, and Xinbo Gao. 2023. Mobile edge computing and machine learning in the internet of unmanned aerial vehicles: a survey. Comput. Surveys 56, 1 (2023), 1–31.
- Ning et al. (2024) Zhaolong Ning, Yuxuan Yang, Xiaojie Wang, Qingyang Song, Lei Guo, and Abbas Jamalipour. 2024. Multi-agent deep reinforcement learning based UAV trajectory optimization for differentiated services. IEEE Trans. Mob. Comput. (2024), 1–17.
- Nwadiugwu et al. (2021) Williams-Paul Nwadiugwu, Seung Hwan Kim, Jae-Min Lee, and Dong-Seong Kim. 2021. Prioritized-MAC model for intelligent UAV-to-BS communication in industrial-WSN systems. In 2021 International Conference on Artificial Intelligence in Information and Communication (ICAIIC) (Jeju Island, Korea (South)). IEEE, 339–343.
- Ouyang et al. (2024) Haibin Ouyang, Leisen Liang, Steven Li, and Liqun Gao. 2024. A joint optimization method for multi-UAV deployment and task scheduling in mobile edge computing with large-scale mobile users. Expert Syst. Appl. 255, 124526 (Dec. 2024), 124526.
- Ouyang et al. (2023) Haibin Ouyang, Kang Liu, Chunliang Zhang, Steven Li, and Liqun Gao. 2023. Large-scale mobile users deployment optimization based on a two-stage hybrid global HS-DE algorithm in multi-UAV-enabled mobile edge computing. Eng. Appl. Artif. Intell. 124, 106608 (Sept. 2023), 106608.
- Ozger et al. (2018) Mustafa Ozger, Michal Vondra, and Cicek Cavdar. 2018. Towards beyond visual line of sight piloting of UAVs with ultra reliable low latency communication. In 2018 IEEE Global Communications Conference (GLOBECOM) (Abu Dhabi, United Arab Emirates). IEEE.
- Pacheco et al. (2021) Lucas Pacheco, Helder Oliveira, Denis Rosario, Zhongliang Zhao, Eduardo Cerqueira, Torsten Braun, and Paulo Mendes. 2021. Towards the future of edge computing in the sky: Outlook and future directions. In 2021 17th International Conference on Distributed Computing in Sensor Systems (DCOSS) (Pafos, Cyprus). IEEE.
- Pak et al. (2024) Henry Pak, Lukas Asmer, Petra Kokus, Bianca I Schuchardt, Albert End, Frank Meller, Karolin Schweiger, Christoph Torens, Carolina Barzantny, Dennis Becker, et al. 2024. Can Urban Air Mobility become reality? Opportunities and challenges of UAM as innovative mode of transport and DLR contribution to ongoing research. CEAS Aeronautical Journal (2024), 1–31.
- Palossi et al. (2019) Daniele Palossi, Andres Gomez, Stefan Draskovic, Andrea Marongiu, Lothar Thiele, and Luca Benini. 2019. Extending the lifetime of nano-blimps via dynamic motor control. J. Signal Process. Syst. 91, 3-4 (March 2019), 339–361.
- Pamuklu et al. (2023) Turgay Pamuklu, Aisha Syed, W Sean Kennedy, and Melike Erol-Kantarci. 2023. Heterogeneous GNN-RL-based task offloading for UAV-aided smart agriculture. IEEE Netw. Lett. 5, 4 (Dec. 2023), 213–217.
- Pan et al. (2024) Lehao Pan, Tong Liu, Jianghua Cheng, Bang Cheng, and Yahui Cai. 2024. AIMED-net: An enhancing infrared small target detection net in UAVs with multi-layer feature enhancement for edge computing. Remote Sens. (Basel) 16, 10 (May 2024), 1776.
- Pan et al. (2022) Wenbin Pan, Yi Liu, and Chao Yang. 2022. Adaptive task offloading of rechargeable UAV edge computing network based on double decision value iteration. Comput. Commun. 193 (Sept. 2022), 136–145.
- Pandey et al. (2021b) Shashi Raj Pandey, Kitae Kim, Madyan Alsenwi, Yan Kyaw Tun, Zhu Han, and Choong Seon Hong. 2021b. Latency-sensitive service delivery with UAV-assisted 5G networks. IEEE Wirel. Commun. Lett. 10, 7 (July 2021), 1518–1522.
- Pandey et al. (2021a) Shashi Raj Pandey, Kitae Kim, Madyan Alsenwi, Yan Kyaw Tun, and Choong Seon Hong. 2021a. A crowd-enabled task execution approach in UAV networks towards fog computing. In 2021 IEEE International Conference on Big Data and Smart Computing (BigComp) (Jeju Island, Korea (South)). IEEE.
- Pang et al. (2024) Shengye Pang, Yi Li, Zhen Qin, Xinkui Zhao, Jintao Chen, Fan Wang, and Jianwei Yin. 2024. Adaptive scheduling of high-availability drone swarms for congestion alleviation in connected automated vehicles. ACM Trans. Auton. Adapt. Syst. (June 2024).
- Panwar et al. (2023) Poonam Panwar, Mohammad Shabaz, Shah Nazir, Ismail Keshta, Ali Rizwan, and R Sugumar. 2023. Generic edge computing system for optimization and computation offloading of unmanned aerial vehicle. Comput. Electr. Eng. 109, 108779 (July 2023), 108779.
- Paranjothi et al. (2019) Anirudh Paranjothi, Urcun Tanik, Yuehua Wang, and Mohammad S Khan. 2019. Hybrid‐Vehfog: A robust approach for reliable dissemination of critical messages in connected vehicles. Trans. Emerg. Telecommun. Technol. 30, 6 (June 2019), e3595.
- Patnayak et al. (2021) Chinmaya Patnayak, James E McClure, and Ryan K Williams. 2021. WASP: A wearable super-computing platform for distributed intelligence in multi-agent systems. In 2021 IEEE High Performance Extreme Computing Conference (HPEC) (Waltham, MA, USA). IEEE.
- Peng et al. (2022) Chaoda Peng, Xumin Huang, Yuan Wu, and Jiawen Kang. 2022. Constrained multi-objective optimization for UAV-enabled mobile edge computing: Offloading optimization and path planning. IEEE Wirel. Commun. Lett. 11, 4 (April 2022), 861–865.
- Peng et al. (2023) Haiying Peng, Yufeng Zheng, Peng He, Yaping Cui, Ruyan Wang, and Dapeng Wu. 2023. Reliability optimization in IRS-assisted UAV networks. Veh. Commun. 44, 100679 (Dec. 2023), 100679.
- Peng et al. (2024) Yuhuai Peng, Xiaoliang Guang, Xinyu Zhang, Lei Liu, Cemulige Wu, and Lei Huang. 2024. A cloud-edge collaborative computing framework using potential games for space-air-ground integrated IoT. EURASIP J. Adv. Signal Process. 2024, 1 (April 2024).
- Perno et al. (2023) Matteo Perno, Lars Hvam, and Anders Haug. 2023. A machine learning digital twin approach for critical process parameter prediction in a catalyst manufacturing line. Computers in Industry 151 (2023), 103987.
- Petersen et al. (2008) Kai Petersen, Robert Feldt, Shahid Mujtaba, and Michael Mattsson. 2008. Systematic mapping studies in software engineering. In 12th international conference on evaluation and assessment in software engineering (EASE). BCS Learning & Development.
- Petrenko et al. (2023) Vyacheslav Petrenko, Fariza Tebueva, Vladimir Antonov, Sergey Ryabtsev, Andrey Pavlov, and Artur Sakolchik. 2023. Method and algorithm for task allocation in a heterogeneous group of UAVs in a clustered field of targets. J. King Saud Univ. - Comput. Inf. Sci. 35, 6 (June 2023), 101580.
- Picano and Fantacci (2024) Benedetta Picano and Romano Fantacci. 2024. A combined stochastic network calculus and matching theory approach for computational offloading in a heterogenous MEC environment. IEEE Trans. Netw. Serv. Manag. 21, 2 (April 2024), 1958–1968.
- Pradhan et al. (2015) Subhav M Pradhan, Abhishek Dubey, Aniruddha Gokhale, and Martin Lehofer. 2015. CHARIOT: a domain specific language for extensible cyber-physical systems. In Proceedings of the Workshop on Domain-Specific Modeling (Pittsburgh PA USA). ACM, New York, NY, USA.
- Pustokhina et al. (2021) Irina V Pustokhina, Denis A Pustokhin, E Laxmi Lydia, Mohamed Elhoseny, and K Shankar. 2021. Energy efficient neuro-fuzzy cluster based topology construction with metaheuristic route planning algorithm for unmanned aerial vehicles. Comput. Netw. 196, 108214 (Sept. 2021), 108214.
- Qadir et al. (2023) Zakria Qadir, Khoa N. Le, Nasir Saeed, and Hafiz Suliman Munawar. 2023. Towards 6G Internet of Things: Recent advances, use cases, and open challenges. ICT Express 9, 3 (2023), 296–312. doi:10.1016/j.icte.2022.06.006
- Qayyum et al. (2021) Tariq Qayyum, Zouheir Trabelsi, Asad Malik, and Kadhim Hayawi. 2021. Trajectory design for UAV-based data collection using clustering model in smart farming. Sensors (Basel) 22, 1 (Dec. 2021), 37.
- Qu et al. (2022) Chengyi Qu, Rounak Singh, Alicia Esquivel-Morel, and Prasad Calyam. 2022. Learning-based multi-drone network edge orchestration for video analytics. In IEEE INFOCOM 2022 - IEEE Conference on Computer Communications (London, United Kingdom). IEEE.
- Qureshi et al. (2024) Khalid Ibrahim Qureshi, Lei Wang, Xuanrui Xiong, and Muhammad Ali Lodhi. 2024. Asynchronous federated learning for resource allocation in software-defined internet of UAVs. IEEE Internet Things J. 11, 12 (June 2024), 20899–20911.
- Rahul et al. (2021) Ashiqur Rahman Rahul, Saifur Rahman Sabuj, Md Sajid Akbar, Han-Shin Jo, and Md Akbar Hossain. 2021. An optimization based approach to enhance the throughput and energy efficiency for cognitive unmanned aerial vehicle networks. Wirel. Netw. 27, 1 (Jan. 2021), 475–493.
- Raj and Simmhan (2023) Suman Raj and Yogesh Simmhan. 2023. Towards a mobile app platform for personalized UAV fleets using edge and cloud. In 2023 IEEE/ACM 23rd International Symposium on Cluster, Cloud and Internet Computing Workshops (CCGridW) (Bangalore, India). IEEE, 331–333.
- Ramasamy et al. (2024) Jayabharathi Ramasamy, E Srividhya, V Vaidehi, S Vimaladevi, N Mohankumar, and S Murugan. 2024. Cloud-enabled isolation forest for anomaly detection in UAV-based power line inspection. In 2024 2nd International Conference on Networking and Communications (ICNWC) (Chennai, India). IEEE.
- Ramos-Ramos et al. (2024) Diego Ramos-Ramos, Alejandro González-Vegas, Javier Berrocal, and Jaime Galán-Jiménez. 2024. Energy-aware microservice-based application deployment in UAV-based networks for rural scenarios. J. Netw. Syst. Manag. 32, 3 (July 2024).
- Rao and Naidu (2024) Kota Nageswara Rao and Kalpana Naidu. 2024. Optimizing energy efficiency in mobile edge computing: Leveraging latency-aware offloading, clustering, and UAV placement strategies. Int. J. Electron. Commun. 185, 155447 (Oct. 2024), 155447.
- Ren et al. (2024) Chao Ren, Chao Gong, and Luchuan Liu. 2024. Task-oriented multimodal communication based on cloud–edge–UAV collaboration. IEEE Internet Things J. 11, 1 (Jan. 2024), 125–136.
- Ren et al. (2017) Rong Ren, Lin Zhu, Qi Xi, and Zhiqiang Wang. 2017. A design and research of digital operation support framework for civil aircrafts. In 2017 International Conference on Sensing, Diagnostics, Prognostics, and Control (SDPC) (Shanghai). IEEE.
- Ritchie et al. (2020) James Ritchie, Andrew Tasso, Adrian Rusu, Amalia Rusu, and Davis Doherty. 2020. Upgrading JavaCat: A distributed system for time-shifted air traffic scenario generation. In 2020 IEEE International Systems Conference (SysCon) (Montreal, QC, Canada). IEEE.
- Rokaha et al. (2019) Bhagawan Rokaha, Bishnu Prasad Gautam, and Tomoya Kitani. 2019. Building a reliable and cost-effective RTK-GNSS infrastructure for precise positioning of IoT applications. In 2019 Twelfth International Conference on Mobile Computing and Ubiquitous Network (ICMU) (Kathmandu, Nepal). IEEE.
- Russell et al. (2018) Luke Russell, Rafik Goubran, and Felix Kwamena. 2018. Computer vision-assisted human-in-the-loop measurements: augmenting qualitative by increasing quantitative analytics for CI situational awareness. In 2018 IEEE International Conference on Computational Intelligence and Virtual Environments for Measurement Systems and Applications (CIVEMSA) (Ottawa, ON, Canada). IEEE.
- Sa et al. (2017) Inkyu Sa, Zetao Chen, Marija Popović, Raghav Khanna, Frank Liebisch, Juan Nieto, and Roland Siegwart. 2017. weednet: Dense semantic weed classification using multispectral images and mav for smart farming. IEEE robotics and automation letters 3, 1 (2017), 588–595.
- Saadaoui et al. (2023) Fatima Zohra Saadaoui, Nawal Cheggaga, and Nour El Houda Djabri. 2023. Multi-sensory system for UAVs detection using Bayesian inference. Appl. Intell. (Nov. 2023).
- Sabino et al. (2024) Arthur Sabino, Luiz Nelson Lima, Carlos Brito, Leonel Feitosa, Marcos F Caetano, Priscila Solis Barreto, and Francisco Airton Silva. 2024. Forest fire monitoring system supported by unmanned aerial vehicles and edge computing: a performance evaluation using petri nets. Cluster Comput. 27, 7 (Oct. 2024), 9735–9755.
- Salehi et al. (2022) Fateme Salehi, Mustafa Ozger, Naaser Neda, and Cicek Cavdar. 2022. Ultra-reliable low-latency communication for aerial vehicles via multi-connectivity. In 2022 Joint European Conference on Networks and Communications & 6G Summit (EuCNC/6G Summit). IEEE, 166–171.
- Sankarasrinivasan et al. (2015) S Sankarasrinivasan, E Balasubramanian, K Karthik, U Chandrasekar, and Rishi Gupta. 2015. Health monitoring of civil structures with integrated UAV and image processing system. Procedia Comput. Sci. 54 (2015), 508–515.
- Sara et al. (2016) Mohamed Sara, Imad Jawhar, and Mohamed Nader. 2016. A softwarization architecture for UAVs and WSNs as part of the cloud environment. In 2016 IEEE International Conference on Cloud Engineering Workshop (IC2EW) (Berlin, Germany). IEEE.
- Schäfer et al. (2018) Lukas Schäfer, Sergio García, and Vassili Srithammavanh. 2018. Simplification of inclusion–exclusion on intersections of unions with application to network systems reliability. Reliab. Eng. Syst. Saf. 173 (May 2018), 23–33.
- Seisa et al. (2022) Achilleas Santi Seisa, Sumeet Gajanan Satpute, Bjorn Lindqvist, and George Nikolakopoulos. 2022. An edge architecture oriented model predictive control scheme for an autonomous UAV mission. In 2022 IEEE 31st International Symposium on Industrial Electronics (ISIE) (Anchorage, AK, USA). IEEE.
- Selim et al. (2021) Mahmoud M Selim, Mohamed Rihan, Yatao Yang, and Junxian Ma. 2021. Optimal task partitioning, Bit allocation and trajectory for D2D-assisted UAV-MEC systems. Peer Peer Netw. Appl. 14, 1 (Jan. 2021), 215–224.
- Seong et al. (2024) Mincheol Seong, Ohyun Jo, and Kyungseop Shin. 2024. Age of information minimization in UAV-assisted data harvesting networks by multi-agent deep reinforcement curriculum learning. Expert Syst. Appl. 255, 124379 (Dec. 2024), 124379.
- Shams et al. (2023) Farshad Shams, Vincenzo Lottici, and Filippo Giannetti. 2023. Joint spectrum and computing resource allocation in fog-assisted drone communications for ambiental services. IEEE Access 11 (2023), 115348–115366.
- Shang et al. (2023) Yuanfeng Shang, Chang Liu, Dawei Qiu, Zixu Zhao, Ruikang Wu, and Shuyuan Tang. 2023. AD-YOLOv5s based UAV detection for low altitude security. Int. J. Micro Air Veh. 15 (Jan. 2023).
- Shao et al. (2023) Ziling Shao, Helin Yang, Liang Xiao, Wei Su, and Zehui Xiong. 2023. Energy and latency-aware resource management for UAV-assisted mobile edge computing against jamming. In GLOBECOM 2023 - 2023 IEEE Global Communications Conference (Kuala Lumpur, Malaysia). IEEE.
- Sharma et al. (2021) Susanta Sharma, Allumallu Veera Venkata Susmitha, Lan-Da Van, and Yu-Chee Tseng. 2021. An edge-controlled outdoor autonomous UAV for colorwise safety helmet detection and counting of workers in construction sites. In 2021 IEEE 94th Vehicular Technology Conference (VTC2021-Fall) (Norman, OK, USA). IEEE.
- Sharma et al. (2019) Vishal Sharma, Ilsun You, Dushantha Nalin K Jayakody, Daniel Gutierrez Reina, and Kim-Kwang Raymond Choo. 2019. Neural-blockchain-based ultrareliable caching for edge-enabled UAV networks. IEEE Trans. Industr. Inform. 15, 10 (Oct. 2019), 5723–5736.
- She et al. (2018) Rui She, Shanyun Liu, and Pingyi Fan. 2018. State variation mining: On information divergence with message importance in big data. In 2018 IEEE Global Communications Conference (GLOBECOM) (Abu Dhabi, United Arab Emirates). IEEE.
- She and Ouyang (2021) Ruifeng She and Yanfeng Ouyang. 2021. Efficiency of UAV-based last-mile delivery under congestion in low-altitude air. Transp. Res. Part C Emerg. Technol. 122, 102878 (Jan. 2021), 102878.
- Shen et al. (2023a) Bo Shen, Qi Gu, and Gang Yang. 2023a. Joint task offloading and UAVs deployment for UAV-assisted mobile edge computing. Comput. Netw. 234, 109943 (Oct. 2023), 109943.
- Shen et al. (2023b) Shuai Shen, Kun Yang, Kezhi Wang, and Guopeng Zhang. 2023b. UAV-aided vehicular short-packet communication and edge computing system under time-varying channel. IEEE Trans. Veh. Technol. 72, 5 (May 2023), 6625–6638.
- Shi et al. (2023) Junling Shi, Chunyu Li, Yunchong Guan, Peiyu Cong, and Jie Li. 2023. Multi-UAV-assisted computation offloading in DT-based networks: A distributed deep reinforcement learning approach. Comput. Commun. 210 (Oct. 2023), 217–228.
- Shrestha et al. (2021) Rakesh Shrestha, Rojeena Bajracharya, and Shiho Kim. 2021. 6G Enabled Unmanned Aerial Vehicle Traffic Management: A Perspective. IEEE Access PP (06 2021), 1–1. doi:10.1109/ACCESS.2021.3092039
- Silva et al. (2024) Francisco Airton Silva, Iure Fe, Carlos Brito, Gabriel Araujo, Leonel Feitosa, Tuan Anh Nguyen, Kwonsu Jeon, Jae-Woo Lee, Dugki Min, and Eunmi Choi. 2024. Aerial computing: Enhancing mobile cloud computing with unmanned aerial vehicles as data bridges—A Markov chain based dependability quantification. ICT Express 10, 2 (April 2024), 406–411.
- Singh et al. (2023) Manish Kumar Singh, Amit Choudhary, Sandeep Gulia, and Anurag Verma. 2023. Multi-objective NSGA-II optimization framework for UAV path planning in an UAV-assisted WSN. J. Supercomput. 79, 1 (Jan. 2023), 832–866.
- Song et al. (2022) Bingbing Song, Haiyang Chen, Jiashun Suo, and Wei Zhou. 2022. Low-power robustness learning framework for adversarial attack on edges. In 2022 18th International Conference on Mobility, Sensing and Networking (MSN) (Guangzhou, China). IEEE.
- Souli et al. (2023) Nicolas Souli, Panayiotis Kolios, and Georgios Ellinas. 2023. Online distributed relative positioning utilizing multiple cooperative autonomous agents. J. Intell. Robot. Syst. 109, 4 (Dec. 2023).
- Su et al. (2022) Weibin Su, Allemar Jhone P Delima, Ivy Kim D Machica, Jan Carlo T Arroyo, and Xiangbin Pan. 2022. An AED-UAV healthcare information robot integration system based on edge computing airport station. In 2022 the 5th International Conference on Robot Systems and Applications (ICRSA) (Shenzhen China). ACM, New York, NY, USA.
- Su (2021) Yu Su. 2021. A trust based scheme to protect 5G UAV communication networks. IEEE Open J. Comput. Soc. 2 (2021), 300–307.
- Subburaj et al. (2023) Brindha Subburaj, Uma Maheswari Jayachandran, Vinothini Arumugham, and Miruna Joe Amali Suthanthira Amalraj. 2023. A self-adaptive trajectory optimization algorithm using fuzzy logic for mobile edge computing system assisted by unmanned aerial vehicle. Drones 7, 4 (April 2023), 266.
- Suganya et al. (2024) B Suganya, R Gopi, A Ranjith Kumar, and Gavendra Singh. 2024. Dynamic task offloading edge-aware optimization framework for enhanced UAV operations on edge computing platform. Sci. Rep. 14, 1 (July 2024), 16383.
- Sun et al. (2020) Han Sun, Jian Yang, Jiaquan Shen, Dong Liang, Liu Ning-Zhong, and Huiyu Zhou. 2020. TIB-net: Drone detection network with tiny iterative backbone. IEEE Access 8 (2020), 130697–130707.
- Sun et al. (2018) Yuxuan Sun, Jinhui Song, Sheng Zhou, Xueying Guo, and Zhisheng Niu. 2018. Task replication for vehicular edge computing: A combinatorial multi-armed bandit based approach. In 2018 IEEE Global Communications Conference (GLOBECOM) (Abu Dhabi, United Arab Emirates). IEEE.
- Tahat et al. (2020) Ashraf Tahat, Talal Edwan, Hamza Al-Sawwaf, Jumana Al-Baw, and Mohammad Amayreh. 2020. Simplistic machine learning-based air-to-ground path loss modeling in an urban environment. In 2020 Fifth International Conference on Fog and Mobile Edge Computing (FMEC) (Paris, France). IEEE, 158–163.
- Taleb et al. (2023) Tarik Taleb, Nassim Sehad, Zinelaabidine Nadir, and Jaeseung Song. 2023. VR-Based Immersive Service Management in B5G Mobile Systems: A UAV Command and Control Use Case. IEEE Internet Things J. 10, 6 (March 2023), 5349–5363.
- Tan et al. (2020) Bo Tan, John Cosmas, Konstantinos Banitsas, Michael Rustell, Joonas Säe, Mikko Valkama, Yue Zhang, Edgardo Montes de Oca, and David Aondoakaa. 2020. Using 5G to Bring More than just Bits to Homes. In 2020 IEEE International Symposium on Broadband Multimedia Systems and Broadcasting (BMSB) (Paris, France). IEEE.
- Tang et al. (2024b) Jianhang Tang, Guoquan Wu, Mohammad Mussadiq Jalalzai, Lin Wang, Bing Zhang, and Yi Zhou. 2024b. Energy-optimal DNN model placement in UAV-enabled edge computing networks. Digit. Commun. Netw. 10, 4 (Aug. 2024), 827–836.
- Tang et al. (2023) Qiang Tang, Chen Dai, Zhiqiang Yu, Dun Cao, and Jin Wang. 2023. An UAV and EV based mobile edge computing system for total delay minimization. Comput. Commun. 212 (Dec. 2023), 104–115.
- Tang et al. (2022a) Qiang Tang, Linjiang Li, Caiyan Jin, Lixin Liu, Jin Wang, Zhuofan Liao, and Yuansheng Luo. 2022a. Heterogeneous UAVs assisted mobile edge computing for energy consumption minimization of the edge side. Comput. Commun. 194 (Oct. 2022), 268–279.
- Tang et al. (2022b) Qiang Tang, Linjiang Li, Jin Wang, Gwang-Jun Kim, and Bin Tang. 2022b. Energy efficiency maximization for UAV and electric vehicle assisted mobile edge computing on the edge side. In Lecture Notes in Computer Science. Springer International Publishing, Cham, 469–481.
- Tang et al. (2024a) Qiang Tang, Xinyu Qu, Jin Wang, and Shiming He. 2024a. A multi-UAV assisted non-orthogonal multiple access based relay system for minimal average receiving rate maximization. Soft Comput. 28, 11-12 (June 2024), 7125–7137.
- Tang et al. (2024c) Xiao Tang, Hongrui Zhang, Ruonan Zhang, Deyun Zhou, Yan Zhang, and Zhu Han. 2024c. Robust trajectory and offloading for energy-efficient UAV edge computing in industrial internet of things. IEEE Trans. Industr. Inform. 20, 1 (Jan. 2024), 38–49.
- Tareq et al. (2018) Md Mostofa Kamal Tareq, Omid Semiari, Mohsen Amini Salehi, and Walid Saad. 2018. Ultra reliable, low latency vehicle-to-infrastructure wireless communications with edge computing. In 2018 IEEE Global Communications Conference (GLOBECOM) (Abu Dhabi, United Arab Emirates). IEEE.
- Tendolkar et al. (2021) Atharv Tendolkar, Amit Choraria, M M Manohara Pai, S Girisha, Gavin Dsouza, and K S Adithya. 2021. Modified crop health monitoring and pesticide spraying system using NDVI and Semantic Segmentation: An AGROCOPTER based approach. In 2021 IEEE International Conference on Autonomous Systems (ICAS) (Montreal, QC, Canada). IEEE.
- Thapliyal et al. (2024) Shardul Thapliyal, Rajoo Pandey, and Chhagan Charan. 2024. Energy-harvesting-enabled UAV relay-assisted NOMA network under finite blocklength. IEEE Internet Things J. 11, 13 (July 2024), 23255–23267.
- Tian et al. (2024) Jinyuan Tian, Li Zhu, Fei Richard, Hongwei Wang, and Tao Tang. 2024. Optimizing edge resources in intelligent railway construction: A two-level game approach. IEEE Trans. Veh. Technol. 73, 6 (June 2024), 7589–7602.
- Tripathi et al. (2023) Shraddha Tripathi, Om Jee Pandey, Linga Reddy Cenkeramaddi, and Rajesh M Hegde. 2023. A socially-aware radio map framework for improving QoS of UAV-assisted MEC networks. IEEE Trans. Netw. Serv. Manag. 20, 1 (March 2023), 342–356.
- Tsamados et al. (2024) Andreas Tsamados, Luciano Floridi, and Mariarosaria Taddeo. 2024. Human control of AI systems: from supervision to teaming. AI and Ethics (2024), 1–14.
- Tsung et al. (2022) Chen-Kun Tsung, Fa-Shian Chang, and Xiu-Yu Liu. 2022. On the construction of an edge-based remote sensing framework: The applications on automated guided vehicles and drones. Electronics (Basel) 11, 7 (March 2022), 1034.
- Vamvakas et al. (2019) Panagiotis Vamvakas, Eirini Eleni Tsiropoulou, and Symeon Papavassiliou. 2019. Exploiting prospect theory and risk-awareness to protect UAV-assisted network operation. EURASIP J. Wirel. Commun. Netw. 2019, 1 (Dec. 2019).
- Vankka (2013) Jouko Vankka. 2013. Performance of satellite gateway over geostationary satellite links. In MILCOM 2013-2013 IEEE Military Communications Conference. IEEE, 289–292.
- Vélez et al. (2024) Sergio Vélez, Mar Ariza-Sentís, Marko Panić, Bojana Ivošević, Dimitrije Stefanović, Jere Kaivosoja, and João Valente. 2024. Speeding up UAV-based crop variability assessment through a data fusion approach using spatial interpolation for site-specific management. Smart Agricultural Technology 8, 100488 (Aug. 2024), 100488.
- Vemulapalli and Mesapam (2021) Sai Charitha Vemulapalli and Shashi Mesapam. 2021. Slope stability analysis for mine hazard assessment using UAV. J. Ind. Soc. Remote Sens. 49, 7 (July 2021), 1483–1491.
- Venkatagiri et al. (2018) Radha Venkatagiri, Karthik Swaminathan, Chung-Ching Lin, Liang Wang, Alper Buyuktosunoglu, Pradip Bose, and Sarita Adve. 2018. Impact of software approximations on the resiliency of a video summarization system. In 2018 48th Annual IEEE/IFIP International Conference on Dependable Systems and Networks (DSN) (Luxembourg City). IEEE.
- Vyasa and Xu (2024) Vikas Vyasa and Zheyuan Xu. 2024. Maintenance in automotive and aerospace applications–An overview. International Journal of Advances in Science Transactions 3, 2 (2024), 349–361.
- Wang et al. (2022a) Benkuan Wang, Bo Chen, Datong Liu, and Xiyuan Peng. 2022a. An embedded instrument for online condition monitoring of unmanned and autonomous systems. IEEE Instrum. Meas. Mag. 25, 5 (Aug. 2022), 23–29.
- Wang et al. (2020b) G Wang, B Lee, J Ahn, and G Cho. 2020b. A UAV-assisted CH election framework for secure data collection in wireless sensor networks. Future Gener. Comput. Syst. 102 (Jan. 2020), 152–162.
- Wang et al. (2020d) Guangying Wang, Xiangbin Yu, Fangcheng Xu, and Jiali Cai. 2020d. Task offloading and resource allocation for UAV-assisted mobile edge computing with imperfect channel estimation over Rician fading channels. EURASIP J. Wirel. Commun. Netw. 2020, 1 (Dec. 2020).
- Wang et al. (2018) Jingrong Wang, Kaiyang Liu, Minming Ni, and Jianping Pan. 2018. Learning based mobility management under uncertainties for mobile edge computing. In 2018 IEEE Global Communications Conference (GLOBECOM) (Abu Dhabi, United Arab Emirates). IEEE.
- Wang et al. (2023b) Juzhen Wang, Xiaoli Zhang, Xingshi He, and Yongqiang Sun. 2023b. Bandwidth allocation and trajectory control in UAV-assisted IoV edge computing using multiagent reinforcement learning. IEEE Trans. Reliab. 72, 2 (June 2023), 599–608.
- Wang et al. (2019c) Peng Wang, Jiaxin Zhang, Xing Zhang, Zhi Yan, Barry G Evans, and Wenbo Wang. 2019c. Convergence of satellite and terrestrial networks: A comprehensive survey. IEEE access 8 (2019), 5550–5588.
- Wang et al. (2024c) Qixu Wang, Peng Xiao, Xiang Li, Yunxiang Qiu, Tao Zheng, and Zhiguang Qin. 2024c. Taas: Trust assessment as a service for secure communication of green edge-assisted UAV network. Peer Peer Netw. Appl. 17, 4 (July 2024), 2221–2236.
- Wang et al. (2020a) Rui Wang, Yong Cao, Adeeb Noor, Thamer A Alamoudi, and Redhwan Nour. 2020a. Agent-enabled task offloading in UAV-aided mobile edge computing. Comput. Commun. 149 (Jan. 2020), 324–331.
- Wang et al. (2021e) Rui Wang, Xiangping Bryce Zhai, Yunlong Zhao, and Xuedong Zhao. 2021e. Delivery optimization for unmanned aerial vehicles based on minimum cost maximum flow with limited battery capacity. In Lecture Notes in Computer Science. Springer International Publishing, Cham, 77–85.
- Wang et al. (2025) Shengdong Wang, Zhen Jia, Zhenbao Liu, Yong Tang, Xinshang Qin, and Xiao Wang. 2025. Self-supervised contrast learning based UAV fault detection and interpretation with spatial–temporal information of multivariate flight data. Expert Systems with Applications 267 (2025), 126156.
- Wang et al. (2022b) Tianhao Wang, Xiaowei Pang, Jie Tang, Nan Zhao, Xiuyin Zhang, and Xianbin Wang. 2022b. Time and energy efficient data collection via UAV. Sci. China Inf. Sci. 65, 8 (Aug. 2022).
- Wang et al. (2019a) Wenbo Wang, Niloofar Okati, Islam Tanash, Taneli Riihonen, and Elena-Simona Lohan. 2019a. Location-based beamforming architecture for efficient farming applications with drones. In 2019 International Conference on Localization and GNSS (ICL-GNSS) (Nuremberg, Germany). IEEE.
- Wang et al. (2023a) Weizheng Wang, Gautam Srivastava, Jerry Chun-Wei Lin, Yaoqi Yang, Mamoun Alazab, and Thippa Reddy Gadekallu. 2023a. Data freshness optimization under CAA in the UAV-aided MECN: A potential game perspective. IEEE Trans. Intell. Transp. Syst. 24, 11 (Nov. 2023), 12912–12921.
- Wang et al. (2021b) Xiaoding Wang, Hui Lin, Hongyan Zhang, Dejun Miao, Qinyang Miao, and Wenxin Liu. 2021b. Intelligent drone-assisted fault diagnosis for B5G-enabled space-air-ground-space networks. IEEE Trans. Netw. Sci. Eng. 8, 4 (Oct. 2021), 2849–2860.
- Wang et al. (2021a) Yuhang Wang, Ying He, and Minhui Dong. 2021a. Resource allocation in vehicular networks with multi-UAV served edge computing. In 2021 IEEE 29th International Conference on Network Protocols (ICNP) (Dallas, TX, USA). IEEE.
- Wang et al. (2022c) Yuntao Wang, Zhou Su, Tom H Luan, Ruidong Li, and Kuan Zhang. 2022c. Federated learning with fair incentives and robust aggregation for UAV-aided crowdsensing. IEEE Trans. Netw. Sci. Eng. 9, 5 (Sept. 2022), 3179–3196.
- Wang et al. (2021c) Yuntao Wang, Zhou Su, Ning Zhang, and Dongfeng Fang. 2021c. Disaster relief wireless networks: Challenges and solutions. IEEE Wirel. Commun. 28, 5 (Oct. 2021), 148–155.
- Wang et al. (2020c) Yangang Wang, Hai Wang, and Xianglin Wei. 2020c. Energy-efficient UAV deployment and task scheduling in multi-UAV edge computing. In 2020 International Conference on Wireless Communications and Signal Processing (WCSP) (Nanjing). IEEE.
- Wang et al. (2024b) Yangang Wang, Xianglin Wei, Hai Wang, Yongyang Hu, Kuang Zhao, and Jianhua Fan. 2024b. Distributed matching theory-based task re-allocating for heterogeneous multi-UAV edge computing. China Commun. 21, 1 (Jan. 2024), 260–278.
- Wang et al. (2024d) Yiqian Wang, Jie Zhu, Haiping Huang, and Fu Xiao. 2024d. Bi-objective ant colony optimization for trajectory planning and task offloading in UAV-assisted MEC systems. IEEE Trans. Mob. Comput. 23, 12 (Dec. 2024), 12360–12377.
- Wang et al. (2024a) Zhihong Wang, Gaocai Wang, and Shuqiang Huang. 2024a. Energy-efficient mobile edge computing assisted by layered UAVs based on convex optimization. Phys. Commun. 65, 102382 (Aug. 2024), 102382.
- Wang et al. (2019b) Zhifang Wang, Jianguo Yu, Shangjing Lin, Junguo Dong, and Zhen Yu. 2019b. Distributed robust adaptive fault-tolerant mechanism for quadrotor UAV real-time wireless network systems with random delay and packet loss. IEEE Access 7 (2019), 134055–134062.
- Wang et al. (2021d) Zhifang Wang, Jianguo Yu, Shangjing Lin, Junguo Dong, and Zheng Yu. 2021d. Distributed robust H adaptive fault-tolerant control of amorphous and flat air-to-ground wireless self-organizing network system. Assem. Autom. 41, 6 (Nov. 2021), 641–658.
- Wassim et al. (2024) Lillian Wassim, Kamal Mohamed, and Ali Hamdi. 2024. LLM-DaaS: LLM-driven Drone-as-a-Service Operations from Text User Requests. arXiv preprint arXiv:2412.11672 (2024).
- Watanabe and Machida (2022) Kengo Watanabe and Fumio Machida. 2022. Availability analysis of a drone system with proactive offloading for software life-extension. In 2022 IEEE International Conference on Omni-layer Intelligent Systems (COINS) (Barcelona, Spain). IEEE, 1–6.
- Wei et al. (2021) Dawei Wei, Jianfeng Ma, Linbo Luo, Yunbo Wang, Lei He, and Xinghua Li. 2021. Computation offloading over multi-UAV MEC network: A distributed deep reinforcement learning approach. Comput. Netw. 199, 108439 (Nov. 2021), 108439.
- Wei et al. (2022) Xianglin Wei, Li Li, Lingfeng Cai, Chaogang Tang, and Suresh Subramaniam. 2022. Joint service-function deployment and task scheduling in UAVFog-assisted data-driven disaster response architecture. World Wide Web 25, 1 (Jan. 2022), 309–333.
- Willers et al. (2020) Oliver Willers, Sebastian Sudholt, Shervin Raafatnia, and Stephanie Abrecht. 2020. Safety concerns and mitigation approaches regarding the use of deep learning in safety-critical perception tasks. In Computer Safety, Reliability, and Security. SAFECOMP 2020 Workshops: DECSoS 2020, DepDevOps 2020, USDAI 2020, and WAISE 2020, Lisbon, Portugal, September 15, 2020, Proceedings 39. Springer, 336–350.
- Wu et al. (2023) Hongyue Wu, Mengchen Wu, Wen Peng, Shizhan Chen, and Zhiyong Feng. 2023. ITS: Improved Tabu search algorithm for path planning in UAV-assisted edge computing systems. In 2023 IEEE International Conference on Web Services (ICWS) (Chicago, IL, USA). IEEE, 340–349.
- Wu et al. (2024) Jiaqi Wu, Rui Jing, Yishuo Bai, Zijian Tian, Wei Chen, Shihao Zhang, F Richard Yu, and Victor C M Leung. 2024. Small insulator defects detection based on multiscale feature interaction transformer for UAV-assisted power IoVT. IEEE Internet Things J. 11, 13 (July 2024), 23410–23427.
- Wu et al. (2019) Qingqing Wu, Liang Liu, and Rui Zhang. 2019. Fundamental trade-offs in communication and trajectory design for UAV-enabled wireless network. IEEE Wireless Communications 26, 1 (2019), 36–44.
- Xia et al. (2023) Xiaoyu Xia, Sheik Mohammad Mostakim Fattah, and Muhammad Ali Babar. 2023. A survey on UAV-enabled edge computing: Resource management perspective. Comput. Surveys 56, 3 (2023), 1–36.
- Xiang et al. (2024) Biqun Xiang, Bo Zhong, Anhua Wang, Wuping Mao, and Liang Liu. 2024. Edge computing collaborative offloading strategy for space‐air‐ground integrated networks. Concurr. Comput. (June 2024).
- Xiong et al. (2019) Jingyu Xiong, Hongzhi Guo, Jiajia Liu, Nei Kato, and Yanning Zhang. 2019. Collaborative computation offloading at UAV-enhanced edge. In 2019 IEEE Global Communications Conference (GLOBECOM) (Waikoloa, HI, USA). IEEE.
- Xu and Zhai (2020) Desheng Xu and Junyi Zhai. 2020. Research on aviation man-made error detection and its impact on flight safety. In 2020 7th IEEE International Conference on Cyber Security and Cloud Computing (CSCloud)/2020 6th IEEE International Conference on Edge Computing and Scalable Cloud (EdgeCom) (New York, NY, USA). IEEE.
- Xu et al. (2020) Fangcheng Xu, Xiangbin Yu, Jiali Cai, and Guangying Wang. 2020. Computation efficiency optimization in UAV-enabled mobile edge computing system with multi-carrier non-orthogonal multiple access. EURASIP J. Wirel. Commun. Netw. 2020, 1 (Dec. 2020).
- Xu et al. (2024b) Frank F Xu, Yufan Song, Boxuan Li, Yuxuan Tang, Kritanjali Jain, Mengxue Bao, Zora Z Wang, Xuhui Zhou, Zhitong Guo, Murong Cao, et al. 2024b. Theagentcompany: Benchmarking llm agents on consequential real world tasks, 2024. URL https://arxiv. org/abs/2412.14161 (2024).
- Xu et al. (2022) Jiajie Xu, Dejuan Li, Wei Gu, and Ying Chen. 2022. UAV-assisted task offloading for IoT in smart buildings and environment via deep reinforcement learning. Build. Environ. 222, 109218 (Aug. 2022), 109218.
- Xu et al. (2024c) Lin Xu, Xinmiao Zhang, Dong Xiao, Beihong Liu, and Aixue Liu. 2024c. Research on heterogeneous multi-UAV collaborative decision-making method based on improved PPO. Appl. Intell. 54, 20 (Oct. 2024), 9892–9905.
- Xu et al. (2024a) Yi Xu, Fengguang Luan, Jonathan Kua, Haoyu Luo, Zhipeng Wang, and Xiao Liu. 2024a. Multi-UAV collaborative face recognition for goods receiver in edge-based smart delivery services. In Lecture Notes in Computer Science. Springer Nature Singapore, Singapore, 217–235.
- Xu et al. (2021) Zhengjia Xu, Ivan Petrunin, and Antonios Tsourdos. 2021. Dynamic spectrum management with network function virtualization for UAV communication. J. Intell. Robot. Syst. 101, 2 (Feb. 2021).
- Xue et al. (2023a) Jianbin Xue, Yuling Ma, Guiying Tian, Jun Dou, and Xiangrui Guan. 2023a. Resource allocation and offloading decision of UAV based on NOMA-MEC. Wirel. Pers. Commun. 133, 1 (Nov. 2023), 259–288.
- Xue et al. (2023c) Jianbin Xue, Tingjuan Zhang, Fei Shao, Jia Yao, and Xiaofeng Xu. 2023c. Joint optimization of trajectory and resource allocation in cellular-connected multi-UAV MEC networks. Phys. Commun. 61, 102209 (Dec. 2023), 102209.
- Xue et al. (2023b) Z Xue, C Xu, R Cao, and S Zhang. 2023b. Joint energy management and trajectory optimization for UAV-enabled crane monitoring system. IET Conf. Proc. 2022, 24 (Feb. 2023), 40–46.
- Yadav and Baranwal (2023) Ravi Yadav and Gaurav Baranwal. 2023. ReTREM: A responsibility based trust revision model for determining trustworthiness of fog nodes. Comput. Commun. 197 (Jan. 2023), 159–172.
- Yan et al. (2024) Yiping Yan, Kai Wang, and Xiaobo Qu. 2024. Urban air mobility (UAM) and ground transportation integration: A survey. Frontiers of Engineering Management (2024), 1–25.
- Yang et al. (2020) Lei Yang, Haipeng Yao, Jingjing Wang, Chunxiao Jiang, Abderrahim Benslimane, and Yunjie Liu. 2020. Multi-UAV-enabled load-balance mobile-edge computing for IoT networks. IEEE Internet Things J. 7, 8 (Aug. 2020), 6898–6908.
- Yang et al. (2019) Xin Yang, Hongcheng Luo, Yuhao Wu, Yang Gao, Chunyuan Liao, and Kwang-Ting Cheng. 2019. Reactive obstacle avoidance of monocular quadrotors with online adapted depth prediction network. Neurocomputing 325 (Jan. 2019), 142–158.
- Yang et al. (2022) Xiaohua Yang, Xiao Ye, Jianing Cao, Rong Yan, Xiaojue Guo, and Jisheng Huang. 2022. High-altitude inspection technology of substation based on fusion of unmanned aerial vehicle and multiple sensors. Sens. Mater. 34, 8 (Aug. 2022), 3191.
- Yang et al. (2017) Yan-Ming Yang, Hui Yu, and Zhi Sun. 2017. Aircraft failure rate forecasting method based on Holt-Winters seasonal model. In 2017 IEEE 2nd International Conference on Cloud Computing and Big Data Analysis (ICCCBDA) (Chengdu, China). IEEE.
- Yang LI (2024) Jinshu SU Baokang ZHAO Yang LI, Ziling WEI. 2024. A multi-agent collaboration scheme for energy-efficient task scheduling in a 3D UAV-MEC space. Frontiers of Information Technology and Electronic Engineering 25, 6 (2024), 824–838. doi:10.1631/FITEE.2300393
- Yao and Ansari (2018) Jingjing Yao and Nirwan Ansari. 2018. Reliability-aware fog resource provisioning for deadline-driven IoT services. In 2018 IEEE Global Communications Conference (GLOBECOM) (Abu Dhabi, United Arab Emirates). IEEE.
- Ye et al. (2020) Weidu Ye, Junzhou Luo, Feng Shan, Wenjia Wu, and Ming Yang. 2020. Offspeeding: Optimal energy-efficient flight speed scheduling for UAV-assisted edge computing. Comput. Netw. 183, 107577 (Dec. 2020), 107577.
- Ye et al. (2022) Weidu Ye, Junzhou Luo, Wenjia Wu, Feng Shan, and Ming Yang. 2022. MUTAA: An online trajectory optimization and task scheduling for UAV-aided edge computing. Comput. Netw. 218, 109405 (Dec. 2022), 109405.
- Yoshimoto et al. (2020) Jo Yoshimoto, Ittetsu Taniguchi, Hiroyuki Tomiyama, and Takao Onoye. 2020. An evaluation of edge computing platform for reliable automated drones. In 2020 International SoC Design Conference (ISOCC) (Yeosu, Korea (South)). IEEE.
- Yu et al. (2022a) Changliang Yu, Wei Du, Fan Ren, and Nan Zhao. 2022a. Deep reinforcement learning for task allocation in UAV-enabled mobile edge computing. In Lecture Notes in Networks and Systems. Springer International Publishing, Cham, 225–232.
- Yu et al. (2023b) Fan Yu, Dingcheng Yang, Fahui Wu, Yapeng Wang, and Hao He. 2023b. Resource optimization for UAV-assisted mobile edge computing system based on deep reinforcement learning. Phys. Commun. 59, 102107 (Aug. 2023), 102107.
- Yu et al. (2023a) Xiaoyong Yu, Liwen Qin, Xin Chen, Lifang Wu, and Biyun Zhang. 2023a. Design and development of an onboard intelligent device based on Jetson NX. In 2023 4th International Conference on Computer Engineering and Intelligent Control (ICCEIC) (Guangzhou, China). IEEE.
- Yu et al. (2022b) Xue-Yong Yu, Wen-Jin Niu, Ye Zhu, and Hong-Bo Zhu. 2022b. UAV-assisted cooperative offloading energy efficiency system for mobile edge computing. Digit. Commun. Netw. (March 2022).
- Yu et al. (2021) Ye Yu, Xiangyuan Bu, Kai Yang, Hongyuan Yang, Xiaozheng Gao, and Zhu Han. 2021. UAV-aided low latency multi-access edge computing. IEEE Trans. Veh. Technol. 70, 5 (May 2021), 4955–4967.
- Yuan et al. (2022) Xinwang Yuan, Zhidong Xie, and Xin Tan. 2022. Computation offloading in UAV-enabled edge computing: A Stackelberg game approach. Sensors (Basel) 22, 10 (May 2022), 3854.
- Yue et al. (2018) Xuejun Yue, Yongxin Liu, Jian Wang, Houbing Song, and Huiru Cao. 2018. Software defined radio and wireless acoustic networking for amateur drone surveillance. IEEE Commun. Mag. 56, 4 (April 2018), 90–97.
- Yuwei and Aijun (2020) Cui Yuwei and Li Aijun. 2020. Research on the key technologies of network distributed flight control system. In 2020 Chinese Automation Congress (CAC) (Shanghai, China). IEEE.
- Zeng et al. (2024b) Kai Zeng, Xin Li, and Tao Shen. 2024b. Energy-stabilized computing offloading algorithm for UAVs with energy harvesting. IEEE Internet Things J. 11, 4 (Feb. 2024), 6020–6031.
- Zeng et al. (2024a) Liekang Zeng, Haowei Chen, Daipeng Feng, Xiaoxi Zhang, and Xu Chen. 2024a. A3D: Adaptive, accurate, and autonomous navigation for edge-assisted drones. IEEE ACM Trans. Netw. 32, 1 (Feb. 2024), 713–728.
- Zeng et al. (2023) Yaoping Zeng, Dongyang Lu, and Jianbo Du. 2023. Joint optimized multi-user access and UAV deployments based on heterogeneous revenue in IoT network. Comput. Netw. 234, 109919 (Oct. 2023), 109919.
- Zhan and Xu (2024) Mengdi Zhan and Shubin Xu. 2024. Delay-aware distributionally robust trajectory UAV-assisted MEC with uncertain task size. In 2024 IEEE Wireless Communications and Networking Conference (WCNC) (Dubai, United Arab Emirates). IEEE, 1–6.
- Zhang et al. (2021a) Chong Zhang, Xiao Liu, Jia Xu, Tianxiang Chen, Gang Li, Frank Jiang, and Xuejun Li. 2021a. An edge based federated learning framework for person re-identification in UAV delivery service. In 2021 IEEE International Conference on Web Services (ICWS) (Chicago, IL, USA). IEEE.
- Zhang et al. (2024e) Degan Zhang, Zhihao Zhang, Jie Zhang, Ting Zhang, Lei Zhang, and Hongtao Chen. 2024e. UAV-assisted task offloading system using dung beetle optimization algorithm & deep reinforcement learning. Ad Hoc Netw. 156, 103434 (April 2024), 103434.
- Zhang et al. (2024a) Gaofeng Zhang, Xiang Li, Liqiang Xu, Ensheng Liu, Liping Zheng, Wenming Wu, and Benzhu Xu. 2024a. UGV-awareness task placement in edge-cloud based urban intelligent video systems. Cluster Comput. 27, 5 (Aug. 2024), 6563–6577.
- Zhang et al. (2022a) Guanghua Zhang, Xiaoxin Wang, Mi Shan, Jingqiu Ren, and Weidang Lu. 2022a. Time allocation improvement method for UAV-based wireless energy transfer cooperative mobile edge. Phys. Commun. 55, 101861 (Dec. 2022), 101861.
- Zhang et al. (2024f) Guoqing Zhang, Yongxiang Zhao, Ping Fu, Wei Luo, Quanqin Shao, Tongzuo Zhang, and Zhongde Yu. 2024f. A reliable unmanned aerial vehicle multi-target tracking system with global motion compensation for monitoring Procapra przewalskii. Ecol. Inform. 81, 102556 (July 2024), 102556.
- Zhang and Bai (2024) Jing Zhang and Jingpan Bai. 2024. Dijkstra algorithm based cooperative caching strategy for UAV-assisted edge computing system. Wirel. Netw. 30, 3 (April 2024), 1201–1219.
- Zhang and Li (2018) Lei Zhang and Chen Li. 2018. Attitude determination system of UAVs based on common-clock GPS/BDS. In 2018 5th IEEE International Conference on Cloud Computing and Intelligence Systems (CCIS) (Nanjing, China). IEEE.
- Zhang et al. (2024b) Longxin Zhang, Runti Tan, Yanfen Zhang, Jiwu Peng, Jing Liu, and Keqin Li. 2024b. UAV-assisted dependency-aware computation offloading in device–edge–cloud collaborative computing based on improved actor–critic DRL. J. Syst. Arch. 154, 103215 (Sept. 2024), 103215.
- Zhang et al. (2018) Long Zhang, Zhen Zhao, Qiwu Wu, Hui Zhao, Haitao Xu, and Xiaobo Wu. 2018. Energy-aware dynamic resource allocation in UAV assisted mobile edge computing over social internet of vehicles. IEEE Access 6 (2018), 56700–56715.
- Zhang et al. (2023d) Peng Zhang, Hui Tian, Hao Luo, Xingwang Li, and Gaofeng Nie. 2023d. A hybrid fast inference approach with distributed neural networks for edge computing enabled UAV swarm. Phys. Commun. 60, 102129 (Oct. 2023), 102129.
- Zhang et al. (2021b) Shangwei Zhang, Jiajia Liu, Yajie Zhu, and Jing Zhang. 2021b. Joint computation offloading and trajectory design for aerial computing. IEEE Wirel. Commun. 28, 5 (Oct. 2021), 88–94.
- Zhang et al. (2023a) Tiankui Zhang, Chaobin Chen, Yu Xu, Jonathan Loo, and Wenjun Xu. 2023a. Joint task scheduling and multi-UAV deployment for aerial computing in emergency communication networks. Sci. China Inf. Sci. 66, 9 (Sept. 2023).
- Zhang et al. (2021c) Weishan Zhang, Haoyun Sun, Dehai Zhao, Liang Xu, Xin Liu, Huansheng Ning, Jiehan Zhou, Yi Guo, and Su Yang. 2021c. A streaming cloud platform for real-time video processing on embedded devices. IEEE Trans. Cloud Comput. 9, 3 (July 2021), 868–880.
- Zhang et al. (2023c) Xuhui Zhang, Wenchao Liu, Huijun Xing, Zhenzhen Jin, Weilin Zang, Shuqiang Wang, Yanyan Shen, and Liang Xue. 2023c. Learning to hybrid offload in space-air-ground integrated mobile edge computing for IoT networks. In 2023 IEEE 13th International Conference on CYBER Technology in Automation, Control, and Intelligent Systems (CYBER) (Qinhuangdao, China). IEEE, 836–841.
- Zhang et al. (2024d) Xiaoqi Zhang, Haijun Zhang, Kai Sun, Keping Long, and Yonghui Li. 2024d. Human-centric irregular RIS-assisted multi-UAV networks with resource allocation and reflecting design for metaverse. IEEE J. Sel. Areas Commun. 42, 3 (March 2024), 603–615.
- Zhang and Zhang (2024) Xuchun Zhang and Xiaokun Zhang. 2024. Application of radio frequency identification technology in maintenance tracking of aircraft electromechanical equipment. In 2024 International Conference on Electrical Drives, Power Electronics & Engineering (EDPEE) (Athens, Greece). IEEE.
- Zhang (2020) Yue Zhang. 2020. Aeroengine fault prediction based on bidirectional LSTM neural network. In 2020 International Conference on Big Data, Artificial Intelligence and Internet of Things Engineering (ICBAIE) (Fuzhou, China). IEEE.
- Zhang et al. (2023b) Yafeng Zhang, Yongzheng Gan, Chunlin Li, Chunping Deng, and Youlong Luo. 2023b. Primary node selection based on node reputation evaluation for PBFT in UAV-assisted MEC environment. Wirel. Netw. 29, 8 (Nov. 2023), 3515–3539.
- Zhang et al. (2024c) Yiguo Zhang, Wei Wang, and Feiyang Shi. 2024c. Reputation-based Raft-Poa layered consensus protocol converging UAV network. Comput. Netw. 240, 110170 (Feb. 2024), 110170.
- Zhang et al. (2023e) Ying Zhang, Supu Xiu, Yiqing Cai, and Pengshan Ren. 2023e. Scheduling of graph neural network and Markov based UAV mobile edge computing networks. Phys. Commun. 60, 102160 (Oct. 2023), 102160.
- Zhang et al. (2022b) Zhuoya Zhang, Fei Xu, Zengshi Qin, and Yue Xie. 2022b. Resource allocation in UAV assisted air ground intelligent inspection system. Cognitive Robotics 2 (2022), 1–12.
- Zhang et al. (2025) Zhiyang Zhang, Fengli Zhang, Minsheng Cao, Chaosheng Feng, and Dajiang Chen. 2025. Enhancing UAV-assisted vehicle edge computing networks through a digital twin-driven task offloading framework. Wirel. Netw. 31, 1 (Jan. 2025), 965–981.
- Zhao et al. (2024) Hui Zhao, Gengyuan Lu, Ying Liu, Zheng Chang, Li Wang, and Timo Hämäläinen. 2024. Safe DQN-based AoI-minimal task offloading for UAV-aided edge computing system. IEEE Internet Things J. (2024), 1–1.
- Zhao et al. (2022) Mingyue Zhao, Xing Zhang, Zezhao Meng, and Xiangwang Hou. 2022. Reliable DNN partitioning for UAV swarm. In 2022 International Wireless Communications and Mobile Computing (IWCMC) (Dubrovnik, Croatia). IEEE.
- Zheng et al. (2024) Chunxiang Zheng, Kai Pan, Jiadong Dong, Lin Chen, Qinghu Guo, Shunfeng Wu, Hanyun Luo, and Xiaolin Zhang. 2024. Multi-agent collaborative optimization of UAV trajectory and latency-aware DAG task offloading in UAV-assisted MEC. IEEE Access 12 (2024), 42521–42534.
- Zheng et al. (2021) Jibin Zheng, Tianyuan Yang, Hongwei Liu, Tao Su, and Liangtian Wan. 2021. Accurate detection and localization of unmanned aerial vehicle swarms-enabled mobile edge computing system. IEEE Trans. Industr. Inform. 17, 7 (July 2021), 5059–5067.
- Zheng et al. (2023) Shiji Zheng, Xiaoqi Zhong, Bang Huang, and Siying Zhuang. 2023. Research on intelligent mobile edge computing and task unloading method of UAV. In 2023 IEEE 6th Information Technology,Networking,Electronic and Automation Control Conference (ITNEC) (Chongqing, China). IEEE.
- Zhi et al. (2020) Yueyan Zhi, Zhangjie Fu, Xingming Sun, and Jingnan Yu. 2020. Security and privacy issues of UAV: A survey. Mobile Networks and Applications 25, 1 (2020), 95–101.
- Zhong et al. (2021) Yerong Zhong, Guoheng Ruan, and Jiaming Jiang. 2021. Deep learning-based comprehensive monitor for smart power station. Int. J. Grid Util. Comput. 12, 4 (2021), 380.
- Zhou and Cao (2018) Fujiao Zhou and Kecai Cao. 2018. Formation fault tolerant control of multi - UAV based on virtual actuator. In 2018 5th IEEE International Conference on Cloud Computing and Intelligence Systems (CCIS) (Nanjing, China). IEEE.
- Zhou et al. (2022b) Jianshan Zhou, Shiyi Liu, Xu Han, Daxin Tian, and Victor C.M. Leung. 2022b. Ultra-Reliable Computation Offloading for Aerial-Ground Cooperative Vehicular Networks with Joint Mobility Optimization and Transmission Scheduling. In Proceedings of the 12th ACM International Symposium on Design and Analysis of Intelligent Vehicular Networks and Applications (Montreal, Quebec, Canada) (DIVANet ’22). Association for Computing Machinery, New York, NY, USA, 9–16. doi:10.1145/3551662.3561200
- Zhou et al. (2023) Xunkuai Zhou, Guidong Yang, Yizhou Chen, Chuangxiang Gao, Benyun Zhao, Li Li, and Ben M Chen. 2023. ADMNet: Anti-drone real-time detection and monitoring. In 2023 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS) (Detroit, MI, USA). IEEE, 3009–3016.
- Zhou et al. (2024c) Xunkuai Zhou, Guidong Yang, Yizhou Chen, Li Li, and Ben M Chen. 2024c. VDTNet: A high-performance visual network for detecting and tracking of intruding drones. IEEE Trans. Intell. Transp. Syst. 25, 8 (Aug. 2024), 9828–9839.
- Zhou et al. (2022a) Yu Zhou, Hui Ge, Bowen Ma, Shuhang Zhang, and Jiwei Huang. 2022a. Collaborative task offloading and resource allocation with hybrid energy supply for UAV-assisted multi-clouds. J. Cloud Comput. Adv. Syst. Appl. 11, 1 (Sept. 2022).
- Zhou et al. (2024a) Yatong Zhou, Xiaoran Kong, Kuo-Ping Lin, and Liangyu Liu. 2024a. Novel task decomposed multi-agent twin delayed deep deterministic policy gradient algorithm for multi-UAV autonomous path planning. Knowl. Based Syst. 287, 111462 (March 2024), 111462.
- Zhou et al. (2024b) Yi Zhou, Zheng Ma, Gang Liu, Zhengquan Zhang, Phee Lep Yeoh, Branka Vucetic, and Yonghui Li. 2024b. Secure multi-layer MEC systems with UAV-enabled reconfigurable intelligent surface against full-duplex eavesdropper. IEEE Trans. Commun. 72, 3 (March 2024), 1565–1577.
- Zhu et al. (2024) Lidong Zhu, Michele Luglio, Zhili Sun, Gengxin Zhang, and Mingchuan Yang. 2024. Resilient satellite communication networks towards highly dynamic and highly reliable transmission. China Communications 21, 2 (2024), iii–v.
- Zou et al. (2022) Yufei Zou, Li Lin, and Lei Zhang. 2022. A task offloading strategy for compute-intensive scenarios in UAV-assisted IoV. In 2022 IEEE 5th International Conference on Electronic Information and Communication Technology (ICEICT) (Hefei, China). IEEE, 427–431.
- Zuo et al. (2022) Yuan Zuo, Wen Yao, Qiang Chang, Xiaozhou Zhu, Jianjun Gui, and Jiangyi Qin. 2022. Voting-based scheme for leader election in lead-follow UAV swarm with constrained communication. Electronics (Basel) 11, 14 (July 2022), 2143.